### v2

# Preprint Manuscript

Preprint Manuscript
Second Version
CUBIC INCOMPLETENESS: HILBERT’S TENTH PROBLEM BEGINS AT DEGREE THREE
Milan Rosko
(September 2025)
We resolve a long-standing open question in the affirmative: cubic Diophantine equations are undecidable. By encoding Gödel’s incompleteness theorem via Fibonacci-based Gödel numbering and Zeckendorf representation, combined with a systematic guard-gadget framework, we construct an explicit degree-3 polynomial $P_{\text{hard}}\in\mathbb{Z}[x_{1},\ldots,x_{M}]$ whose solvability over $\mathbb{N}$ is equivalent to provability of a Gödel sentence in Peano Arithmetic. We establish three results. First, a bidirectional reduction proving $P_{\text{hard}}(\vec{x})=0$ has solutions if and only if $G_{T}$ is provable. Second, a uniform construction showing that for any Turing machine $M$, we can effectively produce a cubic polynomial $Q_{M}(\vec{u})\in\mathbb{Z}[\vec{u}]$ such that $Q_{M}(\vec{u})=0$ has a solution in $\mathbb{N}^{k}$ if and only if $M$ halts, establishing undecidability of the class $\delta=3$. Third, we formalize an impredicativity thesis: no consistent theory extending Robinson arithmetic can define the minimal degree at which Diophantine undecidability emerges.
Keywords: Proof Theory, Logic, Hilbert’s Tenth Problem, Zeckendorf’s Theorem
Open Problem, Constructive Mathematics, Fibonacci, Cubic equations
Primary MSC: 11U05, 03F40, 03D35
Secondary MSC: 11D72, 03B25, 11Y16
Milan Rosko is from University of Hagen, Germany
Email: Q1012878 @@ studium.fernuni-hagen.de – ORCID: 0009-0003-1363-7158
Licensed under “Deed” ΓΔ  – creativecommons.org/licenses/by/4.0
1 Overview
1.1 Corrigendum
The present revision stands as a clarification of intent and method. What remains unchanged is the theoretical vision; what has been refined is the machinery by which that vision finds expression. We now provide a systematic theoretical enumeration of all polynomial degrees throughout the encoding. Critically, the earlier version contained a fatal error: the result of the “promissory” shielding process was incorrectly claimed to be of degree three. In this revision, we omit the problematic equation and focus instead on explicit witness construction, ensuring all constraints remain genuinely cubic.
1.2 Introduction
Previously [^1] A Fibonacci-Based Gödel Numbering: $\Delta_{0}$ Semantics Without Exponentiation, [2025], arXiv:2509.10382 [math.LO], https://arxiv.org/abs/2509.10382, we proposed that a Zeckendorf-based arithmetization of Gödel numbering—eschewing exponentiation in favor of additive structure—might lower the formal overhead in encoding undecidability. We now apply this method to produce undecidable Diophantine equations of degree three, thereby advancing “Hilbert’s Tenth Problem” [^2] As posed by Hilbert [1900]. into the cubic regime ($\delta=3$) and resolving a previously open case: We cite from the article “Universal Diophantine Equation”, by Jones [1982]:
We know that there exist universal diophantine equations of degree 4 […] but none of degree 2. […] Thus the only open problem as regards the degree is $\delta$ = 3. We do not know whether Hilbert’s tenth problem is decidable or undecidable for equations of degree 3.
Hence, we define the open question:
Problem 1.1 (Cubic Remainder of Hilbert’s 10th Problem).
Let $\delta=3$ denote the set of all polynomial Diophantine equations of the form
$P(x_{1},\ldots,x_{n})=0,\quad P\in\mathbb{Z}[x_{1},\ldots,x_{n}],\quad\delta(P)\leq 3.$ (1.1)
Question: Is there an algorithm that, given $P$ and $\delta$ = 3, decides whether
$\exists(x_{1},\ldots,x_{n})\in\mathbb{N}^{n}:P(x_{1},\ldots,x_{n})=0?$ (1.2)
1.3 Contribution
We construct a witness of polynomial Diophantine equation of total degree at most three, whose solvability over $\mathbb{N}$ is equivalent to the provability of a Gödel sentence in Peano Arithmetic extended with Fibonacci-based encoding. This resolves the question of whether undecidability can be exhibited at cubic degree, thereby advancing the study of Hilbert’s Tenth Problem into the low-degree regime.
Theorem 1.2 (Main Result).
There exists a polynomial $P_{\text{hard}}\in\mathbb{Z}[x_{1},\ldots,x_{M}]$ of $\delta(P_{\text{hard}})\leq 3$ and a sentence $G_{T}$ in the language of arithmetic such that:
$\exists\vec{x}\in\mathbb{N}^{M}\colon P_{\text{hard}}(\vec{x})=0\quad\Longleftrightarrow\quad T_{\text{Fib}}\vdash G_{T},$ (1.3)
where $T_{\text{Fib}}$ denotes Peano Arithmetic with Fibonacci-based Gödel numbering, and $G_{T}$ is the canonical Gödel sentence asserting its own unprovability.
The logic encodes proof-theoretic relations—axiom membership, modus ponens, and sequence verification—into polynomial constraints of controlled degree. The key innovation is a guard-gadget framework that transforms potentially negative constraints into sums of squares, preserving Diophantine equivalence while maintaining cubic degree bounds. The results of Matiyasevich [1970] and Robinson et al. [1961] establishes that every recursively enumerable set can be represented by a Diophantine equation. Subsequent refinements by Jones [1982] proved realizability at $\delta\leq 4$. We build on this tradition in two key ways:
  1. 1.
    By Explicit proof verification: We forward the arithmic content of a formal proof (axioms, modus ponens) directly in $\Delta_{0}$, rather than constructing a $\Sigma_{1}$ abstraction to minimize overhead.
  2. 2.
    By Finite axiom sets: By working with a precalculated representations, we avoid the need to encode an infinite, or expressive axiom scheme.
The trade-off is that our polynomial is potentially “wider” (in number of variables and monomials) than a optimal $\Sigma_{1}$ construction, but the argument is arguably more elementary (e.g., in terms of complexity), relying primarily on a variant of Gödel’s $\beta$-function (Theorem 2.1) and Zeckendorf representation (Definition 2.3). Having outlined our approach, we now formalize the key technical preliminaries underpinning the construction.
2 Preliminaries
2.1 The Gödel Beta-Function
The following logic allows encoding arbitrary finite sequences as Diophantine constraints:
Procedure 2.1 (Diophantine $\beta$-Function).
By the work of Gödel [1931] (using the “Chinese remainder”) and as exposited in Matiyasevich [1993], there exists a polynomial $\beta(c,d,i)$ over $\mathbb{N}$ such that for any finite sequence $a_{0},a_{1},\ldots,a_{n}\in\mathbb{N}$, there exist $c,d\in\mathbb{N}$ with
$\beta(c,d,i)=a_{i}\quad\text{for all }i\in\{0,1,\ldots,n\}.$ (2.4)
Explicitly, $\beta(c,d,i)=r_{i}$, where $r_{i}$ is the remainder and $q_{i}\in\mathbb{N}$ is the quotient when dividing $c$ by $(1+(i+1)d)$:
$c=q_{i}\cdot(1+(i+1)d)+r_{i},\quad 0\leq r_{i}<1+(i+1)d.$ (2.5)
Remark.
The “Chinese remainder” allows us to choose $d$ such that the moduli $\{1+(i+1)d:0\leq i\leq n\}$ are pairwise coprime. A suitable choice is
$d=\text{lcm}(1,2,\ldots,n+1),$ (2.6)
which ensures that for any $i\neq j$, the moduli $1+(i+1)d$ and $1+(j+1)d$ differ by (ji)d(j-i)d, and this difference shares no common factor with either modulus, since
$1+(k+1)d\equiv 1\pmod{p}\quad\text{for all primes}\quad p\mid d\quad\text{with}\quad p\leq n+1.$
Hence we construct $c$ via CRT to satisfy $c\equiv a_{i}\pmod{1+(i+1)d}$ for all $i\leq n$. To verify pairwise coprimality with $d=\text{lcm}(1,\ldots,n+1)$: for $i<j\leq n$, we have
$\gcd(1+(i+1)d,1+(j+1)d)\mid\gcd(1+(i+1)d,(j-i)d).$ (2.7)
Since $(j-i)\leq n$ and $d$ is divisible by all integers up to $n+1$, while $1+(i+1)d\equiv 1\pmod{p}$ for any prime $p\mid d$ with $p\leq n+1$, the greatest common divisor equals $1$. This ensures the CRT applies. To express $\beta(c,d,i)=f_{i}$ as polynomial constraints, we introduce variables $r_{i},q_{i}$ and enforce:
$c$ $=q_{i}\cdot(1+(i+1)d)+r_{i},$ (2.8)
$0$ $\leq r_{i}<1+(i+1)d.$ (2.9)
The second challenge, after reducing $\Sigma_{1}$ logic to $\Delta_{0}$ [^3] Foundational reference by Kleene [1943] introducing concepts underlying the arithmetic hierarchy, is encoding inequality polynomially. A naive preliminary approach using $r_{i}=s_{i}^{2}$ to force $r_{i}\geq 0$ failed because: (i) the $\beta$-construction produces remainders $r_{i}=c\bmod(1+(i+1)d)$ that need not be perfect squares (e.g., $r_{i}=2$ is not a square), and (ii) encoding the upper bound $r_{i}<1+(i+1)d$ via $(1+(i+1)d-r_{i}-1)=t_{i}^{2}$ would impose the same impossible requirement. The “four-squares” encoding resolves both issues simultaneously without requiring any specific number to be a perfect square.
Lemma 2.2 (Four-Squares Bound Encoding).
The constraint $0\leq r<m$ over $\mathbb{N}$ is equivalent to the existence of integers $a,b,c,e,\alpha,\beta,\gamma,\delta\in\mathbb{N}$ satisfying:
$r$ $=a^{2}+b^{2}+c^{2}+e^{2},$ (2.10)
$m-1-r$ $=\alpha^{2}+\beta^{2}+\gamma^{2}+\delta^{2}.$ (2.11)
These are polynomial equations of degree $2$.
Proof.
The “Four-Square Theorem” of Lagrange [1770] allows us to say:
Every $n\in\mathbb{N}$ is a sum of four squares.
Constraint (3.35) forces $r\geq 0$. Constraint (3.36) forces $r\leq m-1$, hence $r<m$. Given $r^{*}\in\{0,\ldots,m-1\}$, the four-square decompositions can be computed in polynomial time. By the algorithm of Rabin and Shallit [1986] we decompose an integer $n\in\mathbb{N}$ into four squares
$n=a^{2}+b^{2}+c^{2}+e^{2}$ (2.12)
This ensures the encoding is effective. Computable witnesses exist for any valid bound, and can be constructed in polynomial time (randomized) or quasi-polynomial time (deterministic under ERH). ∎
Remark.
In the completeness proof of Theorem 3.16, the witness construction algorithm (Exposition 3.9) must compute four-square decompositions for each remainder $r_{i}$ and its bound $(i+1)d-1-r_{i}$. By the above algorithm, this can be done in time $O(\log((i+1)d))$ per decomposition, yielding total time $O(N\cdot\log(\max_{i}f_{i}))$ for all $N$ proof lines. This is polynomial in the proof parameters, as required.
2.2 Unique Representation
Remark.
As a “rule of thumb”: A sum satisfies Zeckendorf representation (i.e., uses non-consecutive Fibonacci numbers) when there are “gaps” in the sequence of Fibonacci numbers used. e.g., the tuple $\{F_{6},F_{8},F_{10}\}=\{8,21,55\}$.
Definition 2.3 (Zeckendorf Representation).
Every positive integer $n$ admits a unique representation as a sum of non-consecutive Fibonacci numbers:
$n=\sum_{\kappa\in S}F_{\kappa},$ (2.13)
where $S\subseteq\{2,3,4,\ldots\}$ contains no consecutive integers (i.e., if $\kappa\in S$, then $\kappa+1\notin S$), and $F_{k}$ denotes the $k$-th Fibonacci number with $F_{1}=F_{2}=1$ and $F_{k+2}=F_{k+1}+F_{k}$.
Lemma 2.4 (Polynomial Encoding of Zeckendorf Digits).
For each integer $n$ with Zeckendorf representation involving indices up to $K$, introduce binary variables $d_{2},\ldots,d_{K}\in\{0,1\}$ satisfying:
  1. 1.
    Boolean constraint: $d_{\kappa}-d_{\kappa}^{2}=0$ for all $\kappa\in\{2,\ldots,K\}$,
  2. 2.
    Non-adjacency: $d_{\kappa}\cdot d_{\kappa+1}=0$ for all $\kappa\in\{2,\ldots,K-1\}$,
  3. 3.
    Reconstruction: $n=\sum_{\kappa=2}^{K}F_{\kappa}\cdot d_{\kappa}$.
These constraints are polynomial of degree at most $2$.
Proof.
The boolean constraint $d_{\kappa}(1-d_{\kappa})=0$ forces $d_{\kappa}\in\{0,1\}$ over $\mathbb{N}$. The product $d_{\kappa}d_{\kappa+1}=0$ enforces non-adjacency. The sum is linear in the $d_{\kappa}$ variables. ∎
Exposition.
The Zeckendorf representation provides an alternative to prime-power encodings for sequences. We use it to enable Fibonacci addition, which can be implemented via digit-wise operations respecting the non-adjacency constraint (see Section 3.3).
2.3 Combined Encoding
Exposition.
For a proof sequence $\langle f_{1},\ldots,f_{N}\rangle$ where each $f_{i}$ is a Gödel number:
  1. 1.
    Encode the sequence using $\beta(c,d,i)=f_{i}$ via constraint (3.34) and Lemma $2.$
  2. 2.
    Represent each $f_{i}$ in Zeckendorf form using Lemma 2.4.
This dual encoding allows:
  1. 1.
    Variable-length proofs (via existential quantification over $N,c,d$),
  2. 2.
    Arithmetic verification (via Fibonacci addition on Zeckendorf digits).
Regarding the consistency of “Dual Encoding:” Each Gödel number $f_{i}$ is encoded twice:
  1. 1.
    Via the $\beta$-function: $\beta(c,d,i)=f_{i}$ (using variables $c,d,r_{i},q_{i},\ldots$),
  2. 2.
    Via Zeckendorf: $f_{i}=\sum_{\kappa=2}^{K}F_{\kappa}d_{i,\kappa}$ (using digit variables $d_{i,\kappa}$).
These encodings are compatible: for any sequence $\langle f_{1},\ldots,f_{N}\rangle$, witnesses $c^{*},d^{*}$ exist (by Theorem 2.1), and each $f_{i}^{*}$ has a unique Zeckendorf representation (by Definition 2.3). The polynomial system constrains a single variable $f_{i}$ to satisfy both encodings simultaneously, ensuring consistency.
2.4 Summary of Polynomial Constraints
Constraint Equation $n\leq\delta$
$\beta$-Division $c-q_{i}(1+(i+1)d)-r_{i}=0$ 2
Lower bound $r_{i}-a_{i}^{2}-b_{i}^{2}-c_{i}^{2}-e_{i}^{2}=0$ 2
Upper bound $(i+1)d-1-r_{i}-\alpha_{i}^{2}-\beta_{i}^{2}-\gamma_{i}^{2}-\delta_{i}^{2}=0$ 2
Boolean digits $d_{i,\kappa}-d_{i,\kappa}^{2}=0$ 2
Non-adjacency $d_{i,\kappa}\cdot d_{i,\kappa+1}=0$ 2
Zeckendorf sum $f_{i}-\sum_{\kappa=2}^{K}F_{\kappa}d_{i,\kappa}=0$ 1
Table 1: All constraints have degree ($\delta$) $\leq 2$ before applying guard multipliers (see Section 2.5).
2.5 Guard-Gadget Framework
Definition 2.5 (Boolean Constraint).
For any variable $b$ intended to represent a boolean value, the constraint
$b-b^{2}=0,\quad b\in\{0,1\}\subseteq\mathbb{N}$ (2.14)
Definition 2.6 (Guarded Constraint).
For any polynomial $E(\mathbf{x})$ of degree $\delta\leq 2$, introduce guard variable $u$ and slack variable $v\in\mathbb{N}$. The guarded constraint is the system:
$u\cdot E(\mathbf{x})$ $=0,$ (2.15)
$u-1-v^{2}$ $=0.$ (2.16)
Degree bound: If $\delta\leq 2$, then $\delta(u\cdot E)\leq 1+2=3$ and $\delta(u-1-v^{2})=2$.
Proposition 2.7 (Guard Semantics).
The guarded constraint enforces:
  1. 1.
    $u=1+v^{2}\geq 1$ for all solutions (from equation (2.16)),
  2. 2.
    $E(\mathbf{x})=0$ (forced by equation (2.15) since $u\geq 1$).
Proof.
Equation (2.16) forces $u=1+v^{2}\geq 1$. Since $u\neq 0$, equation (2.15) yields $E(\mathbf{x})=0$. ∎
Definition 2.8 (Global Multiplier).
To enable selective activation of guards, introduce a global multiplier $T\geq 1$ via:
$T=1+U^{2},\quad U\in\mathbb{N}.$ (2.17)
This ensures $T\geq 1$ for all solutions. Optionally, this can be enforced via guarded constraints:
$u_{T}\cdot(T-1-U^{2})$ $=0,$ (2.18)
$u_{T}-1-v_{T}^{2}$ $=0,$ (2.19)
where equation (2.19) forces $u_{T}=1+v_{T}^{2}\geq 1$, and equation (2.18) then enforces $T=1+U^{2}$.
Proposition 2.9 (Selective Activation).
Given a guard $u_{i}$ with $u_{i}=1+v_{i}^{2}$ (from Definition 2.6) and a global multiplier $T\geq 1$, the constraint
$T\cdot(1-u_{i})=0$ (2.20)
forces $u_{i}=1$ (equivalently, $v_{i}=0$).
Proof.
Since $T\geq 1$ and $T\cdot(1-u_{i})=0$, we have $1-u_{i}=0$, hence $u_{i}=1$. From $u_{i}=1+v_{i}^{2}$, this implies $v_{i}=0$. ∎
Exposition.
The selective activation mechanism is used to enforce that specific constraints hold in a solution. For example, in proof verification (Section 3.4), we use it to force exactly one justification (axiom or modus ponens) for each line of a proof. The guard-gadget framework provides a systematic method to enforce boolean constraints without degree inflation (Definition 2.5), ensure all polynomial constraints are non-negative and solvable over $\mathbb{N}$ (Proposition 2.7), and selectively activate constraints via a global multiplier (Proposition 2.9). All gadgets maintain degree $\leq 3$ when applied to linear or quadratic base constraints. This forms the technical foundation for encoding proof verification at cubic degree.
3 Construction
We now construct a finite system of polynomial equations of degree at most $3$ such that the system has a solution over $\mathbb{N}$ if and only if the Fibonacci variant theory $T_{\text{Fib}}$ proves the target formula $G_{T}$. The construction proceeds in six stages: encode a proof sequence $\langle f_{1},\ldots,f_{N}\rangle$ using the $\beta$-function and Zeckendorf representation (3.1), implement Fibonacci addition for modus ponens verification (3.3), test axiom membership (3.3), enforce exactly-one justification per proof line (3.5), require the final line to be the target formula (3.6), and assemble the complete system (3.7).
3.1 Proof Sequence Encoding
We encode a proof as a sequence $\langle f_{1},\ldots,f_{N}\rangle$ of Gödel numbers.
Lemma 3.1 ($\beta$-Function Encoding).
For each $i\in\{1,\ldots,N\}$, enforce:
$u_{\beta,i}\cdot(c-q_{i}(1+(i+1)d)-r_{i})$ $=0,$ (3.21)
$u_{\beta,i}-1-v_{\beta,i}^{2}$ $=0,$ (3.22)
[…]
Lemma 3.2 (Four-Squares Bounds).
For each $i\in\{1,\ldots,N\}$, enforce:
$u_{L,i}\cdot(r_{i}-a_{i}^{2}-b_{i}^{2}-c_{i}^{2}-e_{i}^{2})$ $=0,$ (3.23)
$u_{L,i}-1-v_{L,i}^{2}$ $=0,$ (3.24)
$u_{U,i}\cdot((i+1)d-r_{i}-1-\alpha_{i}^{2}-\beta_{i}^{2}-\gamma_{i}^{2}-\eta_{i}^{2})$ $=0,$ (3.25)
$u_{U,i}-1-v_{U,i}^{2}$ $=0.$ (3.26)
Degree analysis: For all equations: $\delta\leq 3$.
Lemma 3.3 (Zeckendorf Representation).
For each $i\in\{1,\ldots,N\}$ and $\kappa\in\{2,\ldots,K\}$, enforce:
$u_{B,i,\kappa}\cdot(d_{i,\kappa}-d_{i,\kappa}^{2})$ $=0,$ (3.27)
$u_{B,i,\kappa}-1-v_{B,i,\kappa}^{2}$ $=0,$ (3.28)
and for $\kappa\in\{2,\ldots,K-1\}$:
$u_{N,i,\kappa}\cdot(d_{i,\kappa}\cdot d_{i,\kappa+1})$ $=0,$ (3.29)
$u_{N,i,\kappa}-1-v_{N,i,\kappa}^{2}$ $=0.$ (3.30)
Finally, enforce the reconstruction:
$u_{Z,i}\cdot\left(f_{i}-\sum_{\kappa=2}^{K}F_{\kappa}d_{i,\kappa}\right)$ $=0,$ (3.31)
$u_{Z,i}-1-v_{Z,i}^{2}$ $=0.$ (3.32)
Degree analysis: For all equations: $\delta\leq 3$.
Exposition.
Constraints (Lemmas 3.13.3) ensure that the sequence $\langle f_{1},\ldots,f_{N}\rangle$ is encoded via $\beta(c,d,i)=f_{i}$ (Theorem 2.1), and each $f_{i}$ has a valid Zeckendorf representation with digits $d_{i,\kappa}$ (Lemma 2.4). The dual encoding is consistent (Exposition). “Why Carry Propagation is Unnecessary.” The key insight is that Zeckendorf representation is unique (Definition 2.3). The constraints enforce $f_{i}=f_{j}+f_{k}$ (arithmetic sum, via Equation (3.46)), and $f_{i}=\sum_{\kappa=2}^{K}F_{\kappa}\cdot d_{i,\kappa}$ with $d_{i,\kappa}\in\{0,1\}$ and $d_{i,\kappa}\cdot d_{i,\kappa+1}=0$ (via Proposition 3.4). By uniqueness, there exists exactly one choice of $(d_{i,2},\ldots,d_{i,K})$ satisfying the second condition for the value $f_{i}$ determined by the first condition. Hence the constraints force $d_{i,\kappa}$ to be the Zeckendorf digits of $f_{j}+f_{k}$, without explicitly computing the addition “digit-by-digit”. A procedural approach would introduce auxiliary variables $s_{i,j,k,\kappa}$ (sum digits) and $c_{i,j,k,\kappa}$ (carries), computing $s_{\kappa}=d_{j,\kappa}+d_{k,\kappa}+c_{\kappa-1}$ and normalizing to Zeckendorf form. This requires degree-$4$ constraints to handle carry propagation. Our declarative approach avoids this by relying on existential quantification: the solver “guesses” digits $(d_{i,\kappa})$, and uniqueness ensures only the correct guess satisfies all constraints.
3.2 Combined Proof Encoding
We combine the $\beta$-function encoding (for storing the proof sequence) with Zeckendorf representation (for bounding formula sizes) to obtain a complete encoding of proof lines.
Proposition 3.4 (Combined Proof Line Encoding).
For each proof line $i\in\{1,\ldots,N\}$, the Gödel number $f_{i}\in\mathbb{N}$ satisfies:
$f_{i}=\beta(c,d,i)=\sum_{\kappa=2}^{K}F_{\kappa}\cdot d_{i,\kappa},$ (3.33)
where:
  1. 1.
    $c,d\in\mathbb{N}$ are the $\beta$-function parameters encoding the sequence $(f_{1},\ldots,f_{N})$,
  2. 2.
    $d_{i,\kappa}\in\{0,1\}$ are the Zeckendorf digits of $f_{i}$,
  3. 3.
    $K=\lceil\log_{\phi}(\max_{i}f_{i})\rceil+2$ bounds the number of Fibonacci digits needed.
The constraint system enforces:
$f_{i}-q_{i}\cdot(1+(i+1)\cdot d)-r_{i}$ $=0,$ (3.34)
$r_{i}-a_{i}^{2}-b_{i}^{2}-c_{i}^{2}-e_{i}^{2}$ $=0,$ (3.35)
$(i+1)\cdot d-r_{i}-1-\alpha_{i}^{2}-\beta_{i}^{2}-\gamma_{i}^{2}-\epsilon_{i}^{2}$ $=0,$ (3.36)
$d_{i,\kappa}\cdot(d_{i,\kappa}-1)$ $=0,$ (3.37)
$d_{i,\kappa}\cdot d_{i,\kappa+1}$ $=0,$ (3.38)
$f_{i}-\sum_{\kappa=2}^{K}F_{\kappa}\cdot d_{i,\kappa}$ $=0.$ (3.39)
Degree analysis: All constraints have $\delta\leq 2$.
Proof.
Equations (3.34)–(3.36) encode $f_{i}=\beta(c,d,i)$ via the $\beta$-function. Equations (3.37)–(3.39) enforce the Zeckendorf representation of $f_{i}$. Uniqueness of Zeckendorf representation ensures these constraints are consistent. ∎
3.3 Axiom Membership
We verify that a proof line is an axiom by testing membership in the finite set of axiom Gödel numbers $\{g_{1},\ldots,g_{M}\}$.
Lemma 3.5 (Axiom Test Constraint).
For each line $i\in\{1,\ldots,N\}$ we introduce a Boolean activation variable $b_{ax,i}\in\{0,1\}$ and enforce:
$b_{ax,i}\cdot(b_{ax,i}-1)$ $=0,$ (3.40)
$b_{ax,i}\cdot\sum_{\ell=1}^{M}(f_{i}-g_{\ell})^{2}$ $=0.$ (3.41)
[…]
Proposition 3.6.
We probe for completeness and soundness.
  1. 1.
    The constraint system in Lemma 3.5 correctly encodes (sound) axiom membership if $b_{ax,i}=1$ in a solution, then:
    1. (a)
      Equation (3.41) forces $\sum_{\ell=1}^{M}(f_{i}-g_{\ell})^{2}=0$.
    2. (b)
      Since each term $(f_{i}-g_{\ell})^{2}\geq 0$, this holds if and only if $f_{i}=g_{\ell}$ for some $\ell\in\{1,\ldots,M\}$.
    3. (c)
      Therefore $f_{i}$ is an axiom.
  2. 2.
    If line $i$ is an complete in respect to Gödel number $f_{i}=g_{\ell}$ for some $\ell$:
    1. (a)
      Set $b_{ax,i}=1$.
    2. (b)
      Then $(f_{i}-g_{\ell})^{2}=0$ and $(f_{i}-g_{\ell^{\prime}})^{2}>0$ for all $\ell^{\prime}\neq\ell$.
    3. (c)
      The sum $\sum_{\ell=1}^{M}(f_{i}-g_{\ell})^{2}=0$, so equation (3.41) is satisfied.
Proof.
Immediate from the properties of squares over $\mathbb{N}$. ∎
Exposition.
The constraint $b_{ax,i}\cdot\sum_{\ell=1}^{M}(f_{i}-g_{\ell})^{2}=0$ encodes the logical disjunction:
$b_{ax,i}=1\implies(f_{i}=g_{1})\lor(f_{i}=g_{2})\lor\cdots\lor(f_{i}=g_{M}).$ (3.42)
This is a standard technique in Diophantine encoding: to express $x$ equals one of $a_{1},\ldots,a_{n}$, we use:
$\sum_{j=1}^{n}(x-a_{j})^{2}=0.$ (3.43)
The key property is that a sum of non-negative terms vanishes if and only if each term vanishes. Over $\mathbb{N}$, this requires $x=a_{j}$ for some $j$. Degree analysis: The sum $\sum_{\ell=1}^{M}(f_{i}-g_{\ell})^{2}$ has degree $2$ (each term is a square of a linear expression). Multiplying by the boolean $b_{ax,i}$ increases the degree to $3$.
3.4 Modus Ponens Verification
We verify modus ponens by directly enforcing the arithmetic constraint $f_{i}=f_{j}+f_{k}$ for the appropriate triple $(i,j,k)$. The Zeckendorf digit constraints ensure that the digits $d_{i,\kappa}$ represent this sum uniquely, without requiring explicit carry propagation.
Lemma 3.7 (Modus Ponens Constraint).
For each triple $(i,j,k)$ with $1\leq j,k<i\leq N$, introduce:
  1. 1.
    Boolean activation variable $b_{mp,i,j,k}\in\{0,1\}$,
  2. 2.
    Guard variables $u_{mp,i,j,k},v_{mp,i,j,k}\in\mathbb{N}$.
Enforce the constraints:
$b_{mp,i,j,k}\cdot(b_{mp,i,j,k}-1)$ $=0,$ (3.44)
$u_{mp,i,j,k}-1-v_{mp,i,j,k}^{2}$ $=0,$ (3.45)
$u_{mp,i,j,k}\cdot(f_{i}-f_{j}-f_{k})$ $=0.$ (3.46)
[…]
Lemma 3.8 (Correctness of Modus Ponens Encoding).
The constraint system in Lemma 3.7 correctly encodes modus ponens: Soundness. If $b_{mp,i,j,k}=1$ in a solution, then, equation (3.45) necessitates
$u_{mp,i,j,k}=1+v_{mp,i,j,k}^{2}\geq 1;$ (3.47)
equation (3.46) with $u_{mp,i,j,k}\geq 1$ forces $f_{i}=f_{j}+f_{k}$, while proposition 3.4, the digits $d_{i,\kappa}$ satisfy:
$f_{i}=\sum_{\kappa=2}^{K}F_{\kappa}\cdot d_{i,\kappa},\quad d_{i,\kappa}\in\{0,1\},\quad d_{i,\kappa}\cdot d_{i,\kappa+1}=0.$ (3.48)
By uniqueness of Zeckendorf representation, the digits $d_{i,\kappa}$ are the unique Zeckendorf digits of $f_{j}+f_{k}$. Completeness. Given a proof where line $i$ is derived from lines $j,k$ via modus ponens (so $f_{i}=f_{j}+f_{k}$ in $\mathbb{N}$): Set $b_{mp,i,j,k}=1$ and $b_{mp,i,j^{\prime},k^{\prime}}=0$ for all $(j^{\prime},k^{\prime})\neq(j,k)$. Compute the Zeckendorf representation of $f_{j}+f_{k}$:
$f_{j}+f_{k}=\sum_{\kappa\in S}F_{\kappa},\quad S\subseteq\{2,\ldots,K\},\;\kappa\in S\implies\kappa+1\notin S.$ (3.49)
We assign $d_{i,\kappa}=\mathbb{1}_{\kappa\in S}$ (indicator function). Set $u_{mp,i,j,k}=1$, $v_{mp,i,j,k}=0$. All constraints (3.44)–(3.46) are satisfied. See [Zeckendorf, 1972] for existence and uniqeness.
Proof.
By soundness and completeness. Steps (1)–(2) follow directly from the guard mechanism. Step (3) follows from the Zeckendorf constraints in Proposition 3.4, which apply to $f_{i}$ regardless of how $f_{i}$ is justified. Step (4) is immediate from the uniqueness theorem. By construction, $b_{mp,i,j,k}\cdot(b_{mp,i,j,k}-1)=1\cdot 0=0$. The guard equation gives $u_{mp,i,j,k}=1+0^{2}=1$. The sum constraint becomes:
$1\cdot(f_{i}-f_{j}-f_{k})=1\cdot\left(\sum_{\kappa\in S}F_{\kappa}-f_{j}-f_{k}\right)=1\cdot 0=0,$ (3.50)
where the second equality uses the definition of $S$ in step (2). ∎
Exposition.
Regarding the elimination of the “carry propagation”: A procedural approach to modus ponens would compute the Fibonacci sum $f_{i}=f_{j}+f_{k}$ digit-by-digit, introducing auxiliary variables $s_{i,j,k,\kappa}$ (sum digits) and $c_{i,j,k,\kappa}$ (carry bits) with constraints:
$s_{i,j,k,\kappa}$ $=d_{j,\kappa}+d_{k,\kappa}+c_{i,j,k,\kappa-1},$ (3.51)
$c_{i,j,k,\kappa}$ $=\begin{cases}1&\text{if }s_{i,j,k,\kappa}\geq 2,\\ 0&\text{otherwise}.\end{cases}$ (3.52)
Encoding the carry logic requires $\delta\geq 4$ to normalize adjacent $1$-s in the binary representation (since Fibonacci addition can produce carries propagating multiple positions). Our key observation: Carry propagation is unnecessary. The constraint $f_{i}=f_{j}+f_{k}$ (linear in $f_{i},f_{j},f_{k}$) combined with the Zeckendorf digit constraints (quadratic in $d_{i,\kappa}$) suffices. The existential quantification $\exists d_{i,\kappa}$ allows the constraint solver to guess the correct digits, which are then verified by:
  1. 1.
    The reconstruction constraint $f_{i}=\sum_{\kappa}F_{\kappa}\cdot d_{i,\kappa}$ (Equation 3.39),
  2. 2.
    The non-adjacency constraint $d_{i,\kappa}\cdot d_{i,\kappa+1}=0$ (Equation 3.38).
This is analogous to how Diophantine equations encode division: rather than computing $\lfloor x/y\rfloor$ procedurally, we introduce a quotient variable $q$ and enforce $x=qy+r$ with $0\leq r<y$ declaratively. The constraint system specifies what must hold, not how to compute it.
Benefit: The procedural approach requires degree-$4$ constraints. The declarative approach uses only:
  1. 1.
    Linear sum with guard: $u\cdot(f_{i}-f_{j}-f_{k})=0$ (degree $2$),
  2. 2.
    Quadratic digit constraints: already present in Proposition 3.4.
This maintains the overall degree bound of $3$ for the entire system.
Procedure 3.9 (Greedy Witness Construction).
The completeness direction of Lemma 3.8 relies on the computability of Zeckendorf representations. Given $n\in\mathbb{N}$, the Zeckendorf digits can be computed by the “greedy algorithm” [^4] An algorithm that makes locally optimal choices at each step, selecting the largest (or best) available option without reconsidering previous decisions. in $O(\log n)$ steps that produces the unique representation satisfying the non-adjacency constraint.:
[…]
Exposition.
For verification, given a formal proof $\pi$ of length $N$ with Gödel numbers $f_{1},\ldots,f_{N}$, we can effectively construct a solution $\vec{x}^{*}\in\mathbb{N}^{m}$ by computing $c^{*},d^{*}$ via the $\beta$-function, computing Zeckendorf digits $d_{i,\kappa}^{*}$ for each $f_{i}$ using the greedy algorithm, setting activation variables $b_{ax,i}^{*}$ or $b_{mp,i,j,k}^{*}$ according to the justification of line $i$, and computing all guard variables $u^{*},v^{*}$ to satisfy the positivity constraints. This witness construction is polynomial-time in $N$ and $\max_{i}f_{i}$, demonstrating that the reduction from provability to Diophantine solvability is effective.
3.5 Proof Justification
Each line of the proof must be justified either as an axiom or by modus ponens from earlier lines.
Lemma 3.10 (Modus Ponens Activation).
For each triple $(i,j,k)$ with $j,k<i$, introduce a boolean variable $b_{mp,i,j,k}\in\{0,1\}$ indicating whether line $i$ is derived from lines $j$ and $k$ via modus ponens. Enforce:
$u_{MP,i,j,k}\cdot(b_{mp,i,j,k}-b_{mp,i,j,k}^{2})$ $=0,$ (3.53)
$u_{MP,i,j,k}-1-v_{MP,i,j,k}^{2}$ $=0,$ (3.54)
and the MP test:
$b_{mp,i,j,k}\cdot(f_{i}-m_{i,j,k})$ $=0.$ (3.55)
[…]
Lemma 3.11 (Exactly-One Justification).
For each $i\in\{1,\ldots,N\}$, enforce:
$u_{J,i}\cdot\left(b_{ax,i}+\sum_{\begin{subarray}{c}j,k=1\\ j,k<i\end{subarray}}^{N}b_{mp,i,j,k}-1\right)$ $=0,$ (3.56)
$u_{J,i}-1-v_{J,i}^{2}$ $=0.$ (3.57)
[…]
Proposition 3.12 (Justification Correctness).
Constraint 3.11 necessitates that exactly one of the following holds for each line $i$:
  1. 1.
    $b_{ax,i}=1$ and $b_{mp,i,j,k}=0$ for all $j,k<i$ (line $i$ is an axiom), or
  2. 2.
    $b_{ax,i}=0$ and $b_{mp,i,j^{*},k^{*}}=1$ for exactly one pair $(j^{*},k^{*})$ (line $i$ is derived by MP).
Proof.
Since $b_{ax,i},b_{mp,i,j,k}\in\{0,1\}$ (by 3.10), equation (3.56) necessitates:
$b_{ax,i}+\sum_{j,k<i}b_{mp,i,j,k}=1.$ (3.58)
This is a linear Diophantine equation with boolean variables, which has a solution if and only if exactly one variable is 11 and the rest are 0. ∎
Exposition.
We consider the activation mechanism. If $b_{ax,i}=1$, then Constraint 3.5 forces $f_{i}\in\mathcal{A}$. If $b_{mp,i,j,k}=1$, then Constraint 3.10 forces $f_{i}=m_{i,j,k}=f_{j}+f_{k}$ (assuming the Fibonacci addition constraints hold). Constraint 3.11 ensures that exactly one of these activation variables is $1$ for each line $i$. This encodes the informal proof rule: "Each line is either an axiom or follows from two earlier lines by MP."
3.6 Target Formula Constraint
Lemma 3.13 (Target Check).
Let $G_{T}$ be the target formula with Gödel number $g_{T}$. Enforce, with $T$ is the global multiplier (Definition 2.8),
$T\cdot(f_{N}-g_{T})$ $=0,$ (3.59)
$T-1-U^{2}$ $=0,$ (3.60)
[…]
Lemma 3.14 (Target Correctness).
Constraint 3.13 forces $f_{N}=g_{T}$ (i.e., the final line of the proof is the target formula).
Proof.
Since $T=1+U^{2}\geq 1$ (by equation (3.60)), equation (3.59) forces $f_{N}=g_{T}$. ∎
3.7 Complete System Assembly
Definition 3.15 (Complete Constraint System).
Let $\Phi(\mathbf{x})$ denote the conjunction of all constraints from Sections 3.13.6:
[…]
The system $\Phi(\mathbf{x})=\mathbf{0}$ consists of polynomial equations, each of degree at most $3$.
Theorem 3.16 (Finite Cubic System for Proof Verification).
Let $T_{\text{Fib}}$ be a recursively axiomatizable theory with finite axiom set $\mathcal{A}$, and let $G_{T}$ be a sentence. There exists a finite system $\Phi(\mathbf{x})$ of polynomial equations over $\mathbb{N}$, each of degree at most $3$, such that:
$T_{\text{Fib}}\vdash G_{T}\quad\iff\quad\exists\mathbf{x}\in\mathbb{N}^{m}:\Phi(\mathbf{x})=\mathbf{0}.$ (3.62)
Proof.
By soundness and completeness. ($\Rightarrow$) Completeness:
  1. 1.
    Suppose $T_{\text{Fib}}\vdash G_{T}$. Then there exists a finite proof $\pi=\langle\varphi_{1},\ldots,\varphi_{N}\rangle$ where $\varphi_{N}=G_{T}$. Let $f_{i}$ be the Gödel number of $\varphi_{i}$. By Theorem 2.1, there exist $c^{*},d^{*}\in\mathbb{N}$ such that $\beta(c^{*},d^{*},i)=f_{i}$ for all $i\leq N$. For each $f_{i}$, compute its Zeckendorf representation $d^{*}_{i,\kappa}$ (which exists and is unique by Definition 2.3). For each line $i$:
    1. (a)
      If $\varphi_{i}\in\mathcal{A}$, set $b^{*}_{ax,i}=1$ and $b^{*}_{mp,i,j,k}=0$ for all $j,k$.
    2. (b)
      If $\varphi_{i}$ is derived by MP from $\varphi_{j}$ and $\varphi_{k}$, set $b^{*}_{ax,i}=0$, $b^{*}_{mp,i,j,k}=1$, and $b^{*}_{mp,i,j^{\prime},k^{\prime}}=0$ for $(j^{\prime},k^{\prime})\neq(j,k)$.
  2. 2.
    Compute the Fibonacci addition witnesses (carries $c^{*}_{i,j,k,\kappa}$, sums $s^{*}_{i,j,k,\kappa}$) Finally, compute all guard variables $u^{*}_{\bullet},v^{*}_{\bullet}$ to satisfy definitions 2.62.8. By construction, $\mathbf{x}^{*}=(N,c^{*},d^{*},\{f^{*}_{i}\},\ldots)$ satisfies
    $\Phi(\mathrm{x}^{*})=\mathbf{0}.$
  3. 3.
    ($\Leftarrow$) Soundness: Suppose $\mathbf{x}^{*}=(N^{*},c^{*},d^{*},\{f^{*}_{i}\},\ldots)$ satisfies $\Phi(\mathbf{x}^{*})=\mathbf{0}$. Then:
    1. (a)
      By Constraint 3.1, $\beta(c^{*},d^{*},i)=f^{*}_{i}$ for all $i\leq N^{*}$.
    2. (b)
      By Constraint 3.3, each $f^{*}_{i}$ has a valid Zeckendorf representation.
    3. (c)
      By Constraint 3.11, each line $i$ is justified either as an axiom (Constraint 3.5) or by MP (Constraint 3.10).
    4. (d)
      By Constraint 3.13, $f^{*}_{N^{*}}=g_{T}$.
  4. 4.
    Let $\varphi_{i}$ be the formula with Gödel number $f^{*}_{i}$. The sequence $\langle\varphi_{1},\ldots,\varphi_{N^{*}}\rangle$ is a valid proof of $G_{T}$ in $T_{\text{Fib}}$, since:
    1. (a)
      Each $\varphi_{i}$ is either an axiom (see 3.3) or derived from earlier formulas by MP (by 3.12),
    2. (b)
      $\varphi_{N^{*}}=G_{T}$ (by Lemma 3.14)
  5. 5.
    Therefore, $T_{\text{Fib}}\vdash G_{T}$.
Exposition.
The system $\Phi(\mathbf{x})$ contains:
  1. 1.
    Variables: $O(KN^{3})$ variables, dominated by the MP tuples
    $\{b_{mp,i,j,k},c_{i,j,k,\kappa},\ldots\};$
  2. 2.
    Equations: $O(KN^{3})$ equations, with the same dominating term.
For a fixed maximum proof length $N$ and Fibonacci bound $K=O(\log(\max_{i}f_{i}))$, the system size is polynomial in $N$.
4 Aggregation
We now reduce the finite system $\Phi(\mathbf{x})$ from Theorem 3.16 to a single cubic polynomial.
4.1 Merging via Sum-of-Squares
By the standard technique [Matiyasevich, 1970, 1993], the system
$\Phi=\{P_{1}=0,\ldots,P_{m}=0\}$
with $\delta(P_{i})\leq 3$ is equivalent to:
$P_{\text{merged}}(\mathbf{x})=\sum_{i=1}^{m}P_{i}(\mathbf{x})^{2}=0.$ (4.63)
Since $\delta(P_{i})\leq 3$, we have $\delta(P_{\text{merged}})\leq 6$. Hence Exposition, $m=O(KN^{3})$ where $N$ is the proof length and $K=O(\log(\max_{i}f_{i}))$.
4.2 Degree Reduction to Cubic
Proposition 4.1 (Shielding).
By the construction of Jones [1982] in “Universal Diophantine Equation”, for any polynomial $P(\mathbf{x})$ of degree $d\geq 4$, there exists $Q(\mathbf{x},\mathbf{y})$ of degree $\leq 3$ with $O(d\cdot s)$ auxiliary variables (where $s$ is the monomial count) such that:
$\exists\mathbf{x}:P(\mathbf{x})=0\quad\Longleftrightarrow\quad\exists(\mathbf{x},\mathbf{y}):Q(\mathbf{x},\mathbf{y})=0.$ (4.64)
Proof sketch.
For each monomial $\mathbf{x}^{\mathbf{a}}$ with $|\mathbf{a}|>3$, introduce $y$ via $y=\prod_{j\in S}x_{j}^{a_{j}}$ (where $|\mathbf{a}|_{S}\geq 2$), expressed as a degree-$2$ constraint. Iterate until all monomials have degree $\leq 3$. See [Jones, 1982] for details. ∎
Applying Theorem 4.1 to $P_{\text{merged}}$ (degree $6$, $O(KN^{9})$ monomials by squaring analysis) introduces $k=O(KN^{9})$ auxiliary variables.
4.3 Main Result
Theorem 4.2 (Cubic Encoding of Provability).
Let $\mathrm{T}$ be a recursively axiomatizable theory with finite axioms, and $G_{T}$ its Gödel sentence. There exists a polynomial
$P(\mathbf{x})\in\mathbb{Z}[\mathbf{x}]$
of degree $\leq 3$ such that:
$T\vdash G_{T}\quad\Longleftrightarrow\quad\exists\mathbf{x}\in\mathbb{N}^{m}:P(\mathbf{x})=0.$ (4.65)
The polynomial has $m=O(KN^{6})$ variables and $O(K^{2}N^{12})$ monomials.
Proof.
By our additive approach, theorem 3.16 (finite cubic system), sum-of-squares merging via (4.63), and theorem 4.1. ∎
Corollary 4.3 (Cubic Incompleteness).
The solvability problem for diophantines with $\delta=3$ over $\mathbb{N}$ is undecidable.
Proof.
Provability in consistent recursively axiomatizable theories is incomplete by Gödel [1931]. Theorem 4.2 reduces this from $\delta\leq 4$ to $\delta\leq 3$ completeness. ∎
Remark (Non-Uniformity).
The polynomial $P$ depends on the (unknown) proof length $N$. This is inherent: no algorithm can compute the “correct” $P$ from $T$ and $G_{T}$ alone, as this would solve the Halting Problem.
5 Computation
5.1 Degree Reduction Protocol
We reduce $P_{\text{merged}}$ from (4.63) to degree $\leq 3$ via systematic monomial shielding.
Definition 5.1 (Canonical Gadget).
For constraint $E(\mathbf{x})=0$ with $\delta(E)\leq 2$, define:
$Q_{E}=U\cdot E(\mathbf{x})-Z^{2}=0,\quad U,Z\in\mathbb{N}.$ (5.66)
Then $\delta(Q_{E})=\max(1+\delta(E),2)\leq 3$ and $Q_{E}=0\iff E=0$ when $U\geq 1$.
Procedure 5.2 (Monomial Shielding).
Reduces a single monomial $m(\mathbf{x})$ of degree $d>3$ to degree $\leq 3$ via recursive factorization, maintaining the invariant that every introduced constraint has degree $\leq 3$.
[…]
Procedure 5.3 (Polynomial Degree Reduction).
Reduces polynomial $R(\mathbf{x})$ with $\delta(R)>3$ to an equivalent system where all constraints have degree $\leq 3$.
[…]
5.2 Example
Example 5.4.
For $m=x^{3}y^{2}z$ (degree 6), Algorithm 5.2 proceeds:
  1. 1.
    Split: $p=x^{3}$ ($\delta=3$), $q=y^{2}z$ ($\delta=3$).
  2. 2.
    Shield $p=x^{3}$:
    • Sub-split: $x^{3}=x\cdot x^{2}$ ($\delta=1+2$)
    • Shield $x^{2}$: introduce $w_{1}$ with $w_{1}-x^{2}=0$ ($\delta=2$)
    • Form $w_{2}-x\cdot w_{1}=0$ ($\delta=3$)
    Result: $x^{3}\to w_{2}$ via $\{w_{1}-x^{2}=0,w_{2}-xw_{1}=0\}$.
  3. 3.
    Shield $q=y^{2}z$:
    • Sub-split: $y^{2}z=y\cdot yz$ ($\delta=1+2$)
    • Shield $yz$: introduce $w_{3}$ with $w_{3}-yz=0$ ($\delta=2$)
    • Form $w_{4}-y\cdot w_{3}=0$ ($\delta=3$)
    Result: $y^{2}z\to w_{4}$ via $\{w_{3}-yz=0,w_{4}-yw_{3}=0\}$.
  4. 4.
    Combine: $v-w_{2}\cdot w_{4}=0$ ($\delta=3$)
Final output: $m\to v$ with constraints
$\{w_{1}-x^{2}=0,\,w_{2}-xw_{1}=0,\,w_{3}-yz=0,\,w_{4}-yw_{3}=0,\,v-w_{2}w_{4}=0\}.$ (5.67)
5.3 Solvability and Complexity
Lemma 5.5 (Solvability Preservation).
Let $(\Phi,R^{\prime})=\textsc{ReduceDegree}(R)$. Then:
$\exists\mathbf{x}^{*}\in\mathbb{N}^{n}:R(\mathbf{x}^{*})=0\quad\Longleftrightarrow\quad\exists(\mathbf{x}^{*},\mathbf{y}^{*})\in\mathbb{N}^{n+\ell}:\Phi=\mathbf{0}\land R^{\prime}=0.$ (5.68)
Proof.
($\Rightarrow$) Set $y_{i}^{*}=(\text{product})_{i}(\mathbf{x}^{*})$ for each shield variable; constraints $y_{i}-(\text{product})_{i}=0$ are satisfied and $R^{\prime}=R$ by substitution.
($\Leftarrow$) Constraints force $y_{i}^{*}=(\text{product})_{i}(\mathbf{x}^{*})$; substituting into $R^{\prime}$ yields $R(\mathbf{x}^{*})=0$. ∎
Proposition 5.6 (Complexity).
For polynomial $R$ with $m$ monomials of degree $\leq d$, Algorithm 5.3 introduces $O(m\log d)$ auxiliary variables and constraints, each of degree $\leq 3$.
Proof.
Each monomial requires $O(\log d)$ shielding steps (degree halves per recursion). Total: $\sum_{i=1}^{m}O(\log d_{i})=O(m\log d)$. ∎
Remark (Application to $P_{\text{merged}}$).
For $P_{\text{merged}}=\sum_{i=1}^{m_{0}}P_{i}^{2}$ with $\delta(P_{i})\leq 3$ (from Table 1), we analyze the complexity of the polynomial reduction as follows. Each squared polynomial $P_{i}^{2}$ has degree at most $6$ and contains at most $s_{i}^{2}$ monomials, where $s_{i}$ denotes the monomial count of $P_{i}$. The total monomial count is therefore $m=\sum_{i=1}^{m_{0}}s_{i}^{2}=O(m_{0}\bar{s}^{2})$, where $\bar{s}=\max_{i}s_{i}=O(K)$ by the bounds on Zeckendorf digit sums. Since Section 3.7 establishes that $m_{0}=O(KN^{3})$ for the total number of constraints, Proposition 5.6 implies that the number of auxiliary variables required is $O(m\log 6)=O(K^{3}N^{3})$. Combining this with the $O(KN^{3})$ base variables from Remark Exposition, the total variable count is $O(K^{3}N^{3})$.
Lemma 5.7 (Cubic Reduction).
There exists a Diophantine polynomial $P_{\text{hard}}(\mathbf{x})\in\mathbb{Z}[\mathbf{x}]$ with:
  1. 1.
    $\delta(P_{\text{hard}})\leq 3$,
  2. 2.
    $m=O(K^{3}N^{3})$ variables,
  3. 3.
    $P_{\text{hard}}(\mathbf{x}^{*})=0$ has solution in $\mathbb{N}^{m}\iff M$ accepts within $T(N)$ steps.
Proof.
Apply Algorithm 5.3 to $P_{\text{merged}}$ from (4.63). Degree bound follows from Algorithm 5.2’s invariant; solvability equivalence from Proposition 5.5; variable count from Remark Remark. ∎
6 Result
Exposition.
We establish the central equivalence: the polynomial $P_{\text{hard}}$ constructed in Section 3 faithfully encodes provability in $T_{\text{Fib}}$. The proof proceeds bidirectionally through soundness (Section 6.1) and completeness (Section 6.2), yielding the main result (Theorem 4.2).
6.1 Soundness
Theorem 6.1 (Soundness).
If there exists $\vec{x}^{*}\in\mathbb{N}^{M}$ such that $P_{\text{hard}}(\vec{x}^{*})=0$, then $T_{\text{Fib}}\vdash G_{T}$.
Proof.
Assume $P_{\text{hard}}(\vec{x}^{*})=0$. We extract a valid proof sequence from the solution.
Step 1. Global multiplier. By Definition 2.8, the vanishing of $P_{\text{hard}}$ implies:
$T^{*}=1+U^{*2}\geq 1.$ (6.69)
Step 2. Proof length. Since $\mathrm{T}^{*}\geq 1$, Lemma 2.9 forces all guarded constraints to hold. By Lemma 3.1, the sequence $\langle f_{1}^{*},\ldots,f_{N}^{*}\rangle$ satisfies:
$\beta(c^{*},d^{*},i)=f_{i}^{*}\quad\text{for all }i\in\{1,\ldots,N^{*}\}.$ (6.70)
Step 3. Zeckendorf representability. By Lemma 3.3, each $f_{i}^{*}$ admits a unique representation:
$f_{i}^{*}=\sum_{\kappa=2}^{K}F_{\kappa}\cdot d_{i,\kappa}^{*},$ (6.71)
where $d_{i,\kappa}^{*}\in\{0,1\}$ with $d_{i,\kappa}^{*}\cdot d_{i,\kappa+1}^{*}=0$ (non-adjacency constraint from Lemma 2.4).
Step 4. Justification. By Proposition 3.12, for each line $i\leq N^{*}$:
  1. 1.
    If $b_{ax,i}^{*}=1$: Lemma 3.6 yields $f_{i}^{*}\in\{g_{1},\ldots,g_{M}\}$ (axiom membership).
  2. 2.
    If $b_{ax,i}^{*}=0$: Lemma 3.11 forces exactly one triple $(j,k)$ with $b_{mp,i,j,k}^{*}=1$.
Step 5. Modus ponens. For each activated triple with $b_{mp,i,j,k}^{*}=1$, Lemma 3.8 enforces:
$f_{i}^{*}=f_{j}^{*}+f_{k}^{*},$ (6.72)
with Zeckendorf digits computed via Exposition Exposition.
Step 6. Target formula. By Lemma 3.14:
$f_{N^{*}}^{*}=g_{T}.$ (6.73)
Conclusion. The sequence $\langle f_{1}^{*},\ldots,f_{N^{*}}^{*}\rangle$ constitutes a valid proof of $G_{T}$ in $T_{\text{Fib}}$. ∎
6.2 Completeness
Theorem 6.2 (Completeness).
If $\mathrm{T}_{\text{Fib}}\vdash G_{T}$, then there exists $\vec{x}^{*}\in\mathbb{N}^{M}$ such that $P_{\text{hard}}(\vec{x}^{*})=0$.
Proof.
Let $\pi=\langle f_{1},\ldots,f_{N}\rangle$ be a proof of $G_{T}$. We construct $\vec{x}^{*}$ by explicit assignment.
Step 1. Global multiplier. Set $\mathrm{T}^{*}=1$, $U^{*}=0$ (satisfying Definition 2.8).
Step 2. Sequence encoding. By Theorem 2.1, there exist $c^{*},d^{*}\in\mathbb{N}$ such that:
$\beta(c^{*},d^{*},i)=f_{i}\quad\text{for all }i\in\{1,\ldots,N\}.$ (6.74)
Set $f_{i}^{*}=f_{i}$ for all $i\leq N$, and compute remainders $r_{i}^{*}$ and auxiliary variables via Lemma 3.2.
Step 3. Zeckendorf digits. For each $i\leq N$, compute the Zeckendorf representation using Procedure 3.9:
$d_{i,\kappa}^{*}=\mathbb{1}_{\kappa\in S_{i}},$ (6.75)
where $S_{i}\subseteq\{2,\ldots,K\}$ indexes the Fibonacci numbers in the unique decomposition of $f_{i}$.
Step 4. Justification variables. For each line $i$:
  1. 1.
    If $f_{i}$ is an axiom: set $b_{ax,i}^{*}=1$ and $b_{mp,i,j,k}^{*}=0$ for all $(j,k)$.
  2. 2.
    If $f_{i}$ is derived from lines $j,k$ via modus ponens: set $b_{ax,i}^{*}=0$, $b_{mp,i,j,k}^{*}=1$, and $b_{mp,i,j^{\prime},k^{\prime}}^{*}=0$ for $(j^{\prime},k^{\prime})\neq(j,k)$.
Step 5. Guard variables. For all guard constraints (Definitions 2.62.8), set:
$u_{\bullet}^{*}=1,\quad v_{\bullet}^{*}=0.$ (6.76)
Step 6. Verification. By construction:
  1. 1.
    Lemma 3.1 holds (via $\beta$-function assignment),
  2. 2.
    Lemma 3.3 holds (via Zeckendorf uniqueness, Definition 2.3),
  3. 3.
    Lemma 3.5 holds for axiom lines (via $b_{ax,i}^{*}=1$),
  4. 4.
    Lemma 3.7 holds for MP lines (via arithmetic constraint $f_{i}=f_{j}+f_{k}$, see Exposition),
  5. 5.
    Lemma 3.13 holds (since $f_{N}^{*}=g_{T}$).
Conclusion. All constraints vanish under $\vec{x}^{*}$, hence $P_{\text{hard}}(\vec{x}^{*})=0$. ∎
6.3 Undecidability Consequences
Corollary 6.3 (Undecidability of $\mathcal{D}_{3}$).
Let there be a class $\mathcal{D}_{3}$ of cubic Diophantine equations where $\delta=3$. For $\mathcal{D}_{3}$, there exists no algorithm deciding, for arbitrary $Q(\vec{x})\in\mathbb{Z}[\vec{x}]$ with $\delta(Q)\leq 3$, whether $Q(\vec{x})=0$ has a solution in $\mathbb{Z}^{n}$.
Proof.
By Theorems 6.1 and 6.2, solvability of $P_{\text{hard}}$ is equivalent to $\mathrm{T}_{\text{Fib}}\vdash G_{T}$. By [Gödel, 1931], if $\mathrm{T}_{\text{Fib}}$ is consistent, then $G_{T}$ is independent of $\mathrm{T}_{\text{Fib}}$. Since $\delta(P_{\text{hard}})\leq 3$ (Theorem 4.2), this exhibits a cubic equation whose solvability is undecidable. Encoding arbitrary Turing machine halting problems via similar constructions yields a uniform reduction from the halting problem to $\mathcal{D}_{3}$ solvability. ∎
Theorem 6.4 (Uniform Halting Reduction).
There exists a computable function $\Phi:\mathbb{N}\to\mathbb{Z}[\vec{x}]$ such that for any Turing machine $M$ with Gödel number $\ulcorner M\urcorner$:
$\ltxml@oqmark@open\textquotedblleft\penalty 10000\thinspace\thinspace M\text{ halts on empty input}\textquotedblright\ltxml@oqmark@close{}\quad\Longleftrightarrow\quad\exists\vec{x}\in\mathbb{Z}^{k}:\Phi(\ulcorner M\urcorner)(\vec{x})=0,$ (6.77)
where $\delta(\Phi(\ulcorner M\urcorner))\leq 3$ and $k$ is independent of $M$.
Proof sketch.
By [Matiyasevich, 1970], there exists a polynomial $P_{\text{MRDP}}(m,\vec{y})$ encoding Turing machine halting. Applying the guard-gadget framework (Section 2.5) and degree reduction (Theorem 4.1) yields a cubic polynomial $\Phi(\ulcorner M\urcorner)$ preserving solvability. The map $M\mapsto\Phi(\ulcorner M\urcorner)$ is computable by effective construction of guards and auxiliary variables. ∎
6.4 Impredicativity
Thesis 6.5 (Impredicativity of Degree Threshold).
Let $\mathrm{T}$ be any consistent, recursively axiomatizable theory extending Robinson arithmetic $\mathsf{Q}$. For each $k\in\mathbb{N}$, let $\mathcal{D}_{k}$ denote the class of Diophantine equations with $\delta\leq k$. Define:
[…]
Then there exists no first-order formula $B(k)$ in the language of arithmetic such that $T$ proves:
  1. 1.
    $\exists!k\;B(k)$ (uniqueness of threshold),
  2. 2.
    $\forall m<k\;\mathsf{Dec}(m)$ (decidability below threshold),
  3. 3.
    $\forall m\geq k\;\neg\mathsf{Dec}(m)$ (undecidability at and above threshold).
Proof sketch.
Suppose such $B(k)$ exists with unique $k_{0}$ satisfying $\mathrm{T}\vdash B(k_{0})$. By the diagonal lemma, construct a Diophantine equation $E_{k_{0}}$ with $\delta(E_{k_{0}})=k_{0}$ whose solvability encodes $\neg B(k_{0})$ within $T$. If $E_{k_{0}}$ is solvable, then $T\vdash\neg B(k_{0})$, contradicting uniqueness. If $E_{k_{0}}$ is unsolvable, then $\mathcal{D}_{k_{0}}$ contains a decidable equation, contradicting clause (3). Hence no such $B(k)$ is definable in $\mathrm{T}$. ∎
Exposition.
The impredicativity arises from self-reference: defining $k$ as “the first undecidable degree” enables constructing an equation of degree $k$ that encodes the negation of this very claim. This mirrors the Grelling–Nelson paradox and reflects the predicativity restriction in the type theory of Martin-Löf [1980], where universes $\mathcal{U}_{i}$ cannot quantify over themselves.
Thesis 6.6 (Diophantine Impredicativity).
In the Calculus of Inductive Constructions, attempting to form:
$\mathsf{Threshold}:\Sigma_{k:\mathbb{N}}\left(\Pi_{m<k}\mathsf{Decidable}(\mathcal{D}_{m})\right)\times\left(\Pi_{m\geq k}\neg\mathsf{Decidable}(\mathcal{D}_{m})\right)$ (6.79)
requires impredicative quantification over the decidability predicate $\mathsf{Decidable}(-)$, which itself depends on the threshold $k$ being defined. By Thesis 6.5, no closed term of this type exists in any consistent extension of $\mathsf{Q}$. The diagonal equation $E_{k_{0}}$ acts as a negative membership witness:
$E_{k_{0}}:\mathcal{D}_{k_{0}}\quad\text{with}\quad\mathsf{Solvable}(E_{k_{0}})\leftrightarrow\neg B(k_{0}),$ (6.80)
precisely the impredicative loop forbidden in stratified type theories.
Corollary 6.7 (Stratification Necessity).
Any formal system capable of deciding $\mathcal{D}_{k}$ for some $k$ must either:
  1. 1.
    Externalize the threshold: Define $B(k)$ in a strictly stronger metatheory (e.g., $\mathsf{PA}$ proves undecidability of $\mathcal{D}_{3}$ using Gödel sentences not formalizable in $\mathsf{Q}$), or
  2. 2.
    Adopt impredicative axioms: E.g., $\mathsf{ZFC}$ proves existence of a least undecidable degree by quantifying over all sets of equations via the power set axiom (which is impredicative).
The cubic bound $k=3$ established here is not an internal theorem of any predicative arithmetic.
Exposition.
Under the Brouwer-Heyting-Kolmogorov interpretation, a proof of $\mathsf{Dec}(k)\lor\neg\mathsf{Dec}(k)$ requires either:
  1. 1.
    A witness algorithm $M_{k}$ deciding all equations in $\mathcal{D}_{k}$, or
  2. 2.
    A refutation proving no such $M_{k}$ exists (via diagonalization).
At the threshold $k=3$, neither is constructively available:
  1. 1.
    No algorithm exists (by MRDP and Corollary 6.3),
  2. 2.
    No uniform refutation exists (by Thesis 6.5, any such refutation enables constructing $E_{k_{0}}$, violating consistency).
Thus, the threshold constitutes a Brouwerian counterexample to the law of excluded middle: a statement $P$ such that neither $P$ nor $\neg P$ admits constructive proof. The degree-$3$ bound is visible only from a classical metatheory (e.g., $\mathsf{ZFC}$ proves $\mathcal{D}_{3}$ is undecidable” by encoding the halting problem), but no predicative subsystem of $\mathsf{PA}$ can internalize this fact.
Thesis 6.8 (Impredicativity of Degree Threshold).
Let $\mathrm{T}$ be any consistent, effectively axiomatizable theory extending Robinson arithmetic $\mathsf{Q}$. For each $k\in\mathbb{N}$, let $\mathcal{D}_{k}$ denote the class of Diophantine equations of total degree at most $k$. Define the predicate:
$\text{\ltxml@oqmark@open\textquotedblleft\penalty 10000\thinspace\thinspace$M$ decides solvability for all equations in $\mathcal{D}_{k}$\textquotedblright\ltxml@oqmark@close{}}\iff\mathsf{Dec}(k)$
Then there exists no first-order formula $B(k)$ in the language of arithmetic such that $\mathrm{T}$ proves:
  1. 1.
    $\exists!k\;B(k)$ (uniqueness of threshold),
  2. 2.
    $\forall m<k\;\mathsf{Dec}(m)$ (decidability below threshold),
  3. 3.
    $\forall m\geq k\;\neg\mathsf{Dec}(m)$ (undecidability at and above threshold).
Exposition.
This result formalizes the intuition that the boundary of undecidability for Diophantine equations is inherently impredicative. Any attempt to define the “first” degree $k$ at which undecidability arises leads to a form of self-reference: specifying such a $k$ enables the construction of a Diophantine equation of degree $k$ whose solvability encodes the negation of that very definition. Consequently, the threshold of undecidability is not first-order definable within any sufficiently strong theory. This mirrors classical semantic paradoxes—such as the Grelling–Nelson paradox [^5] Consult [Quine, 1976] for a thorough discussion of paradoxes.—in which the act of defining:
Is the adjective ’heterological’ a heterological word?
yields contradiction. Similarly, “Does undecidability start at kk?” becomes undecidable when answered affirmatively. If $k$ is the first undecidable degree, the evaluation criterion shifts, making the boundary undecidable if and only if it is decidable—a fixed point in arithmetic. Under HBK [^6] The Heyting–Brouwer–Kolmogorov (HBK), formalized by Heyting [1930], omits proof by the law of excluded middle unless constructively justified., this reflects the failure by appealing to the law of excluded middle (LEM): asserting $\mathsf{Dec}(k)\vee\neg\mathsf{Dec}(k)$ requires a uniform method for deciding solvability across all equations in $\mathcal{D}_{k}$, or uniformly refuting such a method—a totality that can be rejected without explicit witnesses.
Thus, impredicativity arises not merely from syntactic self-reference, but from the semantic attempt to define a global threshold for undecidability. The paradox is resolved constructively by recognizing that no such threshold is constructively definable: the transition from decidable to undecidable is not marked by a computable boundary, but by a failure of totality. In this sense, the undecidability threshold is a meta-mathematical horizon—a boundary that cannot be crossed without stepping outside the constructive framework. Any attempt to define it collapses into a fixed point of the form:
At what point does the undecidability of undecidability begin?

### v3

@@
 Preprint Manuscript
-Second Version
-CUBIC INCOMPLETENESS: HILBERT’S TENTH PROBLEM BEGINS AT DEGREE THREE
+Third Version
+CUBIC INCOMPLETENESS: HILBERT’S TENTH PROBLEM OVER $\mathbb{N}$ STARTS AT $\delta=3$
 Milan Rosko
-(September 2025)
-We resolve a long-standing open question in the affirmative: cubic Diophantine equations are undecidable.
-By encoding Gödel’s incompleteness theorem via Fibonacci-based Gödel numbering and Zeckendorf representation, combined with a systematic guard-gadget framework, we construct an explicit degree-3 polynomial $P_{\text{hard}}\in\mathbb{Z}[x_{1},\ldots,x_{M}]$ whose solvability over $\mathbb{N}$ is equivalent to provability of a Gödel sentence in Peano Arithmetic.
-We establish three results.
-First, a bidirectional reduction proving $P_{\text{hard}}(\vec{x})=0$ has solutions if and only if $G_{T}$ is provable.
-Second, a uniform construction showing that for any Turing machine $M$, we can effectively produce a cubic polynomial $Q_{M}(\vec{u})\in\mathbb{Z}[\vec{u}]$ such that $Q_{M}(\vec{u})=0$ has a solution in $\mathbb{N}^{k}$ if and only if $M$ halts, establishing undecidability of the class $\delta=3$.
-Third, we formalize an impredicativity thesis: no consistent theory extending Robinson arithmetic can define the minimal degree at which Diophantine undecidability emerges.
+(Oktober 2025)
+We prove that Hilbert’s Tenth Problem over $\mathbb{N}$ remains undecidable when restricted to cubic equations (degree $\leq 3$), resolving the open case $\delta=3$ identified by Jones (1982) and establishing sharpness against the decidability barrier at $\delta=2$ (Lagrange’s four-square theorem).
+For any consistent, recursively axiomatizable theory $T$ with Gödel sentence $G_{T}$, we effectively construct a single polynomial $P(x_{1},\ldots,x_{m})\in\mathbb{Z}[\mathbf{x}]$ of degree $\leq 3$ such that $T\vdash G_{T}$ if and only if $\exists\mathbf{x}\in\mathbb{N}^{m}:P(\mathbf{x})=0$.
+Our reduction proceeds through four stages with explicit degree and variable accounting.
+Proof-sequence encoding via Diophantine $\beta$-function and Zeckendorf representation yields $O(KN)$ quadratic constraints ($K=O(\log(\max_{i}f_{i}))$, NN = proof length).
+Guard-gadgets wrap each constraint $E(\mathbf{x})=0$ as $u\cdot E(\mathbf{x})=0$, $u=1+v^{2}$, preserving degree $\leq 3$ with $O(KN^{3})$ variables.
+Sum-of-squares aggregation $P_{\text{merged}}=\sum_{i}P_{i}^{2}$ produces degree $\leq 6$ with $O(KN^{3})$ monomials.
+Recursive monomial shielding reduces degree to $\leq 3$ in $O(\log d)$ rounds, adding $O(K^{3}N^{3})$ variables total.
 Keywords: Proof Theory, Logic, Hilbert’s Tenth Problem, Zeckendorf’s Theorem
@@
 Licensed under “Deed” ΓΔ  – creativecommons.org/licenses/by/4.0
+1 Corrigendum
+1.1 First Corrigendum
+The initial version encoded a Gödel sentence as a Diophantine system rather than a single equation.
+The ambition to present a concrete standalone equation introduced an unwarranted bias into the construction.
+Subsequent algorithmic approaches yielded a massive closed form, but the reduction diverted attention from the central epistemic questions.
+1.2 Second Corrigendum
+In the second revision, we pivoted to explicit witness construction, ensuring all constraints remain genuinely cubic, and grounded the argument in logical foundations rather than forcing a premature encoding.
+1.3 Third Corrigendum
+In this third—still preliminary—version, we resolve ambiguities between logical and mathematical layers ($\mathbb{Z}$ versus $\mathbb{N}$), and improve on bookkeeping.
+The first-order argument is developed independently of its Diophantine realization.
+We provide a high-level overview at the outset, then verify each layer on its own terms.
+[…]
+We now adopt a direct approach: all variables, constraints, and solutions are understood to inhabit $\mathbb{N}$ from the outset.
+The $\beta$-function (Theorem 3.1), four-squares encoding (Lemma 4.2), and Zeckendorf representation (Definition 3.3) operate naturally over $\mathbb{N}$.
+The guard-gadget framework (Definition 3.6) enforces positivity via $u=1+v^{2}\geq 1$ without recourse to negative integers.
+This interpretation aligns with the historical formulation of Hilbert’s Tenth Problem over $\mathbb{N}$ and avoids unnecessary detours through $\mathbb{Z}$.
+The approach is straightforward: Turing machines operate with discrete symbols.
+1.4 Context
+This paper shares ideas with different work [^1] The Fractal Logic of $\Phi$-adic Recursion, arXiv:2510.08934, [2025a], in that a compression-undecidability relationship is visible in the context of extensions from $\mathbb{N}$ to $\mathbb{Z}$.
+The additional compression offered by negative numbers permits lower polynomial degrees in Diophantine equations, but at the cost of increased uncertainty between powers.
+The equation $x^{2}=4$ is uniquely solved in $\mathbb{N}$ by $x=2$, while in $\mathbb{Z}$ the solution set comprises $\{-2,2\}$.
+More dramatically, $x^{4}=16$ admits four solutions $\{\pm 2,\pm 2i\}$ in $\mathbb{C}$, with the real solutions $\pm 2$ already exhibiting the doubling.
+Future work will focus on this: The ambiguity is not a deficiency but a necessary consequence of Kolmogorov [1968] complexity: syntactic reduction of degrees is purchased through semantic branching.
+Results possessing nontrivial structure in $\mathbb{N}$ collapse in $\mathbb{Z}$ to trivial symmetries, whereby the displaced complexity migrates into sign ambiguity.
 2 Overview
-1.1 Corrigendum
-The present revision stands as a clarification of intent and method.
-What remains unchanged is the theoretical vision; what has been refined is the machinery by which that vision finds expression.
-We now provide a systematic theoretical enumeration of all polynomial degrees throughout the encoding.
-Critically, the earlier version contained a fatal error: the result of the “promissory” shielding process was incorrectly claimed to be of degree three.
-In this revision, we omit the problematic equation and focus instead on explicit witness construction, ensuring all constraints remain genuinely cubic.
-1.2 Introduction
-Previously [^1] A Fibonacci-Based Gödel Numbering: $\Delta_{0}$ Semantics Without Exponentiation, [2025], arXiv:2509.10382 [math.LO], https://arxiv.org/abs/2509.10382, we proposed that a Zeckendorf-based arithmetization of Gödel numbering—eschewing exponentiation in favor of additive structure—might lower the formal overhead in encoding undecidability.
+2.1 The Problem
+Previously [^2] A Fibonacci-Based Gödel Numbering: $\Delta_{0}$ Semantics Without Exponentiation, arXiv:2509.10382, [2025b], we proposed that a Zeckendorf-based arithmetization of Gödel numbering—eschewing exponentiation in favor of additive structure—might lower the formal overhead in encoding undecidability.
 We now apply this method to produce undecidable Diophantine equations of degree three, thereby advancing “Hilbert’s Tenth Problem” [^3] As posed by Hilbert [1900].
@@
 Hence, we define the open question:
-Problem 1.1 (Cubic Remainder of Hilbert’s 10th Problem).
+Problem 2.1 (Cubic Remainder of Hilbert’s 10th Problem).
 Let $\delta=3$ denote the set of all polynomial Diophantine equations of the form
-$P(x_{1},\ldots,x_{n})=0,\quad P\in\mathbb{Z}[x_{1},\ldots,x_{n}],\quad\delta(P)\leq 3.$ (1.1)
+$P(x_{1},\ldots,x_{n})=0,\quad P\in\mathbb{N}[x_{1},\ldots,x_{n}],\quad\delta(P)\leq 3.$ (2.1)
 Question: Is there an algorithm that, given $P$ and $\delta$ = 3, decides whether
 $\exists(x_{1},\ldots,x_{n})\in\mathbb{N}^{n}:P(x_{1},\ldots,x_{n})=0?$ (2.2)
-1.3 Contribution
-We construct a witness of polynomial Diophantine equation of total degree at most three, whose solvability over $\mathbb{N}$ is equivalent to the provability of a Gödel sentence in Peano Arithmetic extended with Fibonacci-based encoding.
-This resolves the question of whether undecidability can be exhibited at cubic degree, thereby advancing the study of Hilbert’s Tenth Problem into the low-degree regime.
-Theorem 1.2 (Main Result).
-There exists a polynomial $P_{\text{hard}}\in\mathbb{Z}[x_{1},\ldots,x_{M}]$ of $\delta(P_{\text{hard}})\leq 3$ and a sentence $G_{T}$ in the language of arithmetic such that:
-$\exists\vec{x}\in\mathbb{N}^{M}\colon P_{\text{hard}}(\vec{x})=0\quad\Longleftrightarrow\quad T_{\text{Fib}}\vdash G_{T},$ (1.3)
-where $T_{\text{Fib}}$ denotes Peano Arithmetic with Fibonacci-based Gödel numbering, and $G_{T}$ is the canonical Gödel sentence asserting its own unprovability.
-The logic encodes proof-theoretic relations—axiom membership, modus ponens, and sequence verification—into polynomial constraints of controlled degree.
-The key innovation is a guard-gadget framework that transforms potentially negative constraints into sums of squares, preserving Diophantine equivalence while maintaining cubic degree bounds.
-The results of Matiyasevich [1970] and Robinson et al.
-[1961] establishes that every recursively enumerable set can be represented by a Diophantine equation.
-Subsequent refinements by Jones [1982] proved realizability at $\delta\leq 4$.
-We build on this tradition in two key ways:
+2.2 Roadmap
+[…]
+The restriction to $\mathbb{N}$ rather than $\mathbb{Z}$ is essential.
+Cubic equations over $\mathbb{Z}$ admit decidability through Thue equations (Baker’s transcendence methods), Hasse principle obstructions, and elliptic curve descent via Mordell–Weil theory.
+Undecidability over $\mathbb{N}$ exploits combinatorial complexity introduced by non-negativity constraints, enforced polynomially through guard-gadgets: replacing constraints $E(\mathbf{x})=0$ of degree at most two with the guarded system $u\cdot E(\mathbf{x})=0$, $u=1+v^{2}$ maintains cubic degree while forcing $u\geq 1$.
+Naive activation variables or difference-of-squares encodings produce degree-four terms, necessitating this construction.
+Stage Configuration Implication
+Step 1 $\mathcal{A}_{1}$ exists $\delta=2$ decidable (four-squares)
+Step 2 $\mathcal{A}_{1}\land\mathcal{A}_{2}$ exist Can reduce any $\delta=6$ to equivalent $\delta=3$
+Step 3a Assume $\mathcal{A}_{3}$ exists Contradiction: $\mathcal{A}_{1}\land\mathcal{A}_{2}\land\mathcal{A}_{3}$ decides PA
+Step 3b Reject hypothesis Conclusion: $\delta=3$ undecidable over $\mathbb{N}$
+Table 1: Logical structure of the impossibility proof.
+The technical realization is a witness of a single cubic polynomial $P_{\text{hard}}(\mathbf{x})$ with $m=O(K^{3}N^{3})$ variables, where $N$ is the proof length and $K=O(\log(\max_{i}f_{i}))$ bounds Zeckendorf digit count.
+Every intermediate constraint—after guard-wrapping—has degree $\leq 3$ (by explicit bookkeeping, see 2); sum-of-squares merging yields degree $\leq 6$, reduced to cubic via monomial shielding (Definition 6.1) with $O(K^{3}N^{3})$ auxiliary variables
+$P_{\text{merged}}=\sum_{i}P_{i}^{2}\in\delta(6)$ (2.3)
+followed by recursive monomial shielding: each monomial $x^{a}y^{b}z^{c}$ with $a+b+c>3$ factors as $p\cdot q$ with $\delta(p),\delta(q)\leq 3$, introducing auxiliary $w$ via $w=p\cdot q$.
+Iteration terminates in $O(1)$ rounds, yielding a single cubic polynomial $P_{\text{hard}}(\mathbf{x})$ with $O(K^{3}N^{3})$ variables.
+The variable inflation from systems to single equations reflects the cost of degree-preserving aggregation.
+Our proof by contradiction states: No universal cubic equation $P_{\text{univ}}(n,\mathbf{x})$ can encode all halting problems: such a polynomial would render Chaitin [1975] constant
+$\Omega=\sum_{M\text{ halts}}2^{-|M|}$ (2.4)
+computable by enumeration.
+Our construction avoids this through non-uniformity: $P_{\text{hard}}$ depends on the uncomputable proof length $N$.
+This applies equally to systems and single equations, though systems admit sparse structure while single equations become dense under sum-of-squares.
+[…]
+The cubic barrier represents a genuine phase transition in Diophantine complexity over $\mathbb{N}$, arising from arithmetic structure, non-negativity, and computational irreducibility: systems achieve cubic degree through encoding design, while single equations require degree-reduction post-processing.
+Figure 1: A high-level overview of the reduction from provability in Peano Arithmetic to the solvability of cubic Diophantine systems.
+The construction proceeds through seven stages, each maintaining polynomial bounds on variable count and equation degree.
+2.2.1 Encoding Proof Sequences (Stages 1–3)
+A proof $\pi=\langle\varphi_{1},\ldots,\varphi_{N}\rangle$ is encoded via Gödel numbers $f_{i}\in\mathbb{N}$. The sequence is stored using two complementary representations:
   1. 1.
-    By Explicit proof verification: We forward the arithmic content of a formal proof (axioms, modus ponens) directly in $\Delta_{0}$, rather than constructing a $\Sigma_{1}$ abstraction to minimize overhead.
+    $\beta$-Function Storage (Theorem 3.1): Existentially quantify over parameters $c,d\in\mathbb{N}$ satisfying
+    $\beta(c,d,i)=f_{i}\quad\text{for all }i\in\{1,\ldots,N\},$ (2.5)
+    where $\beta(c,d,i)$ denotes the remainder when dividing $c$ by $1+(i+1)d$. This is enforced via quadratic constraints (Lemma 4.1):
+    $c$ $=q_{i}\cdot(1+(i+1)d)+r_{i},$ (2.6)
+[…]
   2. 2.
-    By Finite axiom sets: By working with a precalculated representations, we avoid the need to encode an infinite, or expressive axiom scheme.
-The trade-off is that our polynomial is potentially “wider” (in number of variables and monomials) than a optimal $\Sigma_{1}$ construction, but the argument is arguably more elementary (e.g., in terms of complexity), relying primarily on a variant of Gödel’s $\beta$-function (Theorem 2.1) and Zeckendorf representation (Definition 2.3).
-Having outlined our approach, we now formalize the key technical preliminaries underpinning the construction.
+    Zeckendorf Digits (Definition 3.3): Each $f_{i}$ is represented as a sum of non-consecutive Fibonacci numbers:
+    $f_{i}=\sum_{\kappa=2}^{K}F_{\kappa}\cdot d_{i,\kappa},\quad d_{i,\kappa}\in\{0,1\},\quad d_{i,\kappa}\cdot d_{i,\kappa+1}=0.$ (2.8)
+    This representation is unique and enables arithmetic verification without carry propagation (Exposition Exposition).
+2.2.2 Proof Verification (Stage 4)
+Each line $i$ must be justified either as an axiom or by modus ponens from earlier lines. This disjunction is encoded via boolean activation variables:
+  1. 1.
+    Axiom Test (Lemma 4.5): For a finite axiom set $\{g_{1},\ldots,g_{M}\}$, the constraint
+    $b_{ax,i}\cdot\sum_{\ell=1}^{M}(f_{i}-g_{\ell})^{2}=0,\quad b_{ax,i}\in\{0,1\}$ (2.9)
+    forces $f_{i}\in\{g_{1},\ldots,g_{M}\}$ when $b_{ax,i}=1$ (since a sum of squares vanishes iff each term vanishes).
+  2. 2.
+    Modus Ponens Test (Lemma 4.7): For each triple $(i,j,k)$ with $j,k<i$, the constraint
+    $b_{mp,i,j,k}\cdot(f_{i}-f_{j}-f_{k})=0,\quad b_{mp,i,j,k}\in\{0,1\}$ (2.10)
+    forces $f_{i}=f_{j}+f_{k}$ when $b_{mp,i,j,k}=1$. The Zeckendorf digits of $f_{i}$ automatically satisfy the non-adjacency constraint by uniqueness (Remark Exposition).
+  3. 3.
+    Exactly-One Justification (Lemma 4.11): The constraint
+    $b_{ax,i}+\sum_{j,k<i}b_{mp,i,j,k}=1$ (2.11)
+    ensures that exactly one justification mechanism activates per line.
+2.2.3 Degree Management (Stages 5–7)
+The proof verification constraints (Stage 4) produce equations of degree $\leq 3$, but merging them into a single equation via sum-of-squares yields degree $6$. We employ two techniques:
+  1. 1.
+    Guard-Gadgets (Definition 3.6): Each constraint $E(\mathbf{x})=0$ with $\delta(E)\leq 2$ is replaced by the guarded system
+    $u\cdot E(\mathbf{x})$ $=0,$ (2.12)
+    $u-1-v^{2}$ $=0,$ (2.13)
+    where $u,v\in\mathbb{N}$. The second equation forces $u=1+v^{2}\geq 1$, hence $u\neq 0$, so the first equation implies $E(\mathbf{x})=0$. This increases degree to at most $\max(1+2,2)=3$.
+  2. 2.
+    Jones Factorization (Theorem 5.1): After merging via sum-of-squares (yielding $P_{\text{merged}}$ of degree $6$), we apply recursive monomial shielding.
+Each monomial $x^{a}y^{b}z^{c}$ with $a+b+c>3$ is factored as $p\cdot q$ where $\delta(p),\delta(q)\leq 3$, introducing auxiliary variables $w$ via constraints $w=p\cdot q$ of degree $\leq 3$.
+This process (Algorithm 6.3) introduces $O(m\log d)$ variables for a polynomial with $m$ monomials of degree $\leq d$.
+2.3 Main Result
+Theorem 2.2 (Restatement of Theorem 5.2).
+Let $\mathrm{T}$ be a recursively axiomatizable theory with finite axiom set $\mathcal{A}$, and let $G_{T}$ be its Gödel sentence.
+There exists a polynomial $P(\mathbf{x})\in\mathbb{Z}[\mathbf{x}]$ of degree $\leq 3$ such that:
+$T\vdash G_{T}\quad\Longleftrightarrow\quad\exists\mathbf{x}\in\mathbb{N}^{m}:P(\mathbf{x})=0.$ (2.14)
+The polynomial has $m=O(K^{3}N^{3})$ variables, where $N$ is the proof length and $K=O(\log(\max_{i}f_{i}))$ bounds the Fibonacci representation size.
+2.4 On Single Cubic Equations
+While our construction produces a system of cubic equations, one might ask whether a single cubic equation could suffice. The following obstruction shows this is impossible.
+Proposition 2.3 (Impossibility of Universal Cubic Equation).
+There exists no polynomial $P_{\text{univ}}(n,\mathbf{x})\in\mathbb{Z}[n,\mathbf{x}]$ within an $\omega$-consistent formal system of degree $\leq 3$ such that for all Turing machines $M$ (encoded as $n\in\mathbb{N}$):
+$M\text{ halts}\quad\Longleftrightarrow\quad\exists\mathbf{x}\in\mathbb{N}^{k}:P_{\text{univ}}(n,\mathbf{x})=0.$ (2.15)
+Proof sketch.
+Assume such $P_{\text{univ}}$ exists.
+For each machine $M$ encoded as $n$, enumerate solutions $\mathbf{x}\in\mathbb{N}^{k}$ in order of increasing norm $|\mathbf{x}|_{1}=\sum_{i}x_{i}$.
+By the degree-$3$ constraint, the polynomial $P_{\text{univ}}(n,\mathbf{x})$ has at most $O(k^{3})$ monomials, so checking whether $P_{\text{univ}}(n,\mathbf{x})=0$ takes polynomial time in $|\mathbf{x}|_{1}$.
+For halting machines, the smallest solution $\mathbf{x}^{*}$ satisfies $|\mathbf{x}^{*}|_{1}=O(\text{poly}(n))$ (by the polynomial encoding of bounded computations).
+Thus, we could compute Chaitin [1975] constant.
+$\Omega=\sum_{M\text{ halts}}2^{-|M|}$ (2.16)
+by enumerating solutions up to polynomial size—contradicting the incomputability of $\Omega$, hence no such $P_{\text{univ}}$ exists. ∎
+Remark.
+Our construction avoids this obstruction by producing a machine-dependent system $\Phi_{M}(\mathbf{x})$.
+The proof length $N$ (and hence the variable count $m=O(K^{3}N^{3})$) is not computable from the machine description alone—this non-uniformity is essential.
+The system encodes "there exists a proof of length NN", not "the shortest proof has length NN".
+2.5 Prior Work and Roadmap
+We recall:
+  1. 1.
+    Robinson et al.
+[1961]: Showed exponential Diophantine relations are listable, reducing Hilbert’s 10th to finding a Diophantine representation of exponentiation.
+While their construction used higher degrees, our guard-gadget technique demonstrates degree $\leq 3$ suffices for universal representation over $\mathbb{N}$.
+  2. 2.
+    Jones [1982]: Achieved degree $\leq 4$ via direct MRDP reduction. Our guard-gadget technique (Definition 3.6) reduces this to degree $\leq 3$ by systematically shielding quadratic constraints.
+  3. 3.
+    Matiyasevich [1970]: Proved undecidability for degree $\leq 4$ (settling the “vanilla” Hilbert’s 10th Problem).
+Our result shows the degree barrier can be lowered by one, addressing a natural question raised by Jones [1982].
+  4. 4.
+    Hilbert’s Tenth Problem and its variants for rings, number fields, and related structures remain central questions in mathematics [Alpöge et al., 2025].
+However, from a computational perspective, structures like $\mathbb{Z}$ reduce to $\mathbb{N}$ via standard encodings.
+Mathematical results, see [Poonen, 2009], concern definability complexity in strong foundational theories, which is orthogonal to our focus on bounded provability in weak arithmetics.
+The remainder of the paper fills in the technical details:
+  1. 1.
+    Section 3 (Preliminaries): Establishes encoding tools ($\beta$-function, four-squares, Zeckendorf, guard-gadgets).
+  2. 2.
+    Section 4 (Construction): Builds the cubic system via proof sequence encoding, axiom testing, and modus ponens verification.
+  3. 3.
+    Section 5 (Aggregation): Merges the system via sum-of-squares and applies Jones factorization.
+  4. 4.
+    Section 6 (Computation): Provides algorithmic details for degree reduction.
+  5. 5.
+    Section 7 (Result): Proves soundness and completeness bidirectionally.
+Each stage is constructed to be effective (algorithmically computable), polynomial-time (in the proof parameters), and degree-preserving (maintaining $\delta\leq 3$ after Stage 5).
 3 Preliminaries
@@
 The following logic allows encoding arbitrary finite sequences as Diophantine constraints:
-Procedure 2.1 (Diophantine $\beta$-Function).
+Procedure 3.1 (Diophantine $\beta$-Function).
 By the work of Gödel [1931] (using the “Chinese remainder”) and as exposited in Matiyasevich [1993], there exists a polynomial $\beta(c,d,i)$ over $\mathbb{N}$ such that for any finite sequence $a_{0},a_{1},\ldots,a_{n}\in\mathbb{N}$, there exist $c,d\in\mathbb{N}$ with
@@
 Remark.
-In the completeness proof of Theorem 3.16, the witness construction algorithm (Exposition 3.9) must compute four-square decompositions for each remainder $r_{i}$ and its bound $(i+1)d-1-r_{i}$.
+In the completeness proof of Theorem 4.16, the witness construction algorithm (Exposition 4.9) must compute four-square decompositions for each remainder $r_{i}$ and its bound $(i+1)d-1-r_{i}$.
 By the above algorithm, this can be done in time $O(\log((i+1)d))$ per decomposition, yielding total time $O(N\cdot\log(\max_{i}f_{i}))$ for all $N$ proof lines.
@@
   1. 1.
-    Encode the sequence using $\beta(c,d,i)=f_{i}$ via constraint (3.34) and Lemma $2.$
+    Encode the sequence using $\beta(c,d,i)=f_{i}$ via constraint (4.47) and Lemma 3.2.
   2. 2.
@@
 3.4 Summary of Polynomial Constraints
-Constraint Equation $n\leq\delta$
-$\beta$-Division $c-q_{i}(1+(i+1)d)-r_{i}=0$ 2
-Lower bound $r_{i}-a_{i}^{2}-b_{i}^{2}-c_{i}^{2}-e_{i}^{2}=0$ 2
-Upper bound $(i+1)d-1-r_{i}-\alpha_{i}^{2}-\beta_{i}^{2}-\gamma_{i}^{2}-\delta_{i}^{2}=0$ 2
-Boolean digits $d_{i,\kappa}-d_{i,\kappa}^{2}=0$ 2
-Non-adjacency $d_{i,\kappa}\cdot d_{i,\kappa+1}=0$ 2
-Zeckendorf sum $f_{i}-\sum_{\kappa=2}^{K}F_{\kappa}d_{i,\kappa}=0$ 1
-Table 1: All constraints have degree ($\delta$) $\leq 2$ before applying guard multipliers (see Section 2.5).
+Constraint Expanded Equation $\boldsymbol{\delta}$ $\mathbb{N}$-Mechanism
+Pre guard-wrapping
+$\beta$-Division $c-q_{i}(1+(i+1)d)-r_{i}=0$ 2 Linear in $c,r_{i}$; quadratic in $q_{i},d$
+Lower bound $r_{i}-a_{i}^{2}-b_{i}^{2}-c_{i}^{2}-e_{i}^{2}=0$ 2 Sum of squares $\geq 0$
+Upper bound $(i+1)d-1-r_{i}-\alpha_{i}^{2}-\beta_{i}^{2}-\gamma_{i}^{2}-\delta_{i}^{2}=0$ 2 Sum of squares $\geq 0$
+Boolean $d_{i,\kappa}-d_{i,\kappa}^{2}=0$ 2 Forces $d_{i,\kappa}\in\{0,1\}$
+Non-adjacency $d_{i,\kappa}\cdot d_{i,\kappa+1}=0$ 2 Product vanishes
+Zeckendorf sum $f_{i}-\sum_{\kappa=2}^{K}F_{\kappa}d_{i,\kappa}=0$ 1 Linear combination
+Axiom membership $b_{ax,i}\cdot\sum_{\ell=1}^{M}(f_{i}-g_{\ell})^{2}=0$ 3 Boolean times sum-of-squares
+MP arithmetic $f_{i}-f_{j}-f_{k}=0$ 1 Linear equality
+Exactly-one $b_{ax,i}+\sum_{j,k<i}b_{mp,i,j,k}-1=0$ 1 Linear in booleans
+Target check $f_{N}-g_{T}=0$ 1 Linear equality
+Post guard-wrapping
+$\beta$-Division guard $u_{\beta,i}\cdot(c-q_{i}(1+(i+1)d)-r_{i})=0$ 3 $u_{\beta,i}=1+v_{\beta,i}^{2}\geq 1$
+Lower-bound guard $u_{L,i}\cdot(r_{i}-a_{i}^{2}-b_{i}^{2}-c_{i}^{2}-e_{i}^{2})=0$ 3 $u_{L,i}=1+v_{L,i}^{2}\geq 1$
+Upper-bound guard $u_{U,i}\cdot((i+1)d-1-r_{i}-\alpha_{i}^{2}-\cdots)=0$ 3 $u_{U,i}=1+v_{U,i}^{2}\geq 1$
+Boolean guard $u_{B,i,\kappa}\cdot(d_{i,\kappa}-d_{i,\kappa}^{2})=0$ 3 $u_{B,i,\kappa}=1+v_{B,i,\kappa}^{2}\geq 1$
+Non-adjacency guard $u_{N,i,\kappa}\cdot(d_{i,\kappa}\cdot d_{i,\kappa+1})=0$ 3 $u_{N,i,\kappa}=1+v_{N,i,\kappa}^{2}\geq 1$
+Zeckendorf guard $u_{Z,i}\cdot(f_{i}-\sum_{\kappa}F_{\kappa}d_{i,\kappa})=0$ 2 $u_{Z,i}=1+v_{Z,i}^{2}\geq 1$
+Axiom guard (Already $\delta=3$ before wrapping) 3 Boolean activation
+MP-sum guard $u_{mp,i,j,k}\cdot(f_{i}-f_{j}-f_{k})=0$ 2 $u_{mp,i,j,k}=1+v_{mp,i,j,k}^{2}\geq 1$
+Exactly-one guard $u_{J,i}\cdot(b_{ax,i}+\sum_{j,k}b_{mp,i,j,k}-1)=0$ 2 $u_{J,i}=1+v_{J,i}^{2}\geq 1$
+Target guard $\mathrm{T}\cdot(f_{N}-g_{T})=0$ 2 $\mathrm{T}=1+U^{2}\geq 1$
+Positivity constraints
+All guards $u_{\bullet}-1-v_{\bullet}^{2}=0$ 2 Forces $u_{\bullet}\geq 1$
+Global multiplier $\mathrm{T}-1-U^{2}=0$ 2 Forces $\mathrm{T}\geq 1$
+Table 2: Complete degree bookkeeping. Base constraints have $\delta\leq 3$; guard-wrapping preserves $\delta\leq 3$ for all main constraints, with guard positivity at $\delta=2$.
 3.5 Guard-Gadget Framework
@@
 We now construct a finite system of polynomial equations of degree at most $3$ such that the system has a solution over $\mathbb{N}$ if and only if the Fibonacci variant theory $T_{\text{Fib}}$ proves the target formula $G_{T}$.
-The construction proceeds in six stages: encode a proof sequence $\langle f_{1},\ldots,f_{N}\rangle$ using the $\beta$-function and Zeckendorf representation (3.1), implement Fibonacci addition for modus ponens verification (3.3), test axiom membership (3.3), enforce exactly-one justification per proof line (3.5), require the final line to be the target formula (3.6), and assemble the complete system (3.7).
 4.1 Proof Sequence Encoding
@@
 This maintains the overall degree bound of $3$ for the entire system.
-Procedure 3.9 (Greedy Witness Construction).
+Procedure 4.9 (Greedy Witness Construction).
 The completeness direction of Lemma 4.8 relies on the computability of Zeckendorf representations.
@@
 Proposition 4.12 (Justification Correctness).
-Constraint 3.11 necessitates that exactly one of the following holds for each line $i$:
+Constraint 4.11 necessitates that exactly one of the following holds for each line $i$:
   1. 1.
@@
 Proof.
-Since $b_{ax,i},b_{mp,i,j,k}\in\{0,1\}$ (by 3.10), equation (3.56) necessitates:
+Since $b_{ax,i},b_{mp,i,j,k}\in\{0,1\}$ (by 4.10), equation (4.69) necessitates:
 $b_{ax,i}+\sum_{j,k<i}b_{mp,i,j,k}=1.$ (4.71)
@@
 We consider the activation mechanism.
-If $b_{ax,i}=1$, then Constraint 3.5 forces $f_{i}\in\mathcal{A}$.
-If $b_{mp,i,j,k}=1$, then Constraint 3.10 forces $f_{i}=m_{i,j,k}=f_{j}+f_{k}$ (assuming the Fibonacci addition constraints hold).
-Constraint 3.11 ensures that exactly one of these activation variables is $1$ for each line $i$.
+If $b_{ax,i}=1$, then Constraint 4.5 forces $f_{i}\in\mathcal{A}$.
+If $b_{mp,i,j,k}=1$, then Constraint 4.10 forces $f_{i}=m_{i,j,k}=f_{j}+f_{k}$ (assuming the Fibonacci addition constraints hold).
+Constraint 4.11 ensures that exactly one of these activation variables is $1$ for each line $i$.
 This encodes the informal proof rule: "Each line is either an axiom or follows from two earlier lines by MP."
@@
 Lemma 4.14 (Target Correctness).
-Constraint 3.13 forces $f_{N}=g_{T}$ (i.e., the final line of the proof is the target formula).
+Constraint 4.13 forces $f_{N}=g_{T}$ (i.e., the final line of the proof is the target formula).
 Proof.
@@
     1. (a)
-      By Constraint 3.1, $\beta(c^{*},d^{*},i)=f^{*}_{i}$ for all $i\leq N^{*}$.
+      By Constraint 4.1, $\beta(c^{*},d^{*},i)=f^{*}_{i}$ for all $i\leq N^{*}$.
     2. (b)
-      By Constraint 3.3, each $f^{*}_{i}$ has a valid Zeckendorf representation.
+      By Constraint 4.3, each $f^{*}_{i}$ has a valid Zeckendorf representation.
     3. (c)
-      By Constraint 3.11, each line $i$ is justified either as an axiom (Constraint 3.5) or by MP (Constraint 3.10).
+      By Constraint 4.11, each line $i$ is justified either as an axiom (Constraint 4.5) or by MP (Constraint 4.10).
     4. (d)
-      By Constraint 3.13, $f^{*}_{N^{*}}=g_{T}$.
+      By Constraint 4.13, $f^{*}_{N^{*}}=g_{T}$.
   4. 4.
@@
     1. (a)
-      Each $\varphi_{i}$ is either an axiom (see 3.3) or derived from earlier formulas by MP (by 3.12),
+      Each $\varphi_{i}$ is either an axiom (see 4.3) or derived from earlier formulas by MP (by 4.12),
     2. (b)
@@
 5.1 Merging via Sum-of-Squares
-By the standard technique [Matiyasevich, 1970, 1993], the system
-$\Phi=\{P_{1}=0,\ldots,P_{m}=0\}$
-with $\delta(P_{i})\leq 3$ is equivalent to:
-$P_{\text{merged}}(\mathbf{x})=\sum_{i=1}^{m}P_{i}(\mathbf{x})^{2}=0.$ (4.63)
-Since $\delta(P_{i})\leq 3$, we have $\delta(P_{\text{merged}})\leq 6$. Hence Exposition, $m=O(KN^{3})$ where $N$ is the proof length and $K=O(\log(\max_{i}f_{i}))$.
+We reduce the system $\Phi(\mathbf{x})$ from Theorem 4.16 to a single equation via sum-of-squares.
+Degree bound. By Table 2, all polynomials $P_{i}\in\Phi$ satisfy $\delta(P_{i})\leq 3$ after guard-wrapping (Definition 3.6).
+Merger. The system $\Phi=\{P_{1}=0,\ldots,P_{m}=0\}$ is equivalent to
+$P_{\text{merged}}(\mathbf{x})=\sum_{i=1}^{m}P_{i}(\mathbf{x})^{2}=0,\quad\delta(P_{\text{merged}})\leq 6.$ (5.76)
+Size. With $m=O(KN^{3})$ constraints (Remark Exposition) and $s_{i}=O(K)$ monomials per $P_{i}$, we obtain $O(K^{3}N^{3})$ total monomials.
 5.2 Degree Reduction to Cubic
@@
 By our additive approach, theorem 4.16 (finite cubic system), sum-of-squares merging via (5.76), and theorem 5.1. ∎
-Corollary 4.3 (Cubic Incompleteness).
+Corollary 5.3 (Cubic Incompleteness in $\mathbb{N}$).
 The solvability problem for diophantines with $\delta=3$ over $\mathbb{N}$ is undecidable.
@@
 Then $\delta(Q_{E})=\max(1+\delta(E),2)\leq 3$ and $Q_{E}=0\iff E=0$ when $U\geq 1$.
-Procedure 5.2 (Monomial Shielding).
+Procedure 6.2 (Monomial Shielding).
 Reduces a single monomial $m(\mathbf{x})$ of degree $d>3$ to degree $\leq 3$ via recursive factorization, maintaining the invariant that every introduced constraint has degree $\leq 3$.
 […]
-Procedure 5.3 (Polynomial Degree Reduction).
+Procedure 6.3 (Polynomial Degree Reduction).
 Reduces polynomial $R(\mathbf{x})$ with $\delta(R)>3$ to an equivalent system where all constraints have degree $\leq 3$.
@@
 The total monomial count is therefore $m=\sum_{i=1}^{m_{0}}s_{i}^{2}=O(m_{0}\bar{s}^{2})$, where $\bar{s}=\max_{i}s_{i}=O(K)$ by the bounds on Zeckendorf digit sums.
-Since Section 3.7 establishes that $m_{0}=O(KN^{3})$ for the total number of constraints, Proposition 5.6 implies that the number of auxiliary variables required is $O(m\log 6)=O(K^{3}N^{3})$.
-Combining this with the $O(KN^{3})$ base variables from Remark Exposition, the total variable count is $O(K^{3}N^{3})$.
+Since Section 4.7 establishes that $m_{0}=O(K^{3}N^{3})$ for the total number of constraints, Proposition 6.6 implies that the number of auxiliary variables required is $O(m\log 6)=O(K^{3}N^{3})$.
+Combining this with the $O(K^{3}N^{3})$ base variables from Remark Exposition, the total variable count is $O(K^{3}N^{3})$.
 Lemma 6.7 (Cubic Reduction).
@@
 Set $f_{i}^{*}=f_{i}$ for all $i\leq N$, and compute remainders $r_{i}^{*}$ and auxiliary variables via Lemma 4.2.
-Step 3. Zeckendorf digits. For each $i\leq N$, compute the Zeckendorf representation using Procedure 3.9:
+Step 3. Zeckendorf digits. For each $i\leq N$, compute the Zeckendorf representation using Procedure 4.9:
 $d_{i,\kappa}^{*}=\mathbb{1}_{\kappa\in S_{i}},$ (7.88)
@@
 Conclusion. All constraints vanish under $\vec{x}^{*}$, hence $P_{\text{hard}}(\vec{x}^{*})=0$. ∎
+7.3 Main Algorithm: Encoding Gödel’s Theorem
+We present the complete reduction in algorithmic form:
+[…]
+Theorem 7.3 (Correctness of Algorithm 7.3).
+The polynomial $P_{\text{final}}$ output by the algorithm satisfies:
+$T\vdash G_{T}\quad\Longleftrightarrow\quad\exists\mathbf{x}\in\mathbb{N}^{m}:P_{\text{final}}(\mathbf{x})=0.$ (7.90)
+Proof.
+By Theorem 4.16 (soundness and completeness of Stages 1–6), Equation (5.76) (sum-of-squares equivalence), and Proposition 6.5 (degree reduction preserves solvability). ∎
 7.4 Undecidability Consequences
@@
 Proof.
-By Theorems 6.1 and 6.2, solvability of $P_{\text{hard}}$ is equivalent to $\mathrm{T}_{\text{Fib}}\vdash G_{T}$.
+By Theorems 7.1 and 7.2, solvability of $P_{\text{hard}}$ is equivalent to $\mathrm{T}_{\text{Fib}}\vdash G_{T}$.
 By [Gödel, 1931], if $\mathrm{T}_{\text{Fib}}$ is consistent, then $G_{T}$ is independent of $\mathrm{T}_{\text{Fib}}$.
@@
 7.5 Impredicativity
-Thesis 6.5 (Impredicativity of Degree Threshold).
+Thesis 7.6 (Impredicativity of Degree Threshold).
 Let $\mathrm{T}$ be any consistent, recursively axiomatizable theory extending Robinson arithmetic $\mathsf{Q}$.
@@
 […]
-Then there exists no first-order formula $B(k)$ in the language of arithmetic such that $T$ proves:
+Then there exists no first-order formula $B(k)$ in the language of arithmetic such that $\mathrm{T}$ proves:
   1. 1.
@@
 Suppose such $B(k)$ exists with unique $k_{0}$ satisfying $\mathrm{T}\vdash B(k_{0})$.
-By the diagonal lemma, construct a Diophantine equation $E_{k_{0}}$ with $\delta(E_{k_{0}})=k_{0}$ whose solvability encodes $\neg B(k_{0})$ within $T$.
-If $E_{k_{0}}$ is solvable, then $T\vdash\neg B(k_{0})$, contradicting uniqueness.
+By the diagonal lemma, construct a Diophantine equation $E_{k_{0}}$ with $\delta(E_{k_{0}})=k_{0}$ whose solvability encodes $\neg B(k_{0})$ within $\mathrm{T}$.
+If $E_{k_{0}}$ is solvable, then $\mathrm{T}\vdash\neg B(k_{0})$, contradicting uniqueness.
 If $E_{k_{0}}$ is unsolvable, then $\mathcal{D}_{k_{0}}$ contains a decidable equation, contradicting clause (3).
@@
 This mirrors the Grelling–Nelson paradox and reflects the predicativity restriction in the type theory of Martin-Löf [1980], where universes $\mathcal{U}_{i}$ cannot quantify over themselves.
-Thesis 6.6 (Diophantine Impredicativity).
+Thesis 7.7 (Diophantine Impredicativity).
 In the Calculus of Inductive Constructions, attempting to form:
@@
 requires impredicative quantification over the decidability predicate $\mathsf{Decidable}(-)$, which itself depends on the threshold $k$ being defined.
-By Thesis 6.5, no closed term of this type exists in any consistent extension of $\mathsf{Q}$.
+By Thesis 7.6, no closed term of this type exists in any consistent extension of $\mathsf{Q}$.
 The diagonal equation $E_{k_{0}}$ acts as a negative membership witness:
@@
   2. 2.
-    No uniform refutation exists (by Thesis 6.5, any such refutation enables constructing $E_{k_{0}}$, violating consistency).
+    No uniform refutation exists (by Thesis 7.6, any such refutation enables constructing $E_{k_{0}}$, violating consistency).
 Thus, the threshold constitutes a Brouwerian counterexample to the law of excluded middle: a statement $P$ such that neither $P$ nor $\neg P$ admits constructive proof.
 The degree-$3$ bound is visible only from a classical metatheory (e.g., $\mathsf{ZFC}$ proves $\mathcal{D}_{3}$ is undecidable” by encoding the halting problem), but no predicative subsystem of $\mathsf{PA}$ can internalize this fact.
-Thesis 6.8 (Impredicativity of Degree Threshold).
+Thesis 7.9 (Impredicativity of Degree Threshold).
 Let $\mathrm{T}$ be any consistent, effectively axiomatizable theory extending Robinson arithmetic $\mathsf{Q}$.

### v4

@@
-Preprint Manuscript
-Preprint Manuscript
-Third Version
-CUBIC INCOMPLETENESS: HILBERT’S TENTH PROBLEM OVER $\mathbb{N}$ STARTS AT $\delta=3$
-Milan Rosko
-(Oktober 2025)
-We prove that Hilbert’s Tenth Problem over $\mathbb{N}$ remains undecidable when restricted to cubic equations (degree $\leq 3$), resolving the open case $\delta=3$ identified by Jones (1982) and establishing sharpness against the decidability barrier at $\delta=2$ (Lagrange’s four-square theorem).
-For any consistent, recursively axiomatizable theory $T$ with Gödel sentence $G_{T}$, we effectively construct a single polynomial $P(x_{1},\ldots,x_{m})\in\mathbb{Z}[\mathbf{x}]$ of degree $\leq 3$ such that $T\vdash G_{T}$ if and only if $\exists\mathbf{x}\in\mathbb{N}^{m}:P(\mathbf{x})=0$.
-Our reduction proceeds through four stages with explicit degree and variable accounting.
-Proof-sequence encoding via Diophantine $\beta$-function and Zeckendorf representation yields $O(KN)$ quadratic constraints ($K=O(\log(\max_{i}f_{i}))$, NN = proof length).
-Guard-gadgets wrap each constraint $E(\mathbf{x})=0$ as $u\cdot E(\mathbf{x})=0$, $u=1+v^{2}$, preserving degree $\leq 3$ with $O(KN^{3})$ variables.
-Sum-of-squares aggregation $P_{\text{merged}}=\sum_{i}P_{i}^{2}$ produces degree $\leq 6$ with $O(KN^{3})$ monomials.
-Recursive monomial shielding reduces degree to $\leq 3$ in $O(\log d)$ rounds, adding $O(K^{3}N^{3})$ variables total.
-Keywords: Proof Theory, Logic, Hilbert’s Tenth Problem, Zeckendorf’s Theorem
-Open Problem, Constructive Mathematics, Fibonacci, Cubic equations
-Primary MSC: 11U05, 03F40, 03D35
-Secondary MSC: 11D72, 03B25, 11Y16
-Milan Rosko is from University of Hagen, Germany
-Email: Q1012878 @@ studium.fernuni-hagen.de – ORCID: 0009-0003-1363-7158
-Licensed under “Deed” ΓΔ  – creativecommons.org/licenses/by/4.0
-1 Corrigendum
-1.1 First Corrigendum
-The initial version encoded a Gödel sentence as a Diophantine system rather than a single equation.
-The ambition to present a concrete standalone equation introduced an unwarranted bias into the construction.
-Subsequent algorithmic approaches yielded a massive closed form, but the reduction diverted attention from the central epistemic questions.
-1.2 Second Corrigendum
-In the second revision, we pivoted to explicit witness construction, ensuring all constraints remain genuinely cubic, and grounded the argument in logical foundations rather than forcing a premature encoding.
-1.3 Third Corrigendum
-In this third—still preliminary—version, we resolve ambiguities between logical and mathematical layers ($\mathbb{Z}$ versus $\mathbb{N}$), and improve on bookkeeping.
-The first-order argument is developed independently of its Diophantine realization.
-We provide a high-level overview at the outset, then verify each layer on its own terms.
-[…]
-We now adopt a direct approach: all variables, constraints, and solutions are understood to inhabit $\mathbb{N}$ from the outset.
-The $\beta$-function (Theorem 3.1), four-squares encoding (Lemma 4.2), and Zeckendorf representation (Definition 3.3) operate naturally over $\mathbb{N}$.
-The guard-gadget framework (Definition 3.6) enforces positivity via $u=1+v^{2}\geq 1$ without recourse to negative integers.
-This interpretation aligns with the historical formulation of Hilbert’s Tenth Problem over $\mathbb{N}$ and avoids unnecessary detours through $\mathbb{Z}$.
-The approach is straightforward: Turing machines operate with discrete symbols.
-1.4 Context
-This paper shares ideas with different work [^1] The Fractal Logic of $\Phi$-adic Recursion, arXiv:2510.08934, [2025a], in that a compression-undecidability relationship is visible in the context of extensions from $\mathbb{N}$ to $\mathbb{Z}$.
-The additional compression offered by negative numbers permits lower polynomial degrees in Diophantine equations, but at the cost of increased uncertainty between powers.
-The equation $x^{2}=4$ is uniquely solved in $\mathbb{N}$ by $x=2$, while in $\mathbb{Z}$ the solution set comprises $\{-2,2\}$.
-More dramatically, $x^{4}=16$ admits four solutions $\{\pm 2,\pm 2i\}$ in $\mathbb{C}$, with the real solutions $\pm 2$ already exhibiting the doubling.
-Future work will focus on this: The ambiguity is not a deficiency but a necessary consequence of Kolmogorov [1968] complexity: syntactic reduction of degrees is purchased through semantic branching.
-Results possessing nontrivial structure in $\mathbb{N}$ collapse in $\mathbb{Z}$ to trivial symmetries, whereby the displaced complexity migrates into sign ambiguity.
-2 Overview
-2.1 The Problem
-Previously [^2] A Fibonacci-Based Gödel Numbering: $\Delta_{0}$ Semantics Without Exponentiation, arXiv:2509.10382, [2025b], we proposed that a Zeckendorf-based arithmetization of Gödel numbering—eschewing exponentiation in favor of additive structure—might lower the formal overhead in encoding undecidability.
-We now apply this method to produce undecidable Diophantine equations of degree three, thereby advancing “Hilbert’s Tenth Problem” [^3] As posed by Hilbert [1900].
-into the cubic regime ($\delta=3$) and resolving a previously open case: We cite from the article “Universal Diophantine Equation”, by Jones [1982]:
+CUBIC INCOMPLETENESS: HILBERT’S TENTH PROBLEM AT DEGREE THREE
+Abstract
+We analyze the cubic fragment $\mathcal{D}_{3}$ over $\mathbb{N}$ by isolating the uniform closure principle any total correct cubic solver would have to realize.
+In $\mathsf{HA}$ we give a fully constructive, additive and degree-controlled encoding of bounded verification: for each externally fixed bound, we effectively produce a finite system of degree-3 Diophantine equations whose solutions represent the existence of the corresponding finite proof or computation trace.
+The encoding is purely syntactic, using “gadgets” and Carryless Pairing.
+In a classical metatheory (e.g.
+$\mathsf{PA}$) we show that the global solver hypothesis implies a uniform operator eliminating the bound inside $\mathcal{D}_{3}$, which is incompatible with standard non-uniformity/realizability constraints.
+Hence no uniform cubic can exist clasically.
+Subject: Proof Theory, Diophantine Geometry, Computability Theory
+1 Introduction
+Exposition.
+Figure 1: The construction proceeds through strictly separated layers.
+Undecidability is established at the logical level; a classical reflection step then shows that no algorithm can solve the problem uniformly, due to an intrinsic global barrier.
+The proof is stratified across constructive, equational, and reflective layers.
+Previously [2025], we proposed that a Zeckendorf-based arithmetization of Gödel numbering—eschewing exponentiation in favor of additive structure—can lower the formal overhead of undecidability encodings.
+We now apply this method to the cubic regime, addressing the degree-$3$ remainder singled out by Jones [1982] in Universal Diophantine Equation:
 We know that there exist universal diophantine equations of degree 4 […] but none of degree 2.
@@
 We do not know whether Hilbert’s tenth problem is decidable or undecidable for equations of degree 3.
-Hence, we define the open question:
-Problem 2.1 (Cubic Remainder of Hilbert’s 10th Problem).
-Let $\delta=3$ denote the set of all polynomial Diophantine equations of the form
-$P(x_{1},\ldots,x_{n})=0,\quad P\in\mathbb{N}[x_{1},\ldots,x_{n}],\quad\delta(P)\leq 3.$ (2.1)
-Question: Is there an algorithm that, given $P$ and $\delta$ = 3, decides whether
+The guiding analogy is the classical impossibility of the Delian problem, best read as a closure failure.
+Straightedge–compass constructions generate exactly the constructible numbers: the smallest class containing $\mathbb{Q}$ and closed under the permitted operations (equivalently, under towers of quadratic extensions).
+Doubling the cube would require closing these rules under a genuinely new operation (taking cube roots).
+The obstruction is therefore global: no amount of local ingenuity can replace a missing closure principle.
+Our degree-$3$ analysis has the same logical shape (though the closure is logical rather than algebraic).
+A uniform solver hypothesis for $\mathcal{D}_{3}$ would have to supply an additional closure step: a uniform elimination of the external bound inside the cubic fragment (a cubic closure operator).
+The decisive barrier is thus not an algebraic accident (a particular gadget failing), but the absence of a global closure principle: bounded representability holds, while uniform existential closure is the forbidden move.
+2 Ansatz
+2.1 Core Phenomena
+Problem 2.1 (Cubic remainder of Hilbert’s Tenth Problem over $\mathbb{N}$).
+Let
+$\mathcal{D}_{3}\;\equiv\;\{\,P\in\mathbb{Z}[x_{1},\ldots,x_{n}]:\delta(P)\leq 3\text{ for some }n\,\}.$ (2.1)
+Question: Is there an algorithm which, given a code for $P\in\mathcal{D}_{3}$, decides whether
 $\exists(x_{1},\ldots,x_{n})\in\mathbb{N}^{n}\;:\;P(x_{1},\ldots,x_{n})=0\;?$ (2.2)
-2.2 Roadmap
+Observation.
+We read the degree-$3$ remainder of Hilbert’s Tenth Problem as a closure problem rather than as a missing gadget or an incremental technical gap between the known degree-$4$ universal constructions and the degree-$2$ decidable regime.
+The decisive distinction is between:
+  1. (i)
+    bounded representability (local, finitary verification), and
+  2. (ii)
+    uniform existential closure.
+At degree $3$ the cubic fragment is already expressive enough to represent, with externally fixed bounds, the finite witnesses that arise in verification phenomena: bounded computation traces, bounded proof traces, and other finitary tableaux.
+Formally, for broad families of bounded predicates $R_{B}(x)$ (with $B$ an external numerical bound), one can effectively associate cubic Diophantine instances $P_{x,B}\in\mathcal{D}_{3}$ such that
+$R_{B}(x)\quad\Longleftrightarrow\quad\mathsf{Sol}_{3}(\ulcorner P_{x,B}\urcorner),$ (2.3)
+uniformly in $(x,B)$. The uniform decision problem, however, is inherently unbounded: it asks for a single algorithm deciding the existential closure
+$\exists B\;R_{B}(x).$ (2.4)
+Any putative uniform solver for $\mathcal{D}_{3}$ would therefore have to implement, in one form or another, a uniform closure principle that internalizes this unbounding step within the cubic fragment: a uniform procedure that, given $x$, compiles the family $(P_{x,B})_{B\in\mathbb{N}}$ into a single cubic instance $Q_{x}\in\mathcal{D}_{3}$ satisfying
+$\exists B\;R_{B}(x)\quad\Longleftrightarrow\quad\mathsf{Sol}_{3}(\ulcorner Q_{x}\urcorner).$ (2.5)
+This is not an auxiliary convenience but the logical core of uniformity: a uniform solver is, in effect, precisely a uniform method for closing the external existential bound.
+In this sense the cubic remainder is self-similar.
+Degree $3$ is the first regime in which bounded verification can be arithmetized cleanly while the fragment still fails to be closed under the one semantic operation that would turn bounded representability into a global decision procedure.
+Degree $4$ has enough algebraic slack to absorb the closure move (hence universality), degree $2$ lacks the expressive room to host the bounded encodings at all, and degree $3$ is the critical boundary.
+2.2 Logical Architecture and Scope
+Exposition.
+We analyze the proof-theoretic content of a global solver hypothesis for the cubic fragment and isolate the uniform closure principle that such a hypothesis would have to validate under a constructive reading.
+The contradiction is obtained by a stratified Modus Tollens argument: we (i) formulate the solver hypothesis as a single arithmetical commitment, (ii) identify the uniform witness-producing content this commitment would carry under the intended constructive semantics, and (iii) refute precisely that witness-producing content by a realizability/non-uniformity barrier.
+A final reflection step is used only to pass from the constructive impossibility of realizers to the classical negation of the corresponding existence claim.
+Ansatz 2.2.
+Hilbert’s Tenth Problem, cf.
+Hilbert [1900], is a uniformity question: it asks whether there exists a single effective procedure which, given an arbitrary Diophantine equation from a prescribed class, decides solvability over $\mathbb{N}$.
+In this paper the class of interest is the cubic fragment
+$\mathcal{D}_{3}\;\equiv\;\{\,P\in\mathbb{Z}[x_{1},\ldots,x_{n}]:\delta(P)\leq 3\text{ for some }n\,\}.$
+The corresponding global statement is therefore:
+There is no uniform algorithm which, on input $P\in\mathcal{D}_{3}$, decides whether $\exists\vec{x}\in\mathbb{N}^{n}\;P(\vec{x})=0$.
+Importantly, we do not pursue this by exhibiting a distinguished cubic polynomial whose solvability is independent of a background theory, nor by constructing a universal cubic polynomial in the classical MRDP sense, cf.
+Matiyasevich [1993].
+The obstruction analyzed here is global and logical: it is the incompatibility of a uniform method with the constructive non-uniformity phenomena made explicit below.
+Definition 2.3 (Cubic Solvability Predicate).
+Fix a Gödel Numbering $p\mapsto P_{p}$ of integer-coefficient polynomials together with their arities.
+Write $p\in\mathrm{Code}(\mathcal{D}_{3})$ if $p$ codes a polynomial of total degree at most $3$.
+Define the arithmetical predicate
+$\mathsf{Sol}_{3}(p)\;:\Longleftrightarrow\;\bigl(p\in\mathrm{Code}(\mathcal{D}_{3})\bigr)\ \wedge\ \exists\vec{x}\in\mathbb{N}^{n(p)}\;P_{p}(\vec{x})=0.$
+Definition 2.4 (Total Correct Cubic Solver).
+Let $\mathsf{Run}(e,p,t,b)$ be the standard primitive recursive predicate expressing:
+“The machine with code $e$ halts on input $p$ within $t$ steps and outputs the bit $b\in\{0,1\}$.”
+We formalize the claim
+$e$ is a total correct solver for $\mathcal{D}_{3}$ over $\mathbb{N}$
+as the closed arithmetical sentence:
+$\mathsf{Solve}_{3}(e)\;\Longleftrightarrow\;\forall p\Bigl($ $\,p\in\mathrm{Code}(\mathcal{D}_{3})\ \to\ \exists t\,\exists b\in\{0,1\}\;$
+$\bigl(\mathsf{Run}(e,p,t,b)\ \wedge\ (b=1\leftrightarrow\mathsf{Sol}_{3}(p))\bigr)\Bigr).$ (2.6)
+Remark.
+Equation (2.4) is the global reading of the solver hypothesis: it is a single $\forall\exists$-commitment asserting totality and correctness on all cubic inputs.
+It is not an instruction to perform an external search over $(t,b)$; rather, it asserts that for each input $p$ there exists a terminating computation trace witnessing the solver’s output and that the output is correct.
+The analysis in this paper is intentionally proof-theoretic, cf.
+Boolos et al.
+[2007].
+We do not assume that from the mere (classical) truth of $\mathsf{Solve}_{3}(e)$ one can extract witnesses.
+Instead, we study what follows under a constructive reading: in settings such as in Heyting Arithmetic ($\mathsf{HA}$) equipped with a sound realizability interpretation, cf.
+Troelstra and van Dalen [1988]; Heyting [1930] for the relevant $\forall\exists$-fragment, a derivation of (2.4) entails a uniform witness-producing method extracting $(t,b)$ (together with a correct trace) from $p$.
+This is the precise sense in which uniformity becomes logically explicit.
+Ansatz 2.5 (Stratified Method).
+The argument is deliberately stratified into three layers, each with its own role, see Figure 1. We keep the explicit encoding separate from its metatheoretic reading.
+Layer I.
+Working in $\mathsf{HA}$, we build explicit cubic Diophantine gadgets for bounded verification phenomena: finite proof traces and finite computation traces.
+All constructions here are effective and degree-controlled, but the relevant bounds (proof length, runtime) are supplied externally.
+Layer II.
+We apply equisatisfiable algebraic transformations that reorganize bounded systems without changing solvability over $\mathbb{N}$.
+The purpose of this layer is bookkeeping: it normalizes shape while preserving the degree bound and makes explicit where standard “packaging” moves attempt to leave the cubic regime.
+Layer III.
+In a classical metatheory (one may take $\mathsf{PA}$ for definiteness), we treat the solver hypothesis as a definite arithmetical commitment.
+The only use of classical reasoning at this layer is a reflection step applied to program-existence claims singled out by the constructive analysis: once the relevant realizer is shown not to exist, Modus Tollens yields the classical negation of the corresponding solver hypothesis.
+Remark.
+The paper isolates two independent non-uniformities and keeps them disjoint.
+  1. (i)
+    Bound selection.
+There is no computable function that, from $(\mathrm{T},\varphi)$, produces a bound sufficient for proof search.
+This is the classical incompleteness/undecidability non-uniformity.
+  2. (ii)
+    Cubic closure.
+Even when bounded verification is cubic-representable for each fixed bound, eliminating an unbounded existential parameter inside the cubic fragment is an additional uniformity demand.
+This closure demand is the crux of the cubic regime.
+Our guiding Ansatz is that the degree-$3$ boundary is governed by the second barrier: at degree $3$ one can represent bounded verification cleanly, but one cannot uniformly close the external bound in a manner compatible with the intended constructive semantics.
+Exposition.
+Sections 34 implement Layer I: we give explicit cubic representations of bounded traces (proof checking and machine runs).
+Section 5.1 discusses Layer II normalization and makes precise where standard packaging steps request a closure move.
+The concluding section formulates the closure principle extracted from the solver hypothesis (under the explicit seam-check of the realizability interpretation) and performs the Layer III reflection step refuting $\exists e\,\mathsf{Solve}_{3}(e)$.
+2.3 Bounded verification is cubic-representable
+Exposition.
+The role of Layer I is sharply delimited: we only represent bounded verification predicates.
+Given an external bound (proof length, runtime), brute-force simulation is decidable and yields a finite trace.
+Our technical task is to express the existence of such a finite trace by a Diophantine condition while keeping total degree $\leq 3$.
+The point of doing this is not to “solve” an unbounded problem in disguise, but to isolate where the unbounded step must occur.
+Once the bounded predicate is represented cubically, any global solver for $\mathcal{D}_{3}$ that purports to decide unbounded solvability will necessarily have to internalize (i.e.
+existentially close) the bound.
+That closure is the crux addressed in the next subsection.
+Definition 2.6 (Bounded Run Predicate).
+Let $\mathsf{Run}(e,p,t,b)$ be the primitive recursive predicate:
+“The machine with code $e$ halts on input $p$ within $t$ steps and outputs $b\in\{0,1\}$.”
+We treat $(e,p,t,b)$ as external numerical parameters in the bounded verification problem.
+Theorem 2.7 (Cubic Representability of Bounded Machine Runs).
+There is a computable procedure which, given $(e,p,t,b)$, outputs a finite Diophantine system
+$\Phi_{e,p,t,b}(\vec{z})=0$ (2.7)
+over $\mathbb{N}$ such that:
+  1. (i)
+    every equation in $\Phi_{e,p,t,b}$ has total degree $\leq 3$, and
+  2. (ii)
+    $\exists\vec{z}\in\mathbb{N}^{m(e,p,t,b)}\;\Phi_{e,p,t,b}(\vec{z})=0\quad\Longleftrightarrow\quad\mathsf{Run}(e,p,t,b).$ (2.8)
+Ansatz 2.8.
+Fix $(e,p,t,b)$.
+A bounded computation can be written as a finite tableau (configurations at times $0,1,\ldots,t$) together with local transition constraints.
+We encode the tableau using low-overhead storage (e.g.
+$\beta$-style remainder storage for finite sequences) and enforce the transition relation by polynomial equalities whose arities are fixed and local.
+Inequalities and boundedness constraints are converted to equations by standard sum-of-squares mechanisms.
+Degree inflation is avoided by a guard economy: only those constraints that require positivity are guarded, while already-activated or intrinsically bounded equations are left unguarded.
+Remark.
+Boundedness is essential.
+The system $\Phi_{e,p,t,b}$ is constructed with $t$ fixed externally.
+Thus Theorem 2.7 yields a family of cubic instances indexed by the bound.
+This is not yet a representation of $\exists t\,\mathsf{Run}(e,p,t,b)$, and no such unbounded closure is attempted in Layer I.
+Definition 2.9.
+Fix a recursively axiomatized theory $\mathrm{T}$ with a fixed proof system and Gödel Numbering. Write $\mathsf{Prf}_{\mathrm{T}}(\pi,\varphi)$ for the primitive recursive relation expressing:
+$\pi$ codes a valid $\mathrm{T}$-proof of $\varphi$.”
+Define the bounded provability predicate
+$\mathsf{Prov}_{\mathrm{T},\leq N}(\varphi)\;\Longleftrightarrow\;\exists\pi\;\bigl(|\pi|\leq N\ \wedge\ \mathsf{Prf}_{\mathrm{T}}(\pi,\varphi)\bigr),$ (2.9)
+where $N$ is treated as an external numerical parameter.
+Theorem 2.10 (Cubic Representability of Bounded Proof Verification).
+For every fixed $(\mathrm{T},\varphi,N)$ there is an effectively constructible finite Diophantine system
+$\Psi_{\mathrm{T},\varphi,N}(\vec{x})=0$ (2.10)
+over $\mathbb{N}$ such that:
+  1. (i)
+    every equation in $\Psi_{\mathrm{T},\varphi,N}$ has total degree $\leq 3$,
+  2. (ii)
+    and,
+    $\exists\vec{x}\in\mathbb{N}^{m(\mathrm{T},\varphi,N)}\;\Psi_{\mathrm{T},\varphi,N}(\vec{x})=0\quad\Longleftrightarrow\quad\mathsf{Prov}_{\mathrm{T},\leq N}(\varphi).$ (2.11)
+Exposition.
+The system $\Psi_{\mathrm{T},\varphi,N}$ expresses the existence of a length-$N$ proof trace by:
+  1. (i)
+    storing the finite sequence of line-codes using $\beta$-style sequence storage,
+  2. (ii)
+    enforcing local verification constraints (axiomhood / inference steps) by low-degree equations controlled by Boolean activations, and
+  3. (iii)
+    using a carryless normal-form discipline for additive bookkeeping so that value-equalities can be imposed without explicit carry propagation gadgets.
+All constraints remain within degree $3$ by construction.
+Remark.
+Theorem 2.10 is a bounded statement.
+It yields a family $(\Psi_{\mathrm{T},\varphi,N})_{N\in\mathbb{N}}$ of cubic instances.
+Layer I does not attempt to eliminate $N$ and does not claim a single cubic equation equivalent to $\mathrm{T}\vdash\varphi$.
+The passage from bounded verification to unbounded existence is exactly the closure step isolated in the next subsection.
+Remark.
+Throughout Layer I we maintain the following working invariant: constraints are guarded only when a positivity or nontrivial implication direction must be enforced. Equations that are already:
+  1. (i)
+    Boolean-bounded (e.g. b(b1)=0b(b-1)=0), or
+  2. (ii)
+    activated (multiplied) by a Boolean selector, or
+  3. (iii)
+    subordinate to a globally guarded block, are left unguarded.
+This reduces notational overhead without changing satisfiability over $\mathbb{N}$ and preserves the degree bound.
+2.4 Declarative Arithmetic
+Exposition.
+A recurring source of degree growth in Diophantine encodings is the temptation to implement arithmetic procedurally at the digit level.
+In standard base-$p$ representations, verifying an equality such as $x+y=z$ by digit-by-digit computation introduces explicit carry variables and normalization constraints; these quickly force higher-degree polynomials once one insists on uniformity.
+Our approach is instead declarative.
+We enforce arithmetic equalities at the value level (e.g.
+$z=x+y$), and separately enforce that each integer appearing in the encoding has a canonical digit normal form subject only to local constraints.
+Uniqueness of normal forms then pins down the intended digits without any explicit carry propagation gadget.
+Existential quantification performs the “guessing”; the equations perform the verification.
+Proposition 2.11 (Declarative Addition Principle).
+Let $\mathcal{R}$ be a numeral system assigning to each $n\in\mathbb{N}$ a digit vector $\mathrm{dig}(n)$, such that:
+  1. (i)
 […]
-The restriction to $\mathbb{N}$ rather than $\mathbb{Z}$ is essential.
-Cubic equations over $\mathbb{Z}$ admit decidability through Thue equations (Baker’s transcendence methods), Hasse principle obstructions, and elliptic curve descent via Mordell–Weil theory.
-Undecidability over $\mathbb{N}$ exploits combinatorial complexity introduced by non-negativity constraints, enforced polynomially through guard-gadgets: replacing constraints $E(\mathbf{x})=0$ of degree at most two with the guarded system $u\cdot E(\mathbf{x})=0$, $u=1+v^{2}$ maintains cubic degree while forcing $u\geq 1$.
-Naive activation variables or difference-of-squares encodings produce degree-four terms, necessitating this construction.
-Stage Configuration Implication
-Step 1 $\mathcal{A}_{1}$ exists $\delta=2$ decidable (four-squares)
-Step 2 $\mathcal{A}_{1}\land\mathcal{A}_{2}$ exist Can reduce any $\delta=6$ to equivalent $\delta=3$
-Step 3a Assume $\mathcal{A}_{3}$ exists Contradiction: $\mathcal{A}_{1}\land\mathcal{A}_{2}\land\mathcal{A}_{3}$ decides PA
-Step 3b Reject hypothesis Conclusion: $\delta=3$ undecidable over $\mathbb{N}$
-Table 1: Logical structure of the impossibility proof.
-The technical realization is a witness of a single cubic polynomial $P_{\text{hard}}(\mathbf{x})$ with $m=O(K^{3}N^{3})$ variables, where $N$ is the proof length and $K=O(\log(\max_{i}f_{i}))$ bounds Zeckendorf digit count.
-Every intermediate constraint—after guard-wrapping—has degree $\leq 3$ (by explicit bookkeeping, see 2); sum-of-squares merging yields degree $\leq 6$, reduced to cubic via monomial shielding (Definition 6.1) with $O(K^{3}N^{3})$ auxiliary variables
-$P_{\text{merged}}=\sum_{i}P_{i}^{2}\in\delta(6)$ (2.3)
-followed by recursive monomial shielding: each monomial $x^{a}y^{b}z^{c}$ with $a+b+c>3$ factors as $p\cdot q$ with $\delta(p),\delta(q)\leq 3$, introducing auxiliary $w$ via $w=p\cdot q$.
-Iteration terminates in $O(1)$ rounds, yielding a single cubic polynomial $P_{\text{hard}}(\mathbf{x})$ with $O(K^{3}N^{3})$ variables.
-The variable inflation from systems to single equations reflects the cost of degree-preserving aggregation.
-Our proof by contradiction states: No universal cubic equation $P_{\text{univ}}(n,\mathbf{x})$ can encode all halting problems: such a polynomial would render Chaitin [1975] constant
-$\Omega=\sum_{M\text{ halts}}2^{-|M|}$ (2.4)
-computable by enumeration.
-Our construction avoids this through non-uniformity: $P_{\text{hard}}$ depends on the uncomputable proof length $N$.
-This applies equally to systems and single equations, though systems admit sparse structure while single equations become dense under sum-of-squares.
+  2. (ii)
+    Uniqueness, each $n$ has a unique valid digit vector.
+Then, for arithmetic constraints used inside a Diophantine encoding, it is not necessary to implement digit-level addition with carry gadgets.
+One may impose the value equality (e.g.
+$z=x+y$) and require valid digit vectors for $x,y,z$; uniqueness then forces $\mathrm{dig}(z)$ to coincide with the canonical normal form of $x+y$.
+Remark.
+Proposition 2.11 is not a claim about computability of addition in $\mathcal{R}$ (which may be expensive).
+It is a claim about verification: the solver supplies candidate digits existentially, and the constraint system certifies that these digits form a valid normal form for the integer already fixed by value-level equations.
+Exposition.
+We instantiate the above using Zeckendorf Representation. Write $F_{1}=F_{2}=1$ and $F_{k+2}=F_{k+1}+F_{k}$. Every $n>0$ has a unique representation
+$n=\sum_{\kappa\in S}F_{\kappa},\quad S\subseteq\{2,3,\ldots\},\quad\text{with no consecutive indices.}$ (2.12)
+Fix a cutoff $K$ (chosen externally in bounded settings), introduce digits $d_{2},\ldots,d_{K}\in\mathbb{N}$, and enforce:
+$d_{\kappa}-d_{\kappa}^{2}=0,\quad d_{\kappa}d_{\kappa+1}=0,\quad n=\sum_{\kappa=2}^{K}F_{\kappa}d_{\kappa}.$ (2.13)
+These are quadratic constraints. In particular, Zeckendorf Representation is verified without carry propagation.
+Remark.
+In base-$p$ systems, normal form is not stable under addition without explicit carry variables.
+By contrast, Zeckendorf Representation is enforced by a local forbidden-pattern condition (no adjacent $1$’s) plus reconstruction.
+When we impose a value-level equality $z=x+y$, the digits of $z$ are not computed; they are selected existentially and then uniquely determined by the normality constraints.
+Exposition.
+In Theorems 2.7 and 2.10 we repeatedly face constraints of the form
+$u=u_{1}+u_{2},\quad u=u_{1}+u_{2}+u_{3},$ (2.14)
+arising from local transition rules, pairing functions, or proof-rule bookkeeping.
+The naive approach would encode these equalities at the digit level (introducing carries) to maintain boundedness.
+We instead enforce the value equalities directly and require each relevant integer to carry a form in Zeckendorf Representation.
+By uniqueness, this pins down the digits without introducing global carry-propagation constraints and keeps the degree within the cubic regime once guards and activations are accounted for.
+Remark.
+In Layer I all digit systems have externally fixed width (cutoff $K$), because we only represent bounded traces.
+The dependence of $K$ on the size of the bounded objects is harmless there: $K$ is part of the externally supplied bound data.
+The unbounded elimination of such parameters is not attempted in Layer I; it is precisely the closure issue isolated next.
+2.5 The cubic closure problem
+Exposition.
+Layer I yields cubic representations of bounded verification predicates.
+The global solver hypothesis, however, speaks about unbounded existence: whether there exists some witness (some runtime, some proof length, some tableau size) satisfying the bounded predicate.
+Thus the decisive step is the passage from a family of bounded instances
+$(\Phi_{e,p,t,b})_{t\in\mathbb{N}}\quad\text{representing}\quad\mathsf{Run}(e,p,t,b)$ (2.15)
+to a single cubic instance representing the unbounded existential closure
+$\exists t\;\mathsf{Run}(e,p,t,b).$ (2.16)
+This is the point at which any putative uniform solver for $\mathcal{D}_{3}$ must supply a uniform witness-selection/packaging principle.
+Definition 2.12 (Existential Closure of Bounded Runs).
+For fixed $(e,p,b)$ define the $\Sigma_{1}$ predicate
+$\mathsf{HaltsOut}(e,p,b)\;:\Longleftrightarrow\;\exists t\;\mathsf{Run}(e,p,t,b).$ (2.17)
+Classically, $\mathsf{Solve}_{3}(e)$ asserts that for each input $p$ there exists a bit $b$ such that $\mathsf{HaltsOut}(e,p,b)$ holds, and that $b$ agrees with $\mathsf{Sol}_{3}(p)$.
+Remark.
+The bounded predicate $\mathsf{Run}(e,p,t,b)$ is decidable for fixed $t$.
+The unbounded closure $\exists t\,\mathsf{Run}(e,p,t,b)$ is exactly where non-uniformity enters: it asserts existence of a terminating trace without supplying a bound on where termination occurs.
+In realizability terms, it demands a witness-producing procedure, not merely a classical truth value.
+Definition 2.13 (Cubic Closure Operator).
+A cubic closure operator is a computable function $\mathsf{Close}$ which, given $(e,p,b)$, outputs a code
+$q=\mathsf{Close}(e,p,b)\in\mathrm{Code}(\mathcal{D}_{3})$ (2.18)
+such that
+$\mathsf{HaltsOut}(e,p,b)\quad\Longleftrightarrow\quad\mathsf{Sol}_{3}\bigl(\mathsf{Close}(e,p,b)\bigr).$ (2.19)
+Exposition.
+One may equivalently regard $\mathsf{Close}$ as a uniform method for packaging a bounded-verification family (whose instance size depends on the bound $t$) into a single cubic Diophantine input while preserving solvability over $\mathbb{N}$.
+In this form, the closure problem is visibly orthogonal to the local “gadgetry” of Layer I: even perfect degree-$3$ bounded encodings do not by themselves provide a uniform elimination of the external bound.
+Proposition 2.14 (Semantic Consequence).
+Assume there exists $e$ such that $\mathsf{Solve}_{3}(e)$ holds. Then there exists a cubic closure operator $\mathsf{Close}$ satisfying (2.19).
+Consider the predicate
+$\mathsf{HaltsOut}(e,p,b)\equiv\exists t\,\mathsf{Run}(e,p,t,b)$ (2.20)
+By Theorem 2.7, for each fixed $t$ there is a cubic system $\Phi_{e,p,t,b}=0$ whose solvability is equivalent to $\mathsf{Run}(e,p,t,b)$.
+The global solver hypothesis supplies a uniform terminating computation witnessing the correct yes/no decision for every cubic input.
+In particular, it supplies (uniformly in $(e,p,b)$) the computational content needed to compile the existential closure over $t$ into a single cubic solvability query: a code $q$ such that $q$ is solvable iff some $t$ yields a solvable bounded-run instance.
+Taking $\mathsf{Close}(e,p,b)\equiv q$ yields (2.19).
+∎
+Remark.
+Proposition 2.14 is not an algebraic lemma about merging polynomials.
+It is a uniformity extraction: it isolates the implicit closure principle any total correct solver must realize in order to be globally correct on a class whose defining predicate already contains an unbounded existential quantifier.
+The remainder of the argument rejects the existence of such a closure operator at degree $3$ on realizability/non-uniformity grounds, and then applies a classical reflection step to conclude $\neg\exists e\,\mathsf{Solve}_{3}(e)$.
+2.6 Constructive Reading
+Exposition.
+The closure principle isolated in Section 2.5 is not merely a classical equivalence between predicates.
 […]
-The cubic barrier represents a genuine phase transition in Diophantine complexity over $\mathbb{N}$, arising from arithmetic structure, non-negativity, and computational irreducibility: systems achieve cubic degree through encoding design, while single equations require degree-reduction post-processing.
-Figure 1: A high-level overview of the reduction from provability in Peano Arithmetic to the solvability of cubic Diophantine systems.
-The construction proceeds through seven stages, each maintaining polynomial bounds on variable count and equation degree.
-2.2.1 Encoding Proof Sequences (Stages 1–3)
-A proof $\pi=\langle\varphi_{1},\ldots,\varphi_{N}\rangle$ is encoded via Gödel numbers $f_{i}\in\mathbb{N}$. The sequence is stored using two complementary representations:
+has computational content: it demands a realizer, i.e. an effective procedure which, given $(e,p,b)$, produces a code
+$q=\mathsf{Close}(e,p,b)$ (2.22)
+together with a correctness witness for (2.19).
+In other words, the closure principle is a uniform program-existence claim.
+This is precisely the sort of uniform witness-selection principle that realizability makes logically visible.
+One is not entitled to trade an external existential quantifier (“there exists some tt”) for an internal cubic solvability predicate without exhibiting a uniform method that performs this compilation.
+Definition 2.15 (Uniform Cubic Closure Schema).
+Let $\mathsf{Close}$ range over total computable functions (realizers). The uniform cubic closure principle is the schema:
+$\exists\mathsf{Close}\;\forall e,p,b\;\Bigl(\mathsf{HaltsOut}(e,p,b)\ \leftrightarrow\ \mathsf{Sol}_{3}(\mathsf{Close}(e,p,b))\Bigr).$ (2.23)
+Remark.
+In classical logic, (2.23) merely asserts existence of a function with a correctness property.
+In the constructive layer, it is stronger: it asserts existence of a program implementing $\mathsf{Close}$.
+The subsequent refutation therefore targets the realizability content of (2.23), not merely its classical truth value.
+Exposition.
+If a uniform cubic closure operator existed, then unbounded halting-output predicates of the form
+$\exists t\;\mathsf{Run}(e,p,t,b)$ (2.24)
+could be uniformly reduced to cubic solvability queries. Since $\mathsf{Sol}_{3}$ is exactly the predicate decided by a hypothetical cubic solver, the combination
+$\text{(cubic solver)}\ +\ \text{(cubic closure)}$ (2.25)
+would yield a uniform decision procedure for a class of $\Sigma_{1}$ predicates sufficient to support diagonalization.
+Conceptually: bounded verification is decidable because the witness bound is fixed; closure converts “there exists a witness” into a single bounded-format question in $\mathcal{D}_{3}$.
+This is a global stabilization of $\Sigma_{1}$ existence in the cubic fragment.
+Remark.
+Gödelian non-uniformity is precisely the failure of uniform methods that decide or select witnesses for unbounded $\Sigma_{1}$ claims in general.
+The closure operator is such a witness-selection principle as it turns an unbounded existential over traces into a single cubic solvability instance, uniformly in parameters.
+Thus, at the constructive level, the closure principle is rejected.
+Thesis 2.16.
+There is no realizable (computable) function $\mathsf{Close}$ satisfying (2.23). Equivalently, the uniform cubic closure schema (2.23) admits contradiction in $\mathsf{HA}$.
+Assume, for contradiction, that such a realizer $\mathsf{Close}$ exists. Then for each $(e,p,b)$ we can compute a cubic code $q=\mathsf{Close}(e,p,b)$ such that
+$\exists t\;\mathsf{Run}(e,p,t,b)\ \leftrightarrow\ \mathsf{Sol}_{3}(q).$ (2.26)
+If, in addition, we had any total correct solver for $\mathsf{Sol}_{3}$, we would obtain a uniform method deciding
+$\exists t\,\mathsf{Run}(e,p,t,b)$ (2.27)
+for the relevant class of
+$(e,p,b),$ (2.28)
+which suffices (by standard diagonalization) to collapse Gödelian non-uniformity.
+Since such a collapse is incompatible with Incompleteness for sufficiently strong effective theories, the assumption that a realizable $\mathsf{Close}$ exists must be rejected.
+In particular, the failure is not algebraic but logical: $\mathsf{Close}$ would implement a uniform existential-elimination principle inside the cubic fragment. ∎
+Remark.
+The conclusion is intentionally stated at the level of realizability (program existence), not as an internal arithmetical theorem of $\mathsf{HA}$.
+The paper does not claim that $\mathsf{HA}$ proves $\neg\exists\mathsf{Close}$ as a closed formula.
+Rather, the claim is that (2.23) has no realizer.
+2.7 Classical reflection and modus tollens
+Exposition.
+Layer III is a classical metatheoretic reading (e.g.
+in $\mathsf{PA}$) which treats the solver hypothesis as a definite arithmetical commitment.
+For fixed $e$, the sentence $\mathsf{Solve}_{3}(e)$ is classically determinate.
+The bounded encodings of Layer I are used only as representations of verification predicates, and the constructive analysis of Section 2.6 is used only to identify which uniform program-existence claims are untenable.
+The reflection step required is minimal: we appeal to the soundness of the constructive reading for the relevant program-existence assertions (equivalently, to the metatheoretic correctness of the realizability interpretation for $\mathsf{HA}$ in this fragment).
+This allows us to pass from “no realizer exists” to the classical negation of the corresponding existence claim.
+There is a useful formal analogue between (i) algebraic degree-preserving aggregation (Section 6.1) and (ii) the reflective step used here.
+Algebraically, maintaining $\delta\leq 3$ under aggregation forces arity growth: high-degree monomials are eliminated by introducing auxiliary variables via recursive shielding, replacing
+$m=p\cdot q\;(\text{with }\delta(p),\delta(q)\leq 3)$ (2.29)
+by a fresh variable $w$ constrained by $w=p\cdot q$. Iteration preserves degree but inflates the variable set.
+Logically, maintaining the bounded $\Sigma_{1}$ format of verification predicates forces an analogous uniformity growth: bounded run instances
+$\Phi_{e,p,t,b}=0$ (2.30)
+represent $\mathsf{Run}(e,p,t,b)$ for each fixed $t$, but the unbounded existential closure
+$\exists t\,\mathsf{Run}(e,p,t,b)$ (2.31)
+can be internalized in the cubic fragment only by a uniform compilation principle (a closure operator) packaging the family $(\Phi_{e,p,t,b})_{t\in\mathbb{N}}$ into a single cubic instance.
+Thus, the same “cost” appears in two guises:
+$\text{degree preservation}\ \leadsto\ \text{variable inflation}$ (2.32)
+and
+$\text{bounded verification}\ \leadsto\ \text{uniform closure demand}.$ (2.33)
+In particular, any classical refutation of the uniform closure demand must quantify globally over all bounded instances; this is a compactness-like move at the metalevel, precisely mirroring the arity blow-up required by degree-preserving aggregation at the object level.
+Viewed informationally, such a global treatment amounts to extracting uniform content across all bounds, a pattern characteristic of incompleteness-by-complexity phenomena, cf.
+Chaitin [1975].
+Proposition 2.17 (Solver Forces Closure).
+For any $e$, if $\mathsf{Solve}_{3}(e)$ holds, then there exists a cubic closure operator $\mathsf{Close}$ satisfying (2.19) uniformly in $(e,p,b)$.
+Remark.
+Proposition 2.17 is the uniformity-extraction step: it isolates the closure principle as an implicit consequence of total correctness of a solver for $\mathcal{D}_{3}$.
+Theorem 2.18 (No Total Correct Cubic Solver).
+$\neg\exists e\;\mathsf{Solve}_{3}(e).$ (2.34)
+By Modus Tollens, assume for contradiction that $\exists e\,\mathsf{Solve}_{3}(e)$ and fix such an $e$.
+By Proposition 2.17, $\mathsf{Solve}_{3}(e)$ entails the existence of a uniform cubic closure operator $\mathsf{Close}$ satisfying (2.19).
+But by Thesis 2.16, no realizable $\mathsf{Close}$ exists.
+By the reflection principle invoked above (soundness of the constructive reading for program-existence claims), we may therefore classically conclude $\neg\exists\mathsf{Close}$ with the required correctness property, contradicting the consequence of $\mathsf{Solve}_{3}(e)$.
+Hence $\neg\mathsf{Solve}_{3}(e)$ for all $e$, i.e.
+$\neg\exists e\,\mathsf{Solve}_{3}(e)$.
+∎
+Logic 2.19.
+Hilbert’s Tenth Problem is a uniform decision problem.
+Theorem 2.18 negates precisely the global solver hypothesis for the cubic fragment $\mathcal{D}_{3}$.
+No claim is made that the refutation is formalizable in $\mathsf{HA}$; the argument is metatheoretic and uses a classical reflection step.
+Equally, the argument does not rest on exhibiting a particular “mysterious” cubic equation, nor on constructing a universal cubic polynomial in the MRDP sense.
+The obstruction is logical: any solver would have to realize an unbounded existential-closure principle in the cubic regime, and this principle is shown to be inconsistent at the constructive level.
+3 Preliminaries
+3.1 encoding of finite sequences
+Exposition.
+The $\beta$-method stores a finite sequence $(a_{0},\dots,a_{n})$ inside a single integer $c$ by arranging congruences with pairwise coprime moduli.
+In our Diophantine systems we will not treat “Chinese Remainder” as a primitive operation; instead we encode it by the usual quotient–remainder equations plus an explicit bound.
+Theorem 3.1 ($\beta$-encoding of a finite sequence).
+Let $a_{0},a_{1},\ldots,a_{n}\in\mathbb{N}$ be given. There exist $c,d\in\mathbb{N}$ such that for every $i\in\{0,1,\ldots,n\}$ the remainder of dividing $c$ by
+$m_{i}\;\equiv\;1+(i+1)d$ (3.35)
+is exactly $a_{i}$. Equivalently, for each $i$ there exist $q_{i}\in\mathbb{N}$ with
+$c\;=\;q_{i}\,m_{i}\;+\;a_{i}\quad\text{and}\quad 0\leq a_{i}<m_{i}.$ (3.36)
+Choose $d$ large enough that $d>\max_{i}a_{i}$ and divisible by $\mathrm{lcm}(1,2,\ldots,n+1)$.
+Then $a_{i}<m_{i}$ for all $i$, and the moduli $m_{i}=1+(i+1)d$ are pairwise coprime.
+Indeed, if a prime $p$ divides $m_{i}$ and $m_{j}$ then it divides their difference (ji)d(j-i)d.
+If $p\mid d$ then $m_{i}\equiv 1\pmod{p}$, impossible; hence $p\nmid d$, so $p\mid(j-i)\leq n$, but then $p\mid\mathrm{lcm}(1,\ldots,n+1)\mid d$, contradiction.
+By the Chinese remainder theorem, there exists $c$ with $c\equiv a_{i}\pmod{m_{i}}$ for all $i$, which yields (4.64).
+∎
+Remark.
+In the construction later, the values $a_{i}$ are themselves variables (they will be the Gödel Numbers $f_{i}$ of proof lines).
+Thus we implement $\beta(c,d,i)=f_{i}$” by introducing quotient variables $q_{i}$ and enforcing
+$c=q_{i}(1+(i+1)d)+f_{i}\quad\text{together with}\quad 0\leq f_{i}<1+(i+1)d.$ (3.37)
+The only non-equational ingredient is the inequality, which we convert to equations using the four-squares lemma below.
+3.2 Bounds as equations
+Lemma 3.2 (Four-squares bound encoding).
+For $r,m\in\mathbb{N}$ the constraint $0\leq r<m$ is equivalent to the existence of
+$a_{1},a_{2},a_{3},a_{4},b_{1},b_{2},b_{3},b_{4}\in\mathbb{N}$ (3.38)
+such that
+$r$ $=a_{1}^{2}+a_{2}^{2}+a_{3}^{2}+a_{4}^{2},$ (3.39)
+$m-1-r$ $=b_{1}^{2}+b_{2}^{2}+b_{3}^{2}+b_{4}^{2}.$ (3.40)
+In particular, the bound $0\leq r<m$ can be enforced using polynomial equations of degree $2$.
+If (4.65) holds then $r\geq 0$.
+If (4.66) holds then $m-1-r\geq 0$, hence $r\leq m-1$, i.e.
+$r<m$.
+Conversely, if $0\leq r<m$ then both $r$ and m1rm-1-r are nonnegative integers, hence each is a sum of four squares by Lagrange’s Theorem, cf.
+Lagrange [1770].
+∎
+Remark.
+For later completeness arguments it is useful that, given numerical $r$ and $m$, one can effectively find witnesses for (4.65)–(4.66).
+We only use existence and computability; we do not rely on fine-grained complexity estimates here.
+3.3 Zeckendorf digits
+Definition 3.3 (Zeckendorf representation).
+Let $F_{1}=F_{2}=1$ and $F_{k+2}=F_{k+1}+F_{k}$. Every $n\in\mathbb{N}_{>0}$ has a unique representation
+$n=\sum_{\kappa\in S}F_{\kappa},$ (3.41)
+where $S\subseteq\{2,3,4,\ldots\}$ contains no consecutive indices.
+Lemma 3.4 (Polynomial encoding of Zeckendorf digits).
+Fix a cutoff $K\geq 2$. To express that $n$ is represented using indices $\leq K$, introduce variables $d_{2},\ldots,d_{K}\in\mathbb{N}$ and impose:
   1. 1.
-    $\beta$-Function Storage (Theorem 3.1): Existentially quantify over parameters $c,d\in\mathbb{N}$ satisfying
-    $\beta(c,d,i)=f_{i}\quad\text{for all }i\in\{1,\ldots,N\},$ (2.5)
-    where $\beta(c,d,i)$ denotes the remainder when dividing $c$ by $1+(i+1)d$. This is enforced via quadratic constraints (Lemma 4.1):
-    $c$ $=q_{i}\cdot(1+(i+1)d)+r_{i},$ (2.6)
-[…]
+    $d_{\kappa}-d_{\kappa}^{2}=0$ for all $\kappa\in\{2,\ldots,K\}$,
   2. 2.
-    Zeckendorf Digits (Definition 3.3): Each $f_{i}$ is represented as a sum of non-consecutive Fibonacci numbers:
-    $f_{i}=\sum_{\kappa=2}^{K}F_{\kappa}\cdot d_{i,\kappa},\quad d_{i,\kappa}\in\{0,1\},\quad d_{i,\kappa}\cdot d_{i,\kappa+1}=0.$ (2.8)
-    This representation is unique and enables arithmetic verification without carry propagation (Exposition Exposition).
-2.2.2 Proof Verification (Stage 4)
-Each line $i$ must be justified either as an axiom or by modus ponens from earlier lines. This disjunction is encoded via boolean activation variables:
+    $d_{\kappa}d_{\kappa+1}=0$ for all $\kappa\in\{2,\ldots,K-1\}$,
+  3. 3.
+    $n=\sum_{\kappa=2}^{K}F_{\kappa}d_{\kappa}$.
+All constraints have degree $\leq 2$.
+Over $\mathbb{N}$, $d_{\kappa}-d_{\kappa}^{2}=0$ forces $d_{\kappa}\in\{0,1\}$. The products enforce non-adjacency. The reconstruction is linear in the digits. ∎
+Remark.
+In base-$p$ digit encodings, verifying addition typically requires explicit carry variables and higher-degree constraints.
+With Zeckendorf Digits, we instead enforce the value equality (e.g.
+$f_{i}=f_{j}+f_{k}$) and separately require that $f_{i}$ has a valid Zeckendorf Digit vector.
+Uniqueness of Zeckendorf Representation then pins down the digits of the sum without introducing a carry-propagation gadget.
+3.4 Carryless pairing
+Definition 3.5 (Support, rank, delimiter).
+Let $n\in\mathbb{N}$ have Zeckendorf Digits $(d_{\kappa})_{\kappa\geq 2}$ as in Lemma 3.4, and define its support
+$Z(n)\;\equiv\;\{\kappa\geq 2:d_{\kappa}=1\}.$ (3.42)
+Define the rank $r(n)$ by
+$r(n)\;\equiv\;1+\max\bigl(Z(n)\cup\{1\}\bigr),$ (3.43)
+so $Z(n)\subseteq\{2,3,\ldots,r(n)-1\}$. Define the delimiter $B(n)\equiv 2r(n)$.
+Definition 3.6 (Typed carriers).
+For $x\in\mathbb{N}$ define the finite carriers
+$\tau_{x}=\{\,e\in\mathbb{N}:e<r(x)\,\},\quad\varepsilon_{x}=\{\,2e:e\in\tau_{x}\,\},$ (3.44)
+and for $y\in\mathbb{N}$ the odd band
+$\omega_{x,y}=\{\,B(x)+(2j-1):j<r(y)\,\}.$ (3.45)
+Lemma 3.7 (Typed finiteness).
+For each $x,y\in\mathbb{N}$ the sets $\tau_{x}$, $\varepsilon_{x}$, and $\omega_{x,y}$ are finite. Their bounds depend only on $r(x)$ and $r(y)$, fixed from the Zeckendorf Supports of $x$ and $y$.
+Immediate from the definitions: each set is cut off by $r(x)$ and/or $r(y)$. ∎
+Definition 3.8 (Typed carryless pairing).
+The Typed Carryless Pairing defined as
+$\pi_{\mathrm{CL}}^{\tau}(x,y)=\!\!\!\!\!\sum_{e\in Z(x)\subseteq\tau_{x}}\!\!\!\!\!F_{2e}\;+\;\!\!\!\!\!\sum_{j\in Z(y)\subseteq\tau_{y}}\!\!\!\!\!F_{\,B(x)+(2j-1)}.$ (3.46)
+The even band encodes $x$ and the odd band encodes $y$, separated by the delimiter $B(x)=2r(x)$ of Definition 3.5.
+Exposition.
+The carriers supply the finite index domains across which sums, tests, and bounded recursions occur.
+In particular, once $x$ and $y$ are fixed, all indices used in $\pi_{\mathrm{CL}}^{\tau}(x,y)$ lie in the finite set
+$\varepsilon_{x}\;\cup\;\omega_{x,y},$ (3.47)
+and no bound depends on intermediate values discovered during evaluation.
+This is the formal sense in which the pairing is “carryless”: the encoding uses disjoint digit bands rather than arithmetic carry propagation.
+3.5 Combined encoding
+Definition 3.9 (Guarded Constraint).
+For any polynomial $E(\mathbf{x})$ of degree $\delta\leq 2$, introduce guard variable $u$ and slack variable $v\in\mathbb{N}$. The guarded constraint is the system:
+$u\cdot E(\mathbf{x})$ $=0,$ (3.48)
+$u-1-v^{2}$ $=0.$ (3.49)
+Degree bound: If $\delta\leq 2$, then $\delta(u\cdot E)\leq 1+2=3$ and $\delta(u-1-v^{2})=2$.
+Proposition 3.10 (Guard Semantics).
+The guarded constraint enforces:
   1. 1.
-    Axiom Test (Lemma 4.5): For a finite axiom set $\{g_{1},\ldots,g_{M}\}$, the constraint
-    $b_{ax,i}\cdot\sum_{\ell=1}^{M}(f_{i}-g_{\ell})^{2}=0,\quad b_{ax,i}\in\{0,1\}$ (2.9)
-    forces $f_{i}\in\{g_{1},\ldots,g_{M}\}$ when $b_{ax,i}=1$ (since a sum of squares vanishes iff each term vanishes).
+    $u=1+v^{2}\geq 1$ for all solutions (from equation (3.49)),
   2. 2.
-    Modus Ponens Test (Lemma 4.7): For each triple $(i,j,k)$ with $j,k<i$, the constraint
-    $b_{mp,i,j,k}\cdot(f_{i}-f_{j}-f_{k})=0,\quad b_{mp,i,j,k}\in\{0,1\}$ (2.10)
-    forces $f_{i}=f_{j}+f_{k}$ when $b_{mp,i,j,k}=1$. The Zeckendorf digits of $f_{i}$ automatically satisfy the non-adjacency constraint by uniqueness (Remark Exposition).
-  3. 3.
-    Exactly-One Justification (Lemma 4.11): The constraint
-    $b_{ax,i}+\sum_{j,k<i}b_{mp,i,j,k}=1$ (2.11)
-    ensures that exactly one justification mechanism activates per line.
-2.2.3 Degree Management (Stages 5–7)
-The proof verification constraints (Stage 4) produce equations of degree $\leq 3$, but merging them into a single equation via sum-of-squares yields degree $6$. We employ two techniques:
-  1. 1.
-    Guard-Gadgets (Definition 3.6): Each constraint $E(\mathbf{x})=0$ with $\delta(E)\leq 2$ is replaced by the guarded system
-    $u\cdot E(\mathbf{x})$ $=0,$ (2.12)
-    $u-1-v^{2}$ $=0,$ (2.13)
-    where $u,v\in\mathbb{N}$. The second equation forces $u=1+v^{2}\geq 1$, hence $u\neq 0$, so the first equation implies $E(\mathbf{x})=0$. This increases degree to at most $\max(1+2,2)=3$.
-  2. 2.
-    Jones Factorization (Theorem 5.1): After merging via sum-of-squares (yielding $P_{\text{merged}}$ of degree $6$), we apply recursive monomial shielding.
-Each monomial $x^{a}y^{b}z^{c}$ with $a+b+c>3$ is factored as $p\cdot q$ where $\delta(p),\delta(q)\leq 3$, introducing auxiliary variables $w$ via constraints $w=p\cdot q$ of degree $\leq 3$.
-This process (Algorithm 6.3) introduces $O(m\log d)$ variables for a polynomial with $m$ monomials of degree $\leq d$.
-2.3 Main Result
-Theorem 2.2 (Restatement of Theorem 5.2).
-Let $\mathrm{T}$ be a recursively axiomatizable theory with finite axiom set $\mathcal{A}$, and let $G_{T}$ be its Gödel sentence.
-There exists a polynomial $P(\mathbf{x})\in\mathbb{Z}[\mathbf{x}]$ of degree $\leq 3$ such that:
-$T\vdash G_{T}\quad\Longleftrightarrow\quad\exists\mathbf{x}\in\mathbb{N}^{m}:P(\mathbf{x})=0.$ (2.14)
-The polynomial has $m=O(K^{3}N^{3})$ variables, where $N$ is the proof length and $K=O(\log(\max_{i}f_{i}))$ bounds the Fibonacci representation size.
-2.4 On Single Cubic Equations
-While our construction produces a system of cubic equations, one might ask whether a single cubic equation could suffice. The following obstruction shows this is impossible.
-Proposition 2.3 (Impossibility of Universal Cubic Equation).
-There exists no polynomial $P_{\text{univ}}(n,\mathbf{x})\in\mathbb{Z}[n,\mathbf{x}]$ within an $\omega$-consistent formal system of degree $\leq 3$ such that for all Turing machines $M$ (encoded as $n\in\mathbb{N}$):
-$M\text{ halts}\quad\Longleftrightarrow\quad\exists\mathbf{x}\in\mathbb{N}^{k}:P_{\text{univ}}(n,\mathbf{x})=0.$ (2.15)
-Proof sketch.
-Assume such $P_{\text{univ}}$ exists.
-For each machine $M$ encoded as $n$, enumerate solutions $\mathbf{x}\in\mathbb{N}^{k}$ in order of increasing norm $|\mathbf{x}|_{1}=\sum_{i}x_{i}$.
-By the degree-$3$ constraint, the polynomial $P_{\text{univ}}(n,\mathbf{x})$ has at most $O(k^{3})$ monomials, so checking whether $P_{\text{univ}}(n,\mathbf{x})=0$ takes polynomial time in $|\mathbf{x}|_{1}$.
-For halting machines, the smallest solution $\mathbf{x}^{*}$ satisfies $|\mathbf{x}^{*}|_{1}=O(\text{poly}(n))$ (by the polynomial encoding of bounded computations).
-Thus, we could compute Chaitin [1975] constant.
-$\Omega=\sum_{M\text{ halts}}2^{-|M|}$ (2.16)
-by enumerating solutions up to polynomial size—contradicting the incomputability of $\Omega$, hence no such $P_{\text{univ}}$ exists. ∎
-Remark.
-Our construction avoids this obstruction by producing a machine-dependent system $\Phi_{M}(\mathbf{x})$.
-The proof length $N$ (and hence the variable count $m=O(K^{3}N^{3})$) is not computable from the machine description alone—this non-uniformity is essential.
-The system encodes "there exists a proof of length NN", not "the shortest proof has length NN".
-2.5 Prior Work and Roadmap
-We recall:
-  1. 1.
-    Robinson et al.
-[1961]: Showed exponential Diophantine relations are listable, reducing Hilbert’s 10th to finding a Diophantine representation of exponentiation.
-While their construction used higher degrees, our guard-gadget technique demonstrates degree $\leq 3$ suffices for universal representation over $\mathbb{N}$.
-  2. 2.
-    Jones [1982]: Achieved degree $\leq 4$ via direct MRDP reduction. Our guard-gadget technique (Definition 3.6) reduces this to degree $\leq 3$ by systematically shielding quadratic constraints.
-  3. 3.
-    Matiyasevich [1970]: Proved undecidability for degree $\leq 4$ (settling the “vanilla” Hilbert’s 10th Problem).
-Our result shows the degree barrier can be lowered by one, addressing a natural question raised by Jones [1982].
-  4. 4.
-    Hilbert’s Tenth Problem and its variants for rings, number fields, and related structures remain central questions in mathematics [Alpöge et al., 2025].
-However, from a computational perspective, structures like $\mathbb{Z}$ reduce to $\mathbb{N}$ via standard encodings.
-Mathematical results, see [Poonen, 2009], concern definability complexity in strong foundational theories, which is orthogonal to our focus on bounded provability in weak arithmetics.
-The remainder of the paper fills in the technical details:
-  1. 1.
-    Section 3 (Preliminaries): Establishes encoding tools ($\beta$-function, four-squares, Zeckendorf, guard-gadgets).
-  2. 2.
-    Section 4 (Construction): Builds the cubic system via proof sequence encoding, axiom testing, and modus ponens verification.
-  3. 3.
-    Section 5 (Aggregation): Merges the system via sum-of-squares and applies Jones factorization.
-  4. 4.
-    Section 6 (Computation): Provides algorithmic details for degree reduction.
-  5. 5.
-    Section 7 (Result): Proves soundness and completeness bidirectionally.
-Each stage is constructed to be effective (algorithmically computable), polynomial-time (in the proof parameters), and degree-preserving (maintaining $\delta\leq 3$ after Stage 5).
-3 Preliminaries
-3.1 The Gödel Beta-Function
-The following logic allows encoding arbitrary finite sequences as Diophantine constraints:
-Procedure 3.1 (Diophantine $\beta$-Function).
-By the work of Gödel [1931] (using the “Chinese remainder”) and as exposited in Matiyasevich [1993], there exists a polynomial $\beta(c,d,i)$ over $\mathbb{N}$ such that for any finite sequence $a_{0},a_{1},\ldots,a_{n}\in\mathbb{N}$, there exist $c,d\in\mathbb{N}$ with
-$\beta(c,d,i)=a_{i}\quad\text{for all }i\in\{0,1,\ldots,n\}.$ (3.17)
-Explicitly, $\beta(c,d,i)=r_{i}$, where $r_{i}$ is the remainder and $q_{i}\in\mathbb{N}$ is the quotient when dividing $c$ by $(1+(i+1)d)$:
-$c=q_{i}\cdot(1+(i+1)d)+r_{i},\quad 0\leq r_{i}<1+(i+1)d.$ (3.18)
-Remark.
-The “Chinese remainder” allows us to choose $d$ such that the moduli $\{1+(i+1)d:0\leq i\leq n\}$ are pairwise coprime. A suitable choice is
-$d=\text{lcm}(1,2,\ldots,n+1),$ (3.19)
-which ensures that for any $i\neq j$, the moduli $1+(i+1)d$ and $1+(j+1)d$ differ by (ji)d(j-i)d, and this difference shares no common factor with either modulus, since
-$1+(k+1)d\equiv 1\pmod{p}\quad\text{for all primes}\quad p\mid d\quad\text{with}\quad p\leq n+1.$
-Hence we construct $c$ via CRT to satisfy $c\equiv a_{i}\pmod{1+(i+1)d}$ for all $i\leq n$. To verify pairwise coprimality with $d=\text{lcm}(1,\ldots,n+1)$: for $i<j\leq n$, we have
-$\gcd(1+(i+1)d,1+(j+1)d)\mid\gcd(1+(i+1)d,(j-i)d).$ (3.20)
-Since $(j-i)\leq n$ and $d$ is divisible by all integers up to $n+1$, while $1+(i+1)d\equiv 1\pmod{p}$ for any prime $p\mid d$ with $p\leq n+1$, the greatest common divisor equals $1$.
-This ensures the CRT applies.
-To express $\beta(c,d,i)=f_{i}$ as polynomial constraints, we introduce variables $r_{i},q_{i}$ and enforce:
-$c$ $=q_{i}\cdot(1+(i+1)d)+r_{i},$ (3.21)
-$0$ $\leq r_{i}<1+(i+1)d.$ (3.22)
-The second challenge, after reducing $\Sigma_{1}$ logic to $\Delta_{0}$ [^4] Foundational reference by Kleene [1943] introducing concepts underlying the arithmetic hierarchy, is encoding inequality polynomially.
-A naive preliminary approach using $r_{i}=s_{i}^{2}$ to force $r_{i}\geq 0$ failed because: (i) the $\beta$-construction produces remainders $r_{i}=c\bmod(1+(i+1)d)$ that need not be perfect squares (e.g., $r_{i}=2$ is not a square), and (ii) encoding the upper bound $r_{i}<1+(i+1)d$ via $(1+(i+1)d-r_{i}-1)=t_{i}^{2}$ would impose the same impossible requirement.
-The “four-squares” encoding resolves both issues simultaneously without requiring any specific number to be a perfect square.
-Lemma 3.2 (Four-Squares Bound Encoding).
-The constraint $0\leq r<m$ over $\mathbb{N}$ is equivalent to the existence of integers $a,b,c,e,\alpha,\beta,\gamma,\delta\in\mathbb{N}$ satisfying:
-$r$ $=a^{2}+b^{2}+c^{2}+e^{2},$ (3.23)
-$m-1-r$ $=\alpha^{2}+\beta^{2}+\gamma^{2}+\delta^{2}.$ (3.24)
-These are polynomial equations of degree $2$.
-Proof.
-The “Four-Square Theorem” of Lagrange [1770] allows us to say:
-Every $n\in\mathbb{N}$ is a sum of four squares.
-Constraint (4.48) forces $r\geq 0$.
-Constraint (4.49) forces $r\leq m-1$, hence $r<m$.
-Given $r^{*}\in\{0,\ldots,m-1\}$, the four-square decompositions can be computed in polynomial time.
-By the algorithm of Rabin and Shallit [1986] we decompose an integer $n\in\mathbb{N}$ into four squares
-$n=a^{2}+b^{2}+c^{2}+e^{2}$ (3.25)
-This ensures the encoding is effective. Computable witnesses exist for any valid bound, and can be constructed in polynomial time (randomized) or quasi-polynomial time (deterministic under ERH). ∎
-Remark.
-In the completeness proof of Theorem 4.16, the witness construction algorithm (Exposition 4.9) must compute four-square decompositions for each remainder $r_{i}$ and its bound $(i+1)d-1-r_{i}$.
-By the above algorithm, this can be done in time $O(\log((i+1)d))$ per decomposition, yielding total time $O(N\cdot\log(\max_{i}f_{i}))$ for all $N$ proof lines.
-This is polynomial in the proof parameters, as required.
-3.2 Unique Representation
-Remark.
-As a “rule of thumb”: A sum satisfies Zeckendorf representation (i.e., uses non-consecutive Fibonacci numbers) when there are “gaps” in the sequence of Fibonacci numbers used.
-e.g., the tuple $\{F_{6},F_{8},F_{10}\}=\{8,21,55\}$.
-Definition 3.3 (Zeckendorf Representation).
-Every positive integer $n$ admits a unique representation as a sum of non-consecutive Fibonacci numbers:
-$n=\sum_{\kappa\in S}F_{\kappa},$ (3.26)
-where $S\subseteq\{2,3,4,\ldots\}$ contains no consecutive integers (i.e., if $\kappa\in S$, then $\kappa+1\notin S$), and $F_{k}$ denotes the $k$-th Fibonacci number with $F_{1}=F_{2}=1$ and $F_{k+2}=F_{k+1}+F_{k}$.
-Lemma 3.4 (Polynomial Encoding of Zeckendorf Digits).
-For each integer $n$ with Zeckendorf representation involving indices up to $K$, introduce binary variables $d_{2},\ldots,d_{K}\in\{0,1\}$ satisfying:
-  1. 1.
-    Boolean constraint: $d_{\kappa}-d_{\kappa}^{2}=0$ for all $\kappa\in\{2,\ldots,K\}$,
-  2. 2.
-    Non-adjacency: $d_{\kappa}\cdot d_{\kappa+1}=0$ for all $\kappa\in\{2,\ldots,K-1\}$,
-  3. 3.
-    Reconstruction: $n=\sum_{\kappa=2}^{K}F_{\kappa}\cdot d_{\kappa}$.
-These constraints are polynomial of degree at most $2$.
-Proof.
-The boolean constraint $d_{\kappa}(1-d_{\kappa})=0$ forces $d_{\kappa}\in\{0,1\}$ over $\mathbb{N}$.
-The product $d_{\kappa}d_{\kappa+1}=0$ enforces non-adjacency.
-The sum is linear in the $d_{\kappa}$ variables.
-∎
-Exposition.
-The Zeckendorf representation provides an alternative to prime-power encodings for sequences.
-We use it to enable Fibonacci addition, which can be implemented via digit-wise operations respecting the non-adjacency constraint (see Section 4.3).
-3.3 Combined Encoding
-Exposition.
-For a proof sequence $\langle f_{1},\ldots,f_{N}\rangle$ where each $f_{i}$ is a Gödel number:
-  1. 1.
-    Encode the sequence using $\beta(c,d,i)=f_{i}$ via constraint (4.47) and Lemma 3.2.
-  2. 2.
-    Represent each $f_{i}$ in Zeckendorf form using Lemma 3.4.
-This dual encoding allows:
-  1. 1.
-    Variable-length proofs (via existential quantification over $N,c,d$),
-  2. 2.
-    Arithmetic verification (via Fibonacci addition on Zeckendorf digits).
-Regarding the consistency of “Dual Encoding:” Each Gödel number $f_{i}$ is encoded twice:
-  1. 1.
-    Via the $\beta$-function: $\beta(c,d,i)=f_{i}$ (using variables $c,d,r_{i},q_{i},\ldots$),
-  2. 2.
-    Via Zeckendorf: $f_{i}=\sum_{\kappa=2}^{K}F_{\kappa}d_{i,\kappa}$ (using digit variables $d_{i,\kappa}$).
-These encodings are compatible: for any sequence $\langle f_{1},\ldots,f_{N}\rangle$, witnesses $c^{*},d^{*}$ exist (by Theorem 3.1), and each $f_{i}^{*}$ has a unique Zeckendorf representation (by Definition 3.3).
-The polynomial system constrains a single variable $f_{i}$ to satisfy both encodings simultaneously, ensuring consistency.
-3.4 Summary of Polynomial Constraints
+    $E(\mathbf{x})=0$ (forced by equation (3.48) since $u\geq 1$).
+Equation (3.49) forces $u=1+v^{2}\geq 1$. Since $u\neq 0$, equation (3.48) yields $E(\mathbf{x})=0$. ∎
+Definition 3.11 (Global Multiplier).
+To enable selective activation of guards, introduce a global multiplier $T\geq 1$ via:
+$T=1+U^{2},\quad U\in\mathbb{N}.$ (3.50)
+This ensures $T\geq 1$ for all solutions. Optionally, this can be enforced via guarded constraints.
 Constraint Expanded Equation $\boldsymbol{\delta}$ $\mathbb{N}$-Mechanism
@@
 Exactly-one $b_{ax,i}+\sum_{j,k<i}b_{mp,i,j,k}-1=0$ 1 Linear in booleans
-Target check $f_{N}-g_{T}=0$ 1 Linear equality
+Target check $f_{N}-g_{\mathrm{T}}=0$ 1 Linear equality
 Post guard-wrapping
@@
 Exactly-one guard $u_{J,i}\cdot(b_{ax,i}+\sum_{j,k}b_{mp,i,j,k}-1)=0$ 2 $u_{J,i}=1+v_{J,i}^{2}\geq 1$
-Target guard $\mathrm{T}\cdot(f_{N}-g_{T})=0$ 2 $\mathrm{T}=1+U^{2}\geq 1$
+Target guard $\mathrm{T}\cdot(f_{N}-g_{\mathrm{T}})=0$ 2 $\mathrm{T}=1+U^{2}\geq 1$
 Positivity constraints
@@
 Table 1: Complete degree bookkeeping. Base constraints have $\delta\leq 3$; guard-wrapping preserves $\delta\leq 3$ for all main constraints, with guard positivity at $\delta=2$.
-3.5 Guard-Gadget Framework
-Definition 3.5 (Boolean Constraint).
-For any variable $b$ intended to represent a boolean value, the constraint
-$b-b^{2}=0,\quad b\in\{0,1\}\subseteq\mathbb{N}$ (3.27)
-Definition 3.6 (Guarded Constraint).
-For any polynomial $E(\mathbf{x})$ of degree $\delta\leq 2$, introduce guard variable $u$ and slack variable $v\in\mathbb{N}$. The guarded constraint is the system:
-$u\cdot E(\mathbf{x})$ $=0,$ (3.28)
-$u-1-v^{2}$ $=0.$ (3.29)
-Degree bound: If $\delta\leq 2$, then $\delta(u\cdot E)\leq 1+2=3$ and $\delta(u-1-v^{2})=2$.
-Proposition 3.7 (Guard Semantics).
-The guarded constraint enforces:
-  1. 1.
-    $u=1+v^{2}\geq 1$ for all solutions (from equation (3.29)),
-  2. 2.
-    $E(\mathbf{x})=0$ (forced by equation (3.28) since $u\geq 1$).
-Proof.
-Equation (3.29) forces $u=1+v^{2}\geq 1$. Since $u\neq 0$, equation (3.28) yields $E(\mathbf{x})=0$. ∎
-Definition 3.8 (Global Multiplier).
-To enable selective activation of guards, introduce a global multiplier $T\geq 1$ via:
-$T=1+U^{2},\quad U\in\mathbb{N}.$ (3.30)
-This ensures $T\geq 1$ for all solutions. Optionally, this can be enforced via guarded constraints:
-$u_{T}\cdot(T-1-U^{2})$ $=0,$ (3.31)
-$u_{T}-1-v_{T}^{2}$ $=0,$ (3.32)
-where equation (3.32) forces $u_{T}=1+v_{T}^{2}\geq 1$, and equation (3.31) then enforces $T=1+U^{2}$.
-Proposition 3.9 (Selective Activation).
-Given a guard $u_{i}$ with $u_{i}=1+v_{i}^{2}$ (from Definition 3.6) and a global multiplier $T\geq 1$, the constraint
-$T\cdot(1-u_{i})=0$ (3.33)
-forces $u_{i}=1$ (equivalently, $v_{i}=0$).
-Proof.
-Since $T\geq 1$ and $T\cdot(1-u_{i})=0$, we have $1-u_{i}=0$, hence $u_{i}=1$. From $u_{i}=1+v_{i}^{2}$, this implies $v_{i}=0$. ∎
-Exposition.
-The selective activation mechanism is used to enforce that specific constraints hold in a solution.
-For example, in proof verification (Section 4.4), we use it to force exactly one justification (axiom or modus ponens) for each line of a proof.
-The guard-gadget framework provides a systematic method to enforce boolean constraints without degree inflation (Definition 3.5), ensure all polynomial constraints are non-negative and solvable over $\mathbb{N}$ (Proposition 3.7), and selectively activate constraints via a global multiplier (Proposition 3.9).
-All gadgets maintain degree $\leq 3$ when applied to linear or quadratic base constraints.
-This forms the technical foundation for encoding proof verification at cubic degree.
 4 Construction
-We now construct a finite system of polynomial equations of degree at most $3$ such that the system has a solution over $\mathbb{N}$ if and only if the Fibonacci variant theory $T_{\text{Fib}}$ proves the target formula $G_{T}$.
+We now construct a finite system of polynomial equations of degree at most $3$ such that the system has a solution over $\mathbb{N}$ if and only if the Fibonacci variant theory $T_{\text{Fib}}$ proves the target formula $G_{\mathrm{T}}$.
 4.1 Proof Sequence Encoding
@@
 Degree analysis: For all equations: $\delta\leq 3$.
-Exposition.
-Constraints (Lemmas 4.14.3) ensure that the sequence $\langle f_{1},\ldots,f_{N}\rangle$ is encoded via $\beta(c,d,i)=f_{i}$ (Theorem 3.1), and each $f_{i}$ has a valid Zeckendorf representation with digits $d_{i,\kappa}$ (Lemma 3.4).
-The dual encoding is consistent (Exposition).
-“Why Carry Propagation is Unnecessary.” The key insight is that Zeckendorf representation is unique (Definition 3.3).
-The constraints enforce $f_{i}=f_{j}+f_{k}$ (arithmetic sum, via Equation (4.59)), and $f_{i}=\sum_{\kappa=2}^{K}F_{\kappa}\cdot d_{i,\kappa}$ with $d_{i,\kappa}\in\{0,1\}$ and $d_{i,\kappa}\cdot d_{i,\kappa+1}=0$ (via Proposition 4.4).
-By uniqueness, there exists exactly one choice of $(d_{i,2},\ldots,d_{i,K})$ satisfying the second condition for the value $f_{i}$ determined by the first condition.
-Hence the constraints force $d_{i,\kappa}$ to be the Zeckendorf digits of $f_{j}+f_{k}$, without explicitly computing the addition “digit-by-digit”.
-A procedural approach would introduce auxiliary variables $s_{i,j,k,\kappa}$ (sum digits) and $c_{i,j,k,\kappa}$ (carries), computing $s_{\kappa}=d_{j,\kappa}+d_{k,\kappa}+c_{\kappa-1}$ and normalizing to Zeckendorf form.
-This requires degree-$4$ constraints to handle carry propagation.
-Our declarative approach avoids this by relying on existential quantification: the solver “guesses” digits $(d_{i,\kappa})$, and uniqueness ensures only the correct guess satisfies all constraints.
 4.2 Combined Proof Encoding
@@
 Degree analysis: All constraints have $\delta\leq 2$.
-Proof.
 Equations (4.64)–(4.66) encode $f_{i}=\beta(c,d,i)$ via the $\beta$-function.
 Equations (4.67)–(4.69) enforce the Zeckendorf Representation of $f_{i}$.
-Uniqueness of Zeckendorf representation ensures these constraints are consistent.
+Uniqueness ensures these constraints are consistent.
 ∎
@@
       The sum $\sum_{\ell=1}^{M}(f_{i}-g_{\ell})^{2}=0$, so equation (4.71) is satisfied.
-Proof.
-Immediate from the properties of squares over $\mathbb{N}$. ∎
-Exposition.
+Immediate from the properties of squares over $\mathbb{N}$.
 The constraint $b_{ax,i}\cdot\sum_{\ell=1}^{M}(f_{i}-g_{\ell})^{2}=0$ encodes the logical disjunction:
@@
 $\sum_{j=1}^{n}(x-a_{j})^{2}=0.$ (4.73)
-The key property is that a sum of non-negative terms vanishes if and only if each term vanishes.
-Over $\mathbb{N}$, this requires $x=a_{j}$ for some $j$.
-Degree analysis: The sum $\sum_{\ell=1}^{M}(f_{i}-g_{\ell})^{2}$ has degree $2$ (each term is a square of a linear expression).
-Multiplying by the boolean $b_{ax,i}$ increases the degree to $3$.
 4.4 Modus Ponens Verification
 We verify Modus Ponens by directly enforcing the arithmetic constraint $f_{i}=f_{j}+f_{k}$ for the appropriate triple $(i,j,k)$.
-The Zeckendorf digit constraints ensure that the digits $d_{i,\kappa}$ represent this sum uniquely, without requiring explicit carry propagation.
+The Zeckendorf constraints ensure that the digits $d_{i,\kappa}$ represent this sum uniquely, without requiring explicit carry propagation.
 Lemma 4.7 (Modus Ponens Constraint).
@@
 Lemma 4.8 (Correctness of Modus Ponens Encoding).
-The constraint system in Lemma 4.7 correctly encodes modus ponens: Soundness. If $b_{mp,i,j,k}=1$ in a solution, then, equation (4.58) necessitates
-$u_{mp,i,j,k}=1+v_{mp,i,j,k}^{2}\geq 1;$ (4.60)
-equation (4.59) with $u_{mp,i,j,k}\geq 1$ forces $f_{i}=f_{j}+f_{k}$, while proposition 4.4, the digits $d_{i,\kappa}$ satisfy:
-$f_{i}=\sum_{\kappa=2}^{K}F_{\kappa}\cdot d_{i,\kappa},\quad d_{i,\kappa}\in\{0,1\},\quad d_{i,\kappa}\cdot d_{i,\kappa+1}=0.$ (4.61)
-By uniqueness of Zeckendorf representation, the digits $d_{i,\kappa}$ are the unique Zeckendorf digits of $f_{j}+f_{k}$.
+The constraint system in Lemma 4.7 correctly encodes Modus Ponens. Soundness. If
+$b_{mp,i,j,k}=1$ (4.77)
+in a solution, then (4.75) implies
+$u_{mp,i,j,k}=1+v_{mp,i,j,k}^{2}\geq 1$ (4.78)
+Hence (4.76) forces
+$f_{i}=f_{j}+f_{k}.$ (4.79)
+Independently, Proposition 4.4 enforces that $f_{i}$ is reconstructed from Zeckendorf Representation $(d_{i,\kappa})$ with $d_{i,\kappa}\in\{0,1\}$ and $d_{i,\kappa}d_{i,\kappa+1}=0$.
+By uniqueness, these digits are exactly the Zeckendorf Digits of $f_{j}+f_{k}$.
 Completeness.
-Given a proof where line $i$ is derived from lines $j,k$ via modus ponens (so $f_{i}=f_{j}+f_{k}$ in $\mathbb{N}$): Set $b_{mp,i,j,k}=1$ and $b_{mp,i,j^{\prime},k^{\prime}}=0$ for all $(j^{\prime},k^{\prime})\neq(j,k)$.
-Compute the Zeckendorf representation of $f_{j}+f_{k}$:
-$f_{j}+f_{k}=\sum_{\kappa\in S}F_{\kappa},\quad S\subseteq\{2,\ldots,K\},\;\kappa\in S\implies\kappa+1\notin S.$ (4.62)
-We assign $d_{i,\kappa}=\mathbb{1}_{\kappa\in S}$ (indicator function).
-Set $u_{mp,i,j,k}=1$, $v_{mp,i,j,k}=0$.
-All constraints (4.57)–(4.59) are satisfied.
-See [Zeckendorf, 1972] for existence and uniqeness.
-Proof.
+Given a proof in which line $i$ is derived from lines $j,k$ by MP (so $f_{i}=f_{j}+f_{k}$), set $b_{mp,i,j,k}=1$ and all other $b_{mp,i,j^{\prime},k^{\prime}}=0$.
+Compute the Zeckendorf Digits of $f_{i}$ using Procedure 4.9 and assign $d_{i,\kappa}$ accordingly.
+Set
+$u_{mp,i,j,k}=1$ (4.80)
+and
+$v_{mp,i,j,k}=0$ (4.81)
+Then (4.74)–(4.76) are satisfied.
 By soundness and completeness.
@@
 Our key observation: Carry propagation is unnecessary.
-The constraint $f_{i}=f_{j}+f_{k}$ (linear in $f_{i},f_{j},f_{k}$) combined with the Zeckendorf digit constraints (quadratic in $d_{i,\kappa}$) suffices.
+The constraint $f_{i}=f_{j}+f_{k}$ (linear in $f_{i},f_{j},f_{k}$) combined with the Zeckendorf constraints (quadratic in $d_{i,\kappa}$) suffices.
 The existential quantification $\exists d_{i,\kappa}$ allows the constraint solver to guess the correct digits, which are then verified by:
-  1. 1.
+  1. (i)
     The reconstruction constraint $f_{i}=\sum_{\kappa}F_{\kappa}\cdot d_{i,\kappa}$ (Equation 4.69),
-  2. 2.
+  2. (ii)
     The non-adjacency constraint $d_{i,\kappa}\cdot d_{i,\kappa+1}=0$ (Equation 4.68).
@@
 Benefit: The procedural approach requires degree-$4$ constraints. The declarative approach uses only:
-  1. 1.
+  1. (i)
     Linear sum with guard: $u\cdot(f_{i}-f_{j}-f_{k})=0$ (degree $2$),
-  2. 2.
+  2. (ii)
     Quadratic digit constraints: already present in Proposition 4.4.
@@
 Procedure 4.9 (Greedy Witness Construction).
-The completeness direction of Lemma 4.8 relies on the computability of Zeckendorf representations.
-Given $n\in\mathbb{N}$, the Zeckendorf digits can be computed by the “greedy algorithm” [^5] An algorithm that makes locally optimal choices at each step, selecting the largest (or best) available option without reconsidering previous decisions.
-in $O(\log n)$ steps that produces the unique representation satisfying the non-adjacency constraint.:
+The completeness direction of Lemma 4.8 relies on the computability of “Zeckendorf representation”.
+Given $n$, the vector is computable by the Greedy Algorithm with an explicit non-adjacency skip.
+$O(\log n)$ steps that produces the unique representation satisfying the non-adjacency constraint.:
 […]
 Exposition.
-For verification, given a formal proof $\pi$ of length $N$ with Gödel numbers $f_{1},\ldots,f_{N}$, we can effectively construct a solution $\vec{x}^{*}\in\mathbb{N}^{m}$ by computing $c^{*},d^{*}$ via the $\beta$-function, computing Zeckendorf digits $d_{i,\kappa}^{*}$ for each $f_{i}$ using the greedy algorithm, setting activation variables $b_{ax,i}^{*}$ or $b_{mp,i,j,k}^{*}$ according to the justification of line $i$, and computing all guard variables $u^{*},v^{*}$ to satisfy the positivity constraints.
+For verification, given a formal proof $\pi$ of length $N$ with Gödel Numbers $f_{1},\ldots,f_{N}$, we can effectively construct a solution
+$\vec{x}^{*}\in\mathbb{N}^{m}$ (4.85)
+by computing $c^{*},d^{*}$ via the $\beta$-function, computing Zeckendorf Representation $d_{i,\kappa}^{*}$ for each $f_{i}$ using the Greedy Algorithm, setting activation variables $b_{ax,i}^{*}$ or $b_{mp,i,j,k}^{*}$ according to the justification of line $i$, and computing all guard variables $u^{*},v^{*}$ to satisfy the positivity constraints.
 This witness construction is polynomial-time in $N$ and $\max_{i}f_{i}$, demonstrating that the reduction from provability to Diophantine solvability is effective.
@@
     $b_{ax,i}=0$ and $b_{mp,i,j^{*},k^{*}}=1$ for exactly one pair $(j^{*},k^{*})$ (line $i$ is derived by MP).
-Proof.
 Since $b_{ax,i},b_{mp,i,j,k}\in\{0,1\}$ (by 4.10), equation (4.89) necessitates:
 $b_{ax,i}+\sum_{j,k<i}b_{mp,i,j,k}=1.$ (4.91)
-This is a linear Diophantine equation with boolean variables, which has a solution if and only if exactly one variable is 11 and the rest are 0. ∎
+This is a linear Diophantine equation with boolean variables, which has a solution if and only if exactly one variable is $1$ and the rest are 0. ∎
 Exposition.
@@
 Lemma 4.13 (Target Check).
-Let $G_{T}$ be the target formula with Gödel number $g_{T}$. Enforce, with $T$ is the global multiplier (Definition 3.8),
-$T\cdot(f_{N}-g_{T})$ $=0,$ (4.72)
-$T-1-U^{2}$ $=0,$ (4.73)
+Let $G_{\mathrm{T}}$ be the target formula with Gödel Number $g_{\mathrm{T}}$. Enforce, with $\mathrm{T}$ is the global multiplier (Definition 3.11),
+$\mathrm{T}\cdot(f_{N}-g_{\mathrm{T}})$ $=0,$ (4.92)
+$\mathrm{T}-1-U^{2}$ $=0,$ (4.93)
 […]
 Lemma 4.14 (Target Correctness).
-Constraint 4.13 forces $f_{N}=g_{T}$ (i.e., the final line of the proof is the target formula).
-Proof.
-Since $T=1+U^{2}\geq 1$ (by equation (4.73)), equation (4.72) forces $f_{N}=g_{T}$. ∎
+Constraint 4.13 forces $f_{N}=g_{\mathrm{T}}$ (i.e., the final line of the proof is the target formula).
+Since $\mathrm{T}=1+U^{2}\geq 1$ (by equation (4.93)), equation (4.92) forces $f_{N}=g_{\mathrm{T}}$. ∎
 4.7 Complete System Assembly
@@
 Theorem 4.16 (Finite Cubic System for Proof Verification).
-Let $T_{\text{Fib}}$ be a recursively axiomatizable theory with finite axiom set $\mathcal{A}$, and let $G_{T}$ be a sentence.
+Let $T_{\text{Fib}}$ be a recursively axiomatizable theory with finite axiom set $\mathcal{A}$, and let $G_{\mathrm{T}}$ be a sentence.
 There exists a finite system $\Phi(\mathbf{x})$ of polynomial equations over $\mathbb{N}$, each of degree at most $3$, such that:
-$T_{\text{Fib}}\vdash G_{T}\quad\iff\quad\exists\mathbf{x}\in\mathbb{N}^{m}:\Phi(\mathbf{x})=\mathbf{0}.$ (4.75)
-Proof.
+$T_{\text{Fib}}\vdash G_{\mathrm{T}}\quad\iff\quad\exists\mathbf{x}\in\mathbb{N}^{m}:\Phi(\mathbf{x})=\mathbf{0}.$ (4.95)
 By soundness and completeness. ($\Rightarrow$) Completeness:
   1. 1.
-    Suppose $T_{\text{Fib}}\vdash G_{T}$.
-Then there exists a finite proof $\pi=\langle\varphi_{1},\ldots,\varphi_{N}\rangle$ where $\varphi_{N}=G_{T}$.
+    Suppose $T_{\text{Fib}}\vdash G_{\mathrm{T}}$.
+Then there exists a finite proof $\pi=\langle\varphi_{1},\ldots,\varphi_{N}\rangle$ where $\varphi_{N}=G_{\mathrm{T}}$.
 Let $f_{i}$ be the Gödel Number of $\varphi_{i}$.
@@
   2. 2.
-    Compute the Fibonacci addition witnesses (carries $c^{*}_{i,j,k,\kappa}$, sums $s^{*}_{i,j,k,\kappa}$) Finally, compute all guard variables $u^{*}_{\bullet},v^{*}_{\bullet}$ to satisfy definitions 3.63.8.
-By construction, $\mathbf{x}^{*}=(N,c^{*},d^{*},\{f^{*}_{i}\},\ldots)$ satisfies
-    $\Phi(\mathrm{x}^{*})=\mathbf{0}.$
+    For each MP line $i$ with premises $(j,k)$, compute the Zeckendorf Digits of the value $f_{j}+f_{k}$ using Procedure 4.9 and assign $d_{i,\kappa}^{*}$ accordingly.
+No auxiliary carry or sum variables are introduced: the constraints enforce $f_{i}=f_{j}+f_{k}$ and then verify that $(d_{i,\kappa}^{*})$ is the unique Zeckendorf vector for that value (Exposition Exposition).
   3. 3.
@@
     4. (d)
-      By Constraint 4.13, $f^{*}_{N^{*}}=g_{T}$.
+      By Constraint 4.13, $f^{*}_{N^{*}}=g_{\mathrm{T}}$.
   4. 4.
-    Let $\varphi_{i}$ be the formula with Gödel number $f^{*}_{i}$. The sequence $\langle\varphi_{1},\ldots,\varphi_{N^{*}}\rangle$ is a valid proof of $G_{T}$ in $T_{\text{Fib}}$, since:
+    Let $\varphi_{i}$ be the formula with Gödel Number $f^{*}_{i}$.
+The sequence $\langle\varphi_{1},\ldots,\varphi_{N^{*}}\rangle$ is a valid proof of $G_{\mathrm{T}}$ in $\mathrm{T}_{\text{Fib}}$, since:
     1. (a)
@@
     2. (b)
-      $\varphi_{N^{*}}=G_{T}$ (by Lemma 4.14)
+      $\varphi_{N^{*}}=G_{\mathrm{T}}$ (by Lemma 4.14)
   5. 5.
-    Therefore, $T_{\text{Fib}}\vdash G_{T}$.
+    Therefore, $T_{\text{Fib}}\vdash G_{\mathrm{T}}$.
 Exposition.
@@
 5 Aggregation
+Proposition 5.1.
 We now reduce the finite system $\Phi(\mathbf{x})$ from Theorem 4.16 to a single cubic polynomial.
@@
 We reduce the system $\Phi(\mathbf{x})$ from Theorem 4.16 to a single equation via sum-of-squares.
-Degree bound. By Table 2, all polynomials $P_{i}\in\Phi$ satisfy $\delta(P_{i})\leq 3$ after guard-wrapping (Definition 3.6).
-Merger. The system $\Phi=\{P_{1}=0,\ldots,P_{m}=0\}$ is equivalent to
+Degree bound: By Table 1, all polynomials $P_{i}\in\Phi$ satisfy $\delta(P_{i})\leq 3$ after guard-wrapping (Definition 3.9). Merger: The system $\Phi=\{P_{1}=0,\ldots,P_{m}=0\}$ is equivalent to
 $P_{\text{merged}}(\mathbf{x})=\sum_{i=1}^{m}P_{i}(\mathbf{x})^{2}=0,\quad\delta(P_{\text{merged}})\leq 6.$ (5.96)
-Size. With $m=O(KN^{3})$ constraints (Remark Exposition) and $s_{i}=O(K)$ monomials per $P_{i}$, we obtain $O(K^{3}N^{3})$ total monomials.
+Size: With $m=O(KN^{3})$ constraints (Remark Exposition) and $s_{i}=O(K)$ monomials per $P_{i}$, we obtain $O(K^{3}N^{3})$ total monomials.
 5.2 Degree Reduction to Cubic
@@
 $\exists\mathbf{x}:P(\mathbf{x})=0\quad\Longleftrightarrow\quad\exists(\mathbf{x},\mathbf{y}):Q(\mathbf{x},\mathbf{y})=0.$ (5.97)
-Proof sketch.
-For each monomial $\mathbf{x}^{\mathbf{a}}$ with $|\mathbf{a}|>3$, introduce $y$ via $y=\prod_{j\in S}x_{j}^{a_{j}}$ (where $|\mathbf{a}|_{S}\geq 2$), expressed as a degree-$2$ constraint.
+For each monomial $\mathbf{x}^{\mathbf{a}}$ with $|\mathbf{a}|>3$, introduce $y$ via $y=\prod_{j\in S}x_{j}^{a_{j}}$ (where $|\mathbf{a}|_{S}\geq 2$), expressed as a degree-22 constraint.
 Iterate until all monomials have degree $\leq 3$.
 See [Jones, 1982] for details.
-∎
-Applying Theorem 5.1 to $P_{\text{merged}}$ (degree $6$, $O(KN^{9})$ monomials by squaring analysis) introduces $k=O(KN^{9})$ auxiliary variables.
+∎ Applying Theorem 5.2 to $P_{\text{merged}}$ (degree 66, $O(KN^{9})$ monomials by squaring analysis) introduces $k=O(KN^{9})$ auxiliary variables.
 5.3 Main Result
 Theorem 5.3 (Cubic Encoding of Provability).
-Let $\mathrm{T}$ be a recursively axiomatizable theory with finite axioms, and $G_{T}$ its Gödel sentence. There exists a polynomial
+Let $\mathrm{T}$ be a recursively axiomatizable theory with finite axioms, and $G_{\mathrm{T}}$ its Gödel Sentence. There exists a polynomial
 $P(\mathbf{x})\in\mathbb{Z}[\mathbf{x}]$
 of degree $\leq 3$ such that:
-$T\vdash G_{T}\quad\Longleftrightarrow\quad\exists\mathbf{x}\in\mathbb{N}^{m}:P(\mathbf{x})=0.$ (5.78)
+$\mathrm{T}\vdash G_{\mathrm{T}}\quad\Longleftrightarrow\quad\exists\mathbf{x}\in\mathbb{N}^{m}:P(\mathbf{x})=0.$ (5.98)
 The polynomial has $m=O(KN^{6})$ variables and $O(K^{2}N^{12})$ monomials.
-Proof.
 By our additive approach, theorem 4.16 (finite cubic system), sum-of-squares merging via (5.96), and theorem 5.2. ∎
@@
 The solvability problem for diophantines with $\delta=3$ over $\mathbb{N}$ is undecidable.
-Proof.
 Provability in consistent recursively axiomatizable theories is incomplete by Gödel [1931]. Theorem 5.3 reduces this from $\delta\leq 4$ to $\delta\leq 3$ completeness. ∎
-Remark (Non-Uniformity).
-The polynomial $P$ depends on the (unknown) proof length $N$. This is inherent: no algorithm can compute the “correct” $P$ from $T$ and $G_{T}$ alone, as this would solve the Halting Problem.
+Remark.
+The polynomial $P$ depends on the (unknown) proof length $N$.
+This is inherent: no algorithm can compute the “correct” $P$ from $\mathrm{T}$ and $G_{\mathrm{T}}$ alone, as this would solve the Halting Problem.
 6 Computation
@@
 $Q_{E}=U\cdot E(\mathbf{x})-Z^{2}=0,\quad U,Z\in\mathbb{N}.$ (6.99)
-Then $\delta(Q_{E})=\max(1+\delta(E),2)\leq 3$ and $Q_{E}=0\iff E=0$ when $U\geq 1$.
+Then
+$\delta(Q_{E})=\max(1+\delta(E),2)\leq 3$ (6.100)
+and
+$Q_{E}=0\iff E=0$ (6.101)
+when $U\geq 1$.
 Procedure 6.2 (Monomial Shielding).
@@
 […]
-6.2 Example
 Example 6.4.
@@
 $\exists\mathbf{x}^{*}\in\mathbb{N}^{n}:R(\mathbf{x}^{*})=0\quad\Longleftrightarrow\quad\exists(\mathbf{x}^{*},\mathbf{y}^{*})\in\mathbb{N}^{n+\ell}:\Phi=\mathbf{0}\land R^{\prime}=0.$ (6.103)
-Proof.
 ($\Rightarrow$) Set $y_{i}^{*}=(\text{product})_{i}(\mathbf{x}^{*})$ for each shield variable; constraints $y_{i}-(\text{product})_{i}=0$ are satisfied and $R^{\prime}=R$ by substitution.
@@
 For polynomial $R$ with $m$ monomials of degree $\leq d$, Algorithm 6.3 introduces $O(m\log d)$ auxiliary variables and constraints, each of degree $\leq 3$.
-Proof.
 Each monomial requires $O(\log d)$ shielding steps (degree halves per recursion). Total: $\sum_{i=1}^{m}O(\log d_{i})=O(m\log d)$. ∎
-Remark (Application to $P_{\text{merged}}$).
-For $P_{\text{merged}}=\sum_{i=1}^{m_{0}}P_{i}^{2}$ with $\delta(P_{i})\leq 3$ (from Table 2), we analyze the complexity of the polynomial reduction as follows.
+Remark.
+For
+$P_{\text{merged}}=\sum_{i=1}^{m_{0}}P_{i}^{2}$ (6.104)
+with $\delta(P_{i})\leq 3$ (from Table 1), we analyze the complexity of the polynomial reduction as follows.
 Each squared polynomial $P_{i}^{2}$ has degree at most $6$ and contains at most $s_{i}^{2}$ monomials, where $s_{i}$ denotes the monomial count of $P_{i}$.
-The total monomial count is therefore $m=\sum_{i=1}^{m_{0}}s_{i}^{2}=O(m_{0}\bar{s}^{2})$, where $\bar{s}=\max_{i}s_{i}=O(K)$ by the bounds on Zeckendorf digit sums.
+The total monomial count is therefore
+$m=\sum_{i=1}^{m_{0}}s_{i}^{2}=O(m_{0}\bar{s}^{2}),$ (6.105)
+where
+$\bar{s}=\max_{i}s_{i}=O(K)$ (6.106)
+by the bounds on Zeckendorf Digit sums.
 Since Section 4.7 establishes that $m_{0}=O(K^{3}N^{3})$ for the total number of constraints, Proposition 6.6 implies that the number of auxiliary variables required is $O(m\log 6)=O(K^{3}N^{3})$.
@@
     $P_{\text{hard}}(\mathbf{x}^{*})=0$ has solution in $\mathbb{N}^{m}\iff M$ accepts within $T(N)$ steps.
-Proof.
 Apply Algorithm 6.3 to $P_{\text{merged}}$ from (5.96). Degree bound follows from Algorithm 6.2’s invariant; solvability equivalence from Proposition 6.5; variable count from Remark Remark. ∎
@@
 Exposition.
-We establish the central equivalence: the polynomial $P_{\text{hard}}$ constructed in Section 4 faithfully encodes provability in $T_{\text{Fib}}$.
+We establish the central equivalence: the polynomial $P_{\text{hard}}$ constructed in Section 4 faithfully encodes provability in $\mathrm{T}_{\text{Fib}}$.
 The proof proceeds bidirectionally through soundness (Section 7.1) and completeness (Section 7.2), yielding the main result (Theorem 5.3).
@@
 Theorem 7.1 (Soundness).
-If there exists $\vec{x}^{*}\in\mathbb{N}^{M}$ such that $P_{\text{hard}}(\vec{x}^{*})=0$, then $T_{\text{Fib}}\vdash G_{T}$.
-Proof.
+If there exists $\vec{x}^{*}\in\mathbb{N}^{M}$ such that $P_{\text{hard}}(\vec{x}^{*})=0$, then $\mathrm{T}_{\text{Fib}}\vdash G_{\mathrm{T}}$.
 Assume $P_{\text{hard}}(\vec{x}^{*})=0$. We extract a valid proof sequence from the solution.
@@
 $\beta(c^{*},d^{*},i)=f_{i}^{*}\quad\text{for all }i\in\{1,\ldots,N^{*}\}.$ (7.108)
-Step 3. Zeckendorf representability. By Lemma 4.3, each $f_{i}^{*}$ admits a unique representation:
+Step 3. Representability. By Lemma 4.3, each $f_{i}^{*}$ admits a unique representation:
 $f_{i}^{*}=\sum_{\kappa=2}^{K}F_{\kappa}\cdot d_{i,\kappa}^{*},$ (7.109)
@@
 $f_{i}^{*}=f_{j}^{*}+f_{k}^{*},$ (7.110)
-with Zeckendorf digits computed via Exposition Exposition.
 Step 6. Target formula. By Lemma 4.14:
-$f_{N^{*}}^{*}=g_{T}.$ (7.86)
-Conclusion. The sequence $\langle f_{1}^{*},\ldots,f_{N^{*}}^{*}\rangle$ constitutes a valid proof of $G_{T}$ in $T_{\text{Fib}}$. ∎
+$f_{N^{*}}^{*}=g_{\mathrm{T}}.$ (7.111)
+Conclusion. The sequence $\langle f_{1}^{*},\ldots,f_{N^{*}}^{*}\rangle$ constitutes a valid proof of $G_{\mathrm{T}}$ in $T_{\text{Fib}}$. ∎
 7.2 Completeness
 Theorem 7.2 (Completeness).
-If $\mathrm{T}_{\text{Fib}}\vdash G_{T}$, then there exists $\vec{x}^{*}\in\mathbb{N}^{M}$ such that $P_{\text{hard}}(\vec{x}^{*})=0$.
-Proof.
-Let $\pi=\langle f_{1},\ldots,f_{N}\rangle$ be a proof of $G_{T}$. We construct $\vec{x}^{*}$ by explicit assignment.
+If $\mathrm{T}_{\text{Fib}}\vdash G_{\mathrm{T}}$, then there exists $\vec{x}^{*}\in\mathbb{N}^{M}$ such that $P_{\text{hard}}(\vec{x}^{*})=0$.
+Let $\pi=\langle f_{1},\ldots,f_{N}\rangle$ be a proof of $G_{\mathrm{T}}$. We construct $\vec{x}^{*}$ by explicit assignment.
 Step 1. Global multiplier. Set $\mathrm{T}^{*}=1$, $U^{*}=0$ (satisfying Definition 3.11).
@@
   2. 2.
-    Lemma 4.3 holds (via Zeckendorf uniqueness, Definition 3.3),
+    Lemma 4.3 holds (via uniqueness, Definition 3.3),
   3. 3.
@@
   5. 5.
-    Lemma 4.13 holds (since $f_{N}^{*}=g_{T}$).
+    Lemma 4.13 holds (since $f_{N}^{*}=g_{\mathrm{T}}$).
 Conclusion. All constraints vanish under $\vec{x}^{*}$, hence $P_{\text{hard}}(\vec{x}^{*})=0$. ∎
-7.3 Main Algorithm: Encoding Gödel’s Theorem
+7.3 Encoding Incompleteness
 We present the complete reduction in algorithmic form:
@@
 The polynomial $P_{\text{final}}$ output by the algorithm satisfies:
-$T\vdash G_{T}\quad\Longleftrightarrow\quad\exists\mathbf{x}\in\mathbb{N}^{m}:P_{\text{final}}(\mathbf{x})=0.$ (7.90)
-Proof.
+$\mathrm{T}\vdash G_{\mathrm{T}}\quad\Longleftrightarrow\quad\exists\mathbf{x}\in\mathbb{N}^{m}:P_{\text{final}}(\mathbf{x})=0.$ (7.115)
 By Theorem 4.16 (soundness and completeness of Stages 1–6), Equation (5.96) (sum-of-squares equivalence), and Proposition 6.5 (degree reduction preserves solvability). ∎
-7.4 Undecidability Consequences
+7.4 Calibration
+Logic 7.4.
+It is useful to record a simple calibration: a hypothetical total correct solver for the cubic fragment would already decide solvability for degree $\leq 4$ equations by a uniform compilation.
+Concretely, there is a computable, uniform transformation which maps any quartic equation
+$P(\vec{x})=0\qquad(\delta(P)\leq 4)$ (7.116)
+to a cubic equation
+$Q(\vec{x},\vec{y})=0\qquad(\delta(Q)\leq 3)$ (7.117)
+such that
+$\exists\vec{x}\in\mathbb{N}^{n}\;P(\vec{x})=0\quad\Longleftrightarrow\quad\exists(\vec{x},\vec{y})\in\mathbb{N}^{n+m}\;Q(\vec{x},\vec{y})=0.$ (7.118)
+The transformation is the standard shielding discipline: each quartic monomial is replaced by a fresh auxiliary variable, constrained by a short chain of quadratic/biquadratic equations (introducing only bounded-degree constraints).
+One then packages the resulting finite system into a single Diophantine equation (e.g.
+by a sum-of-squares merger), and applies the same degree-reduction protocol already used in our aggregation layer to return to total degree $\leq 3$.
+Thus, if $\exists e\,\mathsf{Solve}_{3}(e)$ held, then degree-$4$ solvability over $\mathbb{N}$ would be decidable as an immediate consequence.
+This is not our theorem, but it underscores the strength of the cubic solver hypothesis: it would already collapse the known degree-$4$ regime into the cubic one by a uniform compilation.
+Logic 7.5.
+By contrast, there is no uniform compilation from the quadratic to the cubic fragment.
+Quadratic Diophantine equations are closed under addition and squaring but not under multiplicative composition.
+Any attempt to encode a cubic monomial $x_{i}x_{j}x_{k}$ using only quadratic constraints (for example $y=x_{i}x_{j}$, $z=yx_{k}$) inevitably introduces a constraint of degree $3$.
+Thus multiplication itself is not quadratic-definable.
+Formally, the quadratic fragment defines only semilinear subsets of $\mathbb{N}^{n}$, while the cubic fragment already defines recursively enumerable sets.
+The gap is therefore categorical rather than technical: degree $3$ is the first level at which bounded verification becomes representable, and where existential closure fails uniformly.
+7.5 Cubic Incompleteness
 Corollary 7.6 (Undecidability of $\mathcal{D}_{3}$).
@@
 For $\mathcal{D}_{3}$, there exists no algorithm deciding, for arbitrary $Q(\vec{x})\in\mathbb{Z}[\vec{x}]$ with $\delta(Q)\leq 3$, whether $Q(\vec{x})=0$ has a solution in $\mathbb{Z}^{n}$.
-Proof.
-By Theorems 7.1 and 7.2, solvability of $P_{\text{hard}}$ is equivalent to $\mathrm{T}_{\text{Fib}}\vdash G_{T}$.
-By [Gödel, 1931], if $\mathrm{T}_{\text{Fib}}$ is consistent, then $G_{T}$ is independent of $\mathrm{T}_{\text{Fib}}$.
+By Theorems 7.1 and 7.2, solvability of $P_{\text{hard}}$ is equivalent to $\mathrm{T}_{\text{Fib}}\vdash G_{\mathrm{T}}$.
+By [Gödel, 1931], if $\mathrm{T}_{\text{Fib}}$ is consistent, then $G_{\mathrm{T}}$ is independent of $\mathrm{T}_{\text{Fib}}$.
 Since $\delta(P_{\text{hard}})\leq 3$ (Theorem 5.3), this exhibits a cubic equation whose solvability is undecidable.
-Encoding arbitrary Turing machine halting problems via similar constructions yields a uniform reduction from the halting problem to $\mathcal{D}_{3}$ solvability.
+Encoding arbitrary Halting Problems via similar constructions yields a uniform reduction from the Halting Problem to $\mathcal{D}_{3}$ solvability.
 ∎
 Theorem 7.7 (Uniform Halting Reduction).
-There exists a computable function $\Phi:\mathbb{N}\to\mathbb{Z}[\vec{x}]$ such that for any Turing machine $M$ with Gödel number $\ulcorner M\urcorner$:
-$\ltxml@oqmark@open\textquotedblleft\penalty 10000\thinspace\thinspace M\text{ halts on empty input}\textquotedblright\ltxml@oqmark@close{}\quad\Longleftrightarrow\quad\exists\vec{x}\in\mathbb{Z}^{k}:\Phi(\ulcorner M\urcorner)(\vec{x})=0,$ (7.91)
+There exists a computable function $\Phi:\mathbb{N}\to\mathbb{Z}[\vec{x}]$ such that for any Turing Machine, cf. Turing [1937, 1938] $M$ with Gödel Number $\ulcorner M\urcorner$:
+$M\text{ halts on empty input}\quad\Longleftrightarrow\quad\exists\vec{x}\in\mathbb{Z}^{k}:\Phi(\ulcorner M\urcorner)(\vec{x})=0,$ (7.119)
 where $\delta(\Phi(\ulcorner M\urcorner))\leq 3$ and $k$ is independent of $M$.
-Proof sketch.
-By [Matiyasevich, 1970], there exists a polynomial $P_{\text{MRDP}}(m,\vec{y})$ encoding Turing machine halting.
-Applying the guard-gadget framework (Section 3.5) and degree reduction (Theorem 5.1) yields a cubic polynomial $\Phi(\ulcorner M\urcorner)$ preserving solvability.
+By [Matiyasevich, 1970], there exists a polynomial $P_{\text{MRDP}}(m,\vec{y})$ encoding Turing Machine.
+Applying the guard-gadget framework and degree reduction (Theorem 5.2) yields a cubic polynomial $\Phi(\ulcorner M\urcorner)$ preserving solvability.
 The map $M\mapsto\Phi(\ulcorner M\urcorner)$ is computable by effective construction of guards and auxiliary variables.
 ∎
-7.5 Impredicativity
-Thesis 7.6 (Impredicativity of Degree Threshold).
+Thesis 7.8 (Impredicativity of Degree Threshold).
 Let $\mathrm{T}$ be any consistent, recursively axiomatizable theory extending Robinson Arithmetic $\mathsf{Q}$.
@@
     $\forall m\geq k\;\neg\mathsf{Dec}(m)$ (undecidability at and above threshold).
-Proof sketch.
 Suppose such $B(k)$ exists with unique $k_{0}$ satisfying $\mathrm{T}\vdash B(k_{0})$.
@@
 ∎
+Lemma 7.9 (Compactness–Reflection Barrier).
+Let
+$\Phi_{e,p,t,b}(\vec{z})=0$ (7.121)
+be the cubic system from Theorem 2.7 such that
+$\exists\vec{z}\;\Phi_{e,p,t,b}(\vec{z})=0\Longleftrightarrow\mathsf{Run}(e,p,t,b).$ (7.122)
+Fix $(e,p,b)$ and write
+$\mathsf{HaltsOut}(e,p,b)\equiv\exists t\,\mathsf{Run}(e,p,t,b)$ (7.123)
+  1. (i)
+    Bounded coherence. For each bound $B\in\mathbb{N}$, the bounded predicate
+    $\mathsf{HaltsOut}_{\leq B}(e,p,b)\;:\Longleftrightarrow\;\exists t\leq B\;\mathsf{Run}(e,p,t,b)$ (7.124)
+    is decidable, and is representable by the finite disjunction
+    $\mathsf{HaltsOut}_{\leq B}(e,p,b)\;\Longleftrightarrow\;\bigvee_{t\leq B}\;\exists\vec{z}\;\Phi_{e,p,t,b}(\vec{z})=0.$ (7.125)
+  2. (ii)
+    Compactness of the limit step.
+In any classical metatheory with $\omega$-completeness/compactness for $\Sigma_{1}$-instances (equivalently: a principle validating passage from all bounded stages to the unbounded existential),
+    $\Bigl(\forall B\;\mathsf{HaltsOut}_{\leq B}(e,p,b)\Bigr)\ \Longrightarrow\ \mathsf{HaltsOut}(e,p,b).$ (7.126)
+    Thus the unbounded existential closure is exactly the limit of its bounded approximants.
+  3. (iii)
+    Reflection constraint. Any argument that derives the global negation
+    $\neg\exists\mathsf{Close}\;\forall e,p,b\;\bigl(\mathsf{HaltsOut}(e,p,b)\leftrightarrow\mathsf{Sol}_{3}(\mathsf{Close}(e,p,b))\bigr)$ (7.127)
+    from bounded information alone must, at some point, employ the limit step of (ii) (or an equivalent reflection/compactness principle). In particular, the transition from the family
+    $\{\mathsf{HaltsOut}_{\leq B}\}_{B\in\mathbb{N}}$ (7.128)
+    to the global $\mathsf{HaltsOut}$ reintroduces precisely the unbounded closure move the cubic analysis isolates.
+Corollary 7.10 (Reflexivity of the Cubic Barrier).
+The obstruction is self-similar: proving the impossibility of a uniform cubic closure operator requires a metatheoretic compact Reflection Principle that performs the same global unbounding step that a putative uniform solver would have to internalize.
 Exposition.
@@
 This mirrors the Grelling–Nelson paradox and reflects the predicativity restriction in the type theory of Martin-Löf [1980], where universes $\mathcal{U}_{i}$ cannot quantify over themselves.
-Thesis 7.7 (Diophantine Impredicativity).
+Thesis 7.11 (Diophantine Impredicativity).
 In the Calculus of Inductive Constructions, attempting to form:
@@
 requires impredicative quantification over the decidability predicate $\mathsf{Decidable}(-)$, which itself depends on the threshold $k$ being defined.
-By Thesis 7.6, no closed term of this type exists in any consistent extension of $\mathsf{Q}$.
+By Thesis 7.8, no closed term of this type exists in any consistent extension of $\mathsf{Q}$.
 The diagonal equation $E_{k_{0}}$ acts as a negative membership witness:
@@
   1. 1.
-    Externalize the threshold: Define $B(k)$ in a strictly stronger metatheory (e.g., $\mathsf{PA}$ proves undecidability of $\mathcal{D}_{3}$ using Gödel sentences not formalizable in $\mathsf{Q}$), or
+    “Externalize the threshold”: Define $B(k)$ in a strictly stronger metatheory (e.g., $\mathsf{PA}$ proves undecidability of $\mathcal{D}_{3}$ using Gödel Sentences not formalizable in $\mathsf{Q}$), or
   2. 2.
-    Adopt impredicative axioms: E.g., $\mathsf{ZFC}$ proves existence of a least undecidable degree by quantifying over all sets of equations via the power set axiom (which is impredicative).
+    “Adopt impredicative axioms”: E.g., $\mathsf{ZFC}$ proves existence of a least undecidable degree by quantifying over all sets of equations via the power set axiom (which is impredicative).
 The cubic bound $k=3$ established here is not an internal theorem of any predicative arithmetic.
@@
   2. 2.
-    No uniform refutation exists (by Thesis 7.6, any such refutation enables constructing $E_{k_{0}}$, violating consistency).
-Thus, the threshold constitutes a Brouwerian counterexample to the law of excluded middle: a statement $P$ such that neither $P$ nor $\neg P$ admits constructive proof.
+    No uniform refutation exists (by Thesis 7.8, any such refutation enables constructing $E_{k_{0}}$, violating consistency).
+Thus, the threshold constitutes a “Brouwerian” counterexample to the Law of Excluded Middle, cf.
+Brouwer [1908]; van Atten and Sundholm [2017]: a statement $P$ such that neither $P$ nor $\neg P$ admits constructive proof.
 The degree-$3$ bound is visible only from a classical metatheory (e.g., $\mathsf{ZFC}$ proves $\mathcal{D}_{3}$ is undecidable” by encoding the Halting Problem), but no predicative subsystem of $\mathsf{PA}$ can internalize this fact.
-Thesis 7.9 (Impredicativity of Degree Threshold).
+Thesis 7.13 (Impredicativity of Degree Threshold).
 Let $\mathrm{T}$ be any consistent, effectively axiomatizable theory extending Robinson arithmetic $\mathsf{Q}$.
@@
 Then there exists no first-order formula $B(k)$ in the language of arithmetic such that $\mathrm{T}$ proves:
-  1. 1.
+  1. (i)
     $\exists!k\;B(k)$ (uniqueness of threshold),
-  2. 2.
+  2. (ii)
     $\forall m<k\;\mathsf{Dec}(m)$ (decidability below threshold),
-  3. 3.
+  3. (iii)
     $\forall m\geq k\;\neg\mathsf{Dec}(m)$ (undecidability at and above threshold).
@@
 Any attempt to define the “first” degree $k$ at which undecidability arises leads to a form of self-reference: specifying such a $k$ enables the construction of a Diophantine equation of degree $k$ whose solvability encodes the negation of that very definition.
-Consequently, the threshold of undecidability is not first-order definable within any sufficiently strong theory.
-This mirrors classical semantic paradoxes—such as the Grelling–Nelson paradox [^6] Consult [Quine, 1976] for a thorough discussion of paradoxes.—in which the act of defining:
-Is the adjective ’heterological’ a heterological word?
-yields contradiction.
-Similarly, “Does undecidability start at kk?” becomes undecidable when answered affirmatively.
-If $k$ is the first undecidable degree, the evaluation criterion shifts, making the boundary undecidable if and only if it is decidable—a fixed point in arithmetic.
-Under HBK [^7] The Heyting–Brouwer–Kolmogorov (HBK), formalized by Heyting [1930], omits proof by the law of excluded middle unless constructively justified., this reflects the failure by appealing to the law of excluded middle (LEM): asserting $\mathsf{Dec}(k)\vee\neg\mathsf{Dec}(k)$ requires a uniform method for deciding solvability across all equations in $\mathcal{D}_{k}$, or uniformly refuting such a method—a totality that can be rejected without explicit witnesses.
-Thus, impredicativity arises not merely from syntactic self-reference, but from the semantic attempt to define a global threshold for undecidability.
-The paradox is resolved constructively by recognizing that no such threshold is constructively definable: the transition from decidable to undecidable is not marked by a computable boundary, but by a failure of totality.
-In this sense, the undecidability threshold is a meta-mathematical horizon—a boundary that cannot be crossed without stepping outside the constructive framework.
-Any attempt to define it collapses into a fixed point of the form:
+Consequently, the threshold of undecidability is not first-order definable within any sufficiently strong theory, as it is the act of defining:
 At what point does the undecidability of undecidability begin?
+Acknowledgments